### v1

Abstract
In a Jacobi–Davidson (JD) type method for singular value decomposition (SVD) problems, called JDSVD, a large symmetric and generally indefinite correction equation is approximately solved iteratively at each outer iteration, which constitutes the inner iterations and dominates the overall efficiency of JDSVD. In this paper, a convergence analysis is made on the minimal residual (MINRES) method for the correction equation. Motivated by the results obtained, a preconditioned correction equation is derived that extracts useful information from current searching subspaces to construct effective preconditioners for the correction equation and is proved to retain the same convergence of outer iterations of JDSVD. The resulting method is called inner preconditioned JDSVD (IPJDSVD) method. Convergence results show that MINRES for the preconditioned correction equation can converge much faster when there is a cluster of singular values closest to a given target, so that IPJDSVD is more efficient than JDSVD. A new thick-restart IPJDSVD algorithm with deflation and purgation is proposed that simultaneously accelerates the outer and inner convergence of the standard thick-restart JDSVD and computes several singular triplets of a large matrix. Numerical experiments justify the theory and illustrate the considerable superiority of IPJDSVD to JDSVD.
keywords:
singular value decomposition, singular value, singular vector, Jacobi–Davidson method, outer iteration, inner iteration, correction equation, MINRES, convergence, preconditioning
{AMS}
65F15, 15A18, 65F10
1 Introduction
Let $A\in\mathbb{R}^{M\times N}$ be large and possibly sparse with $M\geq N$, and its singular value decomposition (SVD) be
(1.1) $A=U\Sigma V^{T},$
[…]
Given a target $\tau>0$ that is not equal to any singular value of $A$, assume that the singular values of $A$ are labeled by
(1.2) $|\sigma_{1}-\tau|<\dots<|\sigma_{\ell}-\tau|<|\sigma_{\ell+1}-\tau|\leq\dots\leq|\sigma_{N}-\tau|.$
We are interested in the $\ell$ singular triplets $(\sigma_{i},u_{i},v_{i})$, $i=1\dots,\ell$. Throughout the paper, we suppose that $\tau$ lies in the interval $(\sigma_{\min},\sigma_{\max})$ with $\sigma_{\min}$ and $\sigma_{\max}$ the smallest and largest singular values of $A$, respectively. Specifically, if $\tau$ is neither close to $\sigma_{\min}$ nor to $\sigma_{\max}$, then $(\sigma_{i},u_{i},v_{i})$, $i=1\dots,\ell$ are called interior singular triplets of $A$; otherwise they are called extreme, i.e., the largest or smallest, singular triplets. For some applications of extreme and interior singular triplets, see [17].
Projection methods [4, 21, 22, 26, 27] for large eigenvalue problems have been adapted to the computation of a partial SVD of $A$. There are four kinds of projection methods for computing approximate singular triplets of $A$: the standard extraction [2, 6, 7, 8, 14, 29], the harmonic extraction [7, 8, 9, 13, 18, 20], the refined extraction [8, 9, 10, 14, 28, 29], and the refined harmonic extraction [10, 15, 28, 29].
For all projection methods on the computation of partial SVDs, a common core task is the construction and expansion of left and right searching subspaces so that they contain as much information on the desired singular vectors as possible. There are mainly three types of approaches to construct searching subspaces for the large SVD computation. Next we review two well-known approaches, and leave out the very new but complicated one that is used in the FEAST SVDsolvers [16, 17], which compute the singular triplets with all the singular values in a given interval. The first approach is the Lanczos bidiagonalization (LBD) process [2, 4]. It reduces $A$ to a sequence of small bidiagonal matrices and generates two sets of orthonormal bases of the left and right Krylov subspaces formed by $AA^{T}$ and $A^{T}A$ with relevant starting vectors. Such searching subspaces generally favor the left and right singular vectors associated with the extreme singular values of $A$; that is, the distances between these singular vectors and the corresponding Krylov subspaces generally tend to zero fast as the LBD process proceeds. As a consequence, combined with appropriate extraction approaches, state-of-the-art LBD type algorithms [8, 14, 15, 18, 19, 24] can compute the extreme singular triplets of $A$ efficiently. However, the LBD type methods do not suit well for the computation of interior singular values of $A$.
The second approach is that used in Jacobi–Davidson (JD) type methods. The JD method proposed by Sleijpen and van der Vorst [23] for large eigenvalue problems is extended by Hochstenbach [7, 8] to the large SVD computation, referred to as JDSVD method. Wu and Stathopoulos [29] present a hybrid two-stage JDSVD method, called PHSVD. At the first stage, PHSVD transforms the SVD problem of $A$ into the mathematically equivalent eigenproblem of the cross-product matrix $A^{T}A$ and computes the desired singular values and right singular vectors with relatively low accuracy by some projection method. At the second stage, in order to obtain more accurate approximate singular triplets of $A$, it implements the JD method on the eigenproblem of the augmented $\begin{bmatrix}\begin{smallmatrix}\bm{0}&A\\ A^{T}&\bm{0}\end{smallmatrix}\end{bmatrix}$ by using the approximate singular vectors obtained in the first stage as an initial guess. Wu, Romero and Stathopoulos [28] develop the PRIMME__\__SVDS software that implements the PHSVD method with standard, refined and refined harmonic extractions.
At each step of JDSVD and the second stage of PHSVD, one must solve some large shifted linear system, called correction equation, that involves $\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ and the current approximate singular triplet, and uses its solution to expand the searching subspaces. The correction equations are approximately solved iteratively, called the inner iterations, while the outer iterations compute approximate singular triplets with respect to the searching subspaces. The least solution accuracy requirement on the correction equations and convergence speed of the inner iterations critically affect the outer convergence and dominate the overall efficiency of JDSVD, respectively. The least solution accuracy means that the outer iterations of the resulting inexact JDSVD methods behave as if all the correction equations were solved exactly. The authors [10] have figured out such accuracy requirement and proved that it generally suffices to solve all the correction equations with low or modest accuracy, i.e., the relative errors $\varepsilon\in[10^{-4},10^{-2}]$ of approximate solutions.
As will turn out, the correction equations involved in JDSVD methods are generally indefinite and may be ill conditioned, which is particularly true when the desired singular values are clustered at $\tau$; the correction equation is positive definite only when $\sigma_{1}=\sigma_{\max}$ is of interest. As a result, it is generally very hard to iteratively solve them even with low accuracy. To this end, for the overall efficiency of JDSVD, effective preconditioning correction equations is absolutely necessary. For a general large linear system $Bx=z$, the preconditioning is to transform it into a mathematically equivalent one that is easier to be solved by a suitable iterative solver, e.g., the conjugate gradient (CG) method, the minimal residual (MINRES) method, the generalized MINRES (GMRES) method, and the biconjugate gradient stabilized (BiCGstab) method [1, 5, 22]. There are a lot of preconditioning procedures [22], and among others two major classes of popular algebraic preconditioning techniques are incomplete LU factorization (ILU)-based preconditioners $LU\approx B$ and sparse approximate inverse (SAI) preconditioners $C\approx B^{-1}$.
In JDSVD-type methods, since the coefficient matrices in the correction equations involve $\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ and are generally symmetric indefinite, the ILU and SAI precondioning procedures encounter severe difficulties and are not appropriate choices in several aspects. For instance, (i) constructing an SAI preconditioner is costly. In a JDSVD method with adaptively changing shifts, the coefficient matrix in the correction equation changes at each outer iteration, which requires that one construct a new SAI preconditioner from scratch. Therefore, it may be unaffordable to do so. (ii) ILU preconditioners are generally ineffective and have intrinsic numerical instability for indefinite symmetric or highly nonsymmetric systems. (iii) The symmetry of correction equations enables us to fully exploit the merits of MINRES, but ILU and SAI preconditioners destroy the symmetry, so that we have to use a much inferior nonsymmetric solver GMRES or BiCGstab to solve the preconditioned unsymmetric correction equations. Numerically, we have observed that BiCGstab with ILU preconditioners is very often inefficient and performs even worse than the unpreconditioned MINRES, in terms of total inner iterations and total CPU time [10]. Therefore, in the context of JDSVD, other effective preconditioning approaches for the correction equation are extremely appealing.
For symmetric positive definite and indefinite linear systems, the CG and MINRES are preferences among Krylov iterative solvers, respectively. Setting aside initial vectors, their convergence is uniquely determined by the eigenvalues of an underlying coefficient matrix. These two methods converge fast when the eigenvalues are clustered, and they may converge very slowly when the coefficient matrix has relatively small eigenvalues in magnitude [5]. For such a linear system, favorable preconditioning is to preserve the symmetry of the preconditioned linear system and, meanwhile, shrink the eigenvalue inclusion interval of the preconditioned definite matrix or the two intervals containing negative and positive eigenvalues of the preconditioned indefinite coefficient matrix. For the correction equation in JDSVD, as will be clear, an ideal preconditioning procedure is to annihilate the $m$ clustered smallest eigenvalues in magnitude of the coefficient matrix that are well separated from the other eigenvalues, so that the eigenvalue distribution of the preconditioned coefficient matrix is favorable and MINRES converges fast. Our careful investigation will show that this is equivalent to forming a new, called preconditioned, correction equation, in which the orthogonal projector is now formed by the left and right singular vectors $u_{i}$ and $v_{i}$ associated with the $m$ clustered singular values $\sigma_{i},i=1,2,\ldots,m$ rather than only $\sigma_{1}$. Stacking these vectors, we obtain an $m$-dimensional eigenspace of the coefficient matrix in the correction equation associated with its $m$ clustered smallest eigenvalues in magnitude, which will be shown to eliminate the effects of these smallest eigenvalues so that MINRES converges much faster than it does for the original one.
We have observed from extensive numerical experiments that, as the outer iterations of JDSVD proceed, the current subspaces also contain increasingly richer information on the left and right singular vectors $u_{i}$ and $v_{i}$, $i=1,\ldots,m$ corresponding to the $m$ clustered singular values. These reasonably good approximate singular vectors form a good approximate eigenspace of the coefficient matrix in the correction equation corresponding to its $m$ clustered smallest eigenvalues in magnitude. We will fully exploit such rich information to derive a symmetric preconditioned correction equation that eliminates the effects of those smallest eigenvalues in magnitude of the coefficient matrix, so that MINRES can be accelerated greatly.
We first make a detailed convergence analysis on MINRES for the correction equation in JDSVD for computing one singular triplet $(\sigma_{1},u_{1},v_{1})$. Then we consider the case that there are $m$ clustered singular values closest to the target $\tau$, and show that MINRES may converge very slowly. For the preconditioned correction equation, we prove an extremely important property: its solution has the same quality as the solution to the standard correction equation when expanding the current subspaces; that is, the solution to the preconditioned correction equation expands the subspaces as effectively as the solution to the standard one and, therefore, and does not spoil the subsequent outer iterations. Then we establish convergence results on MINRES for the preconditioned correction equation, and prove that it can converge substantially faster than it does for the standard one. In computations, we shall set up some necessary criteria and propose an adaptive approach to pick up appropriate approximate singular triplets to form the preconditioned correction equation. The resulting method is called the inner-preconditioned JDSVD (IPJDSVD) method. On the other hand, importantly and practically, we propose a new thick-restart IPJDSVD algorithm with deflation and purgation that computes $\ell$ singular triplets of $A$. Whenever $m>k_{\min}$, the initial subspace dimension of the standard thick-restart, the new thick-restart is more effective than the standard one and saves outer iterations.
Since the goal of this paper is to propose an effective preconditioning procedure for the correction equation in JDSVD and a new thick-restart strategy, we shall only consider the standard extraction-based JDSVD method and the shifts in all the correction equations are $\tau$. However, the theoretical results and analysis on the correction equations and preconditioning are directly applicable to the harmonic, refined and refined harmonic extraction-based JDSVD methods and their adaptively-changing-shift variants [10], so are the corresponding thick-restart algorithms.
In section 2 we review the standard extraction-based JDSVD method. In section 3 we make a convergence analysis on MINRES for the correction equation in JDSVD. In section 4 we derive the preconditioned correction equation, and extend the results and analysis in section 3 to the preconditioned correction equation. We consider how to dynamically form a preconditioned correction equation in computations. In section 5 we propose the thick-restart IPJDSVD with deflation and purgation to compute $\ell$ singular triplets of $A$. Numerical experiments are reported in section 6 to illustrate the high performance of our preconditioning procedure and the thick-restart IPJDSVD. We conclude the paper in section 7.
Throughout this paper, denote by $\|X\|$ the 2-norm of a vector or matrix $X$, by $\kappa(X)=\sigma_{\max}(X)/\sigma_{\min}(X)$ the condition number of $X$ with $\sigma_{\max}(X)$ and $\sigma_{\min}(X)$ the largest and smallest singular values of $X$, respectively, and by $I_{i}$ and $\bm{0}_{i,j}$ the identity matrix of order $i$ and zero matrix of size $i\times j$, respectively. The subscripts of identity and zero matrices are omitted whenever their sizes are clear from the context.
2 The basic JDSVD method
We review the basic JDSVD method [7] that is based on the standard extraction to compute the singular triplet $(\sigma_{1},u_{1},v_{1})$ of $A$.
At the $k$th outer iteration, given two $k$-dimensional left and right searching subspaces $\mathcal{U}\subset\mathbb{R}^{M}$ and $\mathcal{V}\subset\mathbb{R}^{N}$, the standard extraction approach seeks for scalars $\theta_{i}>0$ and unit-length vectors $\tilde{u}_{i}\in\mathcal{U}$ and $\tilde{v}_{i}\in\mathcal{V}$ that satisfy the double-orthogonality
(2.1) $\qquad A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i}\perp\mathcal{U}\quad\mbox{and}\quad A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i}\perp\mathcal{V},\qquad\qquad i=1,\dots,k.$
The $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ are called the Ritz approximations of $A$ with respect to the left and right subspaces $\mathcal{U}$ and $\mathcal{V}$. In [7], condition (2.1) is denoted as $(A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i},A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i})\perp\perp(\mathcal{U},\mathcal{V}),\ i=1,\dots,k.$ Label $|\theta_{1}-\tau|\leq|\theta_{2}-\tau|\leq\dots\leq|\theta_{k}-\tau|$. The method takes $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ as an approximation to $(\sigma_{1},u_{1},v_{1})$.
Let $\widetilde{U}\in\mathbb{R}^{M\times k}$ and $\widetilde{V}\in\mathbb{R}^{N\times k}$ be orthonormal basis matrices of $\mathcal{U}$ and $\mathcal{V}$. Then the left and right Ritz vectors $\tilde{u}_{i}=\widetilde{U}c_{i}$ and $\tilde{v}_{i}=\widetilde{V}d_{i}$ with unit-length $c_{i}\in\mathbb{R}^{k}$ and $d_{i}\in\mathbb{R}^{k}$. Define $H=\widetilde{U}^{T}A\widetilde{V}$. Then (2.1) is equivalent to
(2.2) $\qquad Hd_{i}=\theta_{i}c_{i}\qquad\mbox{and}\qquad H^{T}c_{i}=\theta_{i}d_{i},\qquad\qquad i=1,\dots,k.$
This means that the $(\theta_{i},c_{i},d_{i})$ are the singular triplets of $H$ and
(2.3) $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=(\theta_{1},\widetilde{U}c_{1},\widetilde{V}d_{1}),$
whose residual is defined as
(2.4) $r=r(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=\begin{bmatrix}A\tilde{v}_{1}-\theta_{1}\tilde{u}_{1}\\ A^{T}\tilde{u}_{1}-\theta_{1}\tilde{v}_{1}\end{bmatrix}.$
By (2.1), we particularly have
(2.5) $r\perp\perp(\tilde{u}_{1},\tilde{v}_{1}).$
$r=\bm{0}$ if and only if $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ is an exact singular triplet of $A$. If
(2.6) $\|r\|\leq\|A\|\cdot\textit{tol}$
for a user prescribed tolerance $\textit{tol}>0$, we claim that $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ converges.
If $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ does not yet converge, JDSVD expands $\mathcal{U}$ and $\mathcal{V}$ in the following way: It first solves the large symmetric and generally indefinite correction equation
(2.7) $\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r$
for $(s,t)\perp\perp(\tilde{u}_{1},\tilde{v}_{1})$ approximately using the preferable MINRES method, and then uses the approximate solution $\tilde{s}$ and $\tilde{t}$ to expand $\mathcal{U}$ and $\mathcal{V}$, respectively. Specifically, we orthonormalize $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ and obtain the expansion vectors
$u_{+}=\frac{(I-\widetilde{U}\widetilde{U}^{T})\tilde{s}}{\|(I-\widetilde{U}\widetilde{U}^{T})\tilde{s}\|}\qquad\mbox{and}\qquad v_{+}=\frac{(I-\widetilde{V}\widetilde{V}^{T})\tilde{t}}{\|(I-\widetilde{V}\widetilde{V}^{T})\tilde{t}\|}.$
We then form the orthonormal basis matrices
(2.8) $\widetilde{U}_{\mathrm{new}}=[\widetilde{U},u_{+}]\qquad\mbox{and}\qquad\widetilde{V}_{\mathrm{new}}=[\widetilde{V},v_{+}],$
and generate the expanded $(k+1)$-dimensional left and right searching subspaces
$\mathcal{U}_{\mathrm{new}}=\mathrm{span}\{\widetilde{U}_{\mathrm{new}}\}=\mathrm{span}\{\widetilde{U},\tilde{s}\}\quad\mbox{and}\quad\mathcal{V}_{\mathrm{new}}=\mathrm{span}\{\widetilde{V}_{\mathrm{new}}\}=\mathrm{span}\{\widetilde{V},\tilde{t}\}.$
The basic JDSVD seeks for a new Ritz approximation to $(\sigma_{1},u_{1},v_{1})$ with respect to $\mathcal{U}_{\mathrm{new}}$ and $\mathcal{V}_{\mathrm{new}}$, and proceeds in such a way until $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ converges.
3 Convergence analysis on MINRES for solving (2.7)
Observe that the coefficient matrix of system (2.7) is the restriction of $B=\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ to the subspace
$(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}:=\{(x,y)|x\in\mathbb{R}^{M},y\in\mathbb{R}^{N},x\perp\tilde{u}_{1},y\perp\tilde{v}_{1}\},$
and maps the subspace to itself. Let $\widetilde{P}_{1}\in\mathbb{R}^{M\times(M-1)}$ and $\widetilde{Q}_{1}\in\mathbb{R}^{N\times(N-1)}$ be such that the matrices $[\tilde{u}_{1},\widetilde{P}_{1}]$ and $[\tilde{v}_{1},\widetilde{Q}_{1}]$ are orthogonal. Then the columns of
(3.1) $\widetilde{W}_{1}=\begin{bmatrix}\widetilde{P}_{1}&\\ &\widetilde{Q}_{1}\end{bmatrix}\in\mathbb{R}^{(M+N)\times(M+N-2)}$
form an orthonormal basis of $(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$, and the orthogonal projector in (2.7) is
$\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}=\begin{bmatrix}\widetilde{P}_{1}\widetilde{P}_{1}^{T}&\\ &\widetilde{Q}_{1}\widetilde{Q}_{1}^{T}\end{bmatrix}=\widetilde{W}_{1}\widetilde{W}_{1}^{T}.$
Denote by $w=(s^{T},t^{T})^{T}$. Then the correction equation (2.7) is
(3.2) $\widetilde{B}_{1}w=-r,$
where
(3.3) $\widetilde{B}_{1}=\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\quad\ \mbox{with}\quad\ \widetilde{B}_{1}^{\prime}=\widetilde{W}_{1}^{T}B\widetilde{W}_{1}=\begin{bmatrix}-\tau I_{M-1}&\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}\\ \widetilde{Q}_{1}^{T}A^{T}\widetilde{P}_{1}&-\tau I_{N-1}\end{bmatrix}.$
Let $\tilde{w}_{j}=(\tilde{s}_{j}^{T},\tilde{t}_{j}^{T})^{T}$ be the approximate solution of (3.2) computed by MINRES at the $j$th iteration. Then the residual of $\tilde{w}_{j}$ is
(3.4) $r_{\mathrm{in},j}=\widetilde{B}_{1}\tilde{w}_{j}+r.$
We now present an equivalent form of (2.7) that will be used later.
Theorem 3.1.
Let $\widetilde{B}_{1}^{\prime}$ be defined as in (3.3). The solution $w=\widetilde{W}_{1}z$ of (2.7), where $z\in\mathbb{R}^{M+N-2}$ is the solution to the linear system
(3.5) $\widetilde{B}_{1}^{\prime}z=-r^{\prime}\qquad\mbox{with}\qquad r^{\prime}=\widetilde{W}_{1}^{T}r.$
If MINRES is applied to (3.2) and (3.5) with the initial guesses $\tilde{w}_{0}$ and $\tilde{z}_{0}$ related by $\widetilde{w}_{0}=\widetilde{W}_{1}\tilde{z}_{0}$, then the $j$th approximate solutions $\tilde{w}_{j}$ and $\tilde{z}_{j}$ satisfy $\tilde{w}_{j}=\widetilde{W}_{1}\tilde{z}_{j}$, $j=1,2,\ldots,M+N-2$. Write $r_{\mathrm{in},j}^{\prime}=\widetilde{B}_{1}^{\prime}\tilde{z}_{j}+r^{\prime}$. Then
(3.6) $\|r_{\mathrm{in},j}\|=\|r^{\prime}_{\mathrm{in},j}\|,\qquad j=0,1,\ldots,M+N-2.$
Proof.
By (2.1) and (2.7), since $r\in(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$ and $w\in(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$, we have
$r=\widetilde{W}_{1}\widetilde{W}_{1}^{T}r=\widetilde{W}_{1}r^{\prime}\qquad\mbox{and}\qquad w=\widetilde{W}_{1}\widetilde{W}_{1}^{T}w=\widetilde{W}_{1}z$
with $r^{\prime}=\widetilde{W}_{1}^{T}r$ and $z=\widetilde{W}_{1}^{T}w$. Substituting them and (3.3) into (3.2) yields
(3.7) $\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\widetilde{W}_{1}z=\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}z=-\widetilde{W}_{1}r^{\prime},$
premultiplying which by $\widetilde{W}_{1}^{T}$ leads to (3.5).
Write the $j$-dimensional Krylov subspace
$\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})=\mathrm{span}\{r_{\mathrm{in},0},\widetilde{B}_{1}r_{\mathrm{in},0},\widetilde{B}_{1}^{2}r_{\mathrm{in},0},\dots,\widetilde{B}_{1}^{j-1}r_{\mathrm{in},0}\}.$
Since $\tilde{w}_{j}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})$ satisfies
(3.8) $\|r_{\mathrm{in},j}\|=\|\widetilde{B}_{1}\tilde{w}_{j}+r\|=\min_{\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})}\|\widetilde{B}_{1}\tilde{w}+r\|,$
by definition it is straightforward to verify from (3.3) and (3.4) that
$\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})=\mathcal{K}_{j}(\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T},\widetilde{W}_{1}r^{\prime}_{\mathrm{in},0})=\widetilde{W}_{1}\mathcal{K}_{j}(\widetilde{B}_{1}^{\prime},r_{\mathrm{in},0}^{\prime});$
that is, for any $\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})$, there exists $\tilde{z}\in\mathcal{K}_{j}(\widetilde{B}_{1}^{\prime},r_{\mathrm{in},0}^{\prime})$ such that $\tilde{w}=\widetilde{W}_{1}\tilde{z}$ and thus $\tilde{z}=\widetilde{W}_{1}^{T}\tilde{w}$, $j=1,2,\ldots,M+N-2$. Therefore, from (3.8) we have
$\|r_{\mathrm{in},j}\|=\min_{\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})}\|\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\tilde{w}+\widetilde{W}_{1}r^{\prime}\|=\min_{\tilde{z}\in\mathcal{K}_{j}(\widetilde{B}_{1}^{\prime},r_{\mathrm{in},0}^{\prime})}\|\widetilde{B}_{1}^{\prime}\tilde{z}+r^{\prime}\|=\|r_{\mathrm{in},j}^{\prime}\|,$
which proves (3.6) and $\tilde{w}_{j}=\widetilde{W}_{1}\tilde{z}_{j}$. ∎
Theorem 3.1 shows that the convergence of MINRES for (3.2) and (3.5) is identical. Without loss of generality, we take initial guesses $\tilde{w}_{0}=\bm{0}$ and $\tilde{z}_{0}=\bm{0}$ in MINRES for (3.2) and (3.5), so that their initial residuals are
(3.9) $r_{\mathrm{in},0}=r\qquad\mbox{and}\qquad r_{\mathrm{in},0}^{\prime}=r^{\prime}.$
Denote by $\tilde{\lambda}_{i},i=1,\dots,M+N-2$ the eigenvalues of $\widetilde{B}_{1}^{\prime}$ in (3.3). It is known from [5] that the residual norm of the approximate solution $\tilde{z}_{j}$ of (3.5) satisfies
(3.10) $\frac{\|r_{\mathrm{in},j}^{\prime}\|}{\|r^{\prime}\|}\leq\min_{\pi\in\Pi_{j}}\max_{1\leq i\leq M+N-2}\left|\pi\left(\tilde{\lambda}_{i}\right)\right|,$
where $\Pi_{j}$ denotes the set of polynomials of degree no more than $j$ with $\pi(0)=1$ for $\pi\in\Pi_{j}$. Keep in mind (3.6). Then there are the following estimates [5, Chapter 3].
Theorem 3.2 ([5, Chapter 3]).
(i) If the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in the interval $[-\beta,-\alpha]$ or $[\alpha,\beta]$ with $0<\alpha<\beta$, then
(3.11) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{\beta}/\sqrt{\alpha}}\right)^{j};$
(ii) If the $\tilde{\lambda}_{i}\in[-\beta_{1},-\alpha_{1}]\cup[\alpha_{2},\beta_{2}]$ with $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}>0$ and $\beta_{2}-\alpha_{2}=\beta_{1}-\alpha_{1}$, then
(3.12) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{\beta_{1}\beta_{2}}/\sqrt{\alpha_{1}\alpha_{2}}}\right)^{\left[\frac{j}{2}\right]},$
where $\left[\frac{j}{2}\right]$ denotes the integer part of $\frac{j}{2}$.
This theorem indicates that the smaller $\beta-\alpha$ or $\beta_{1}-\alpha_{1}$ is, the closer is $\beta/\alpha$ or $(\beta_{1}\beta_{2})/(\alpha_{1}\alpha_{2})$ to one and thus the faster MINRES converges. Bound (3.11) is the same as that for the CG method that solves the same linear system [5], and the convergence factor is $\frac{\sqrt{\kappa(\widetilde{B}_{1}^{\prime})}-1}{\sqrt{\kappa(\widetilde{B}_{1}^{\prime})}+1}$ with $\kappa(\widetilde{B}_{1}^{\prime})=\beta/\alpha$. Suppose that $[-\beta_{1},-\alpha_{1}]$ and $[\alpha_{2},\beta_{2}]$ are symmetric with respect to the origin, i.e., $\alpha_{1}=\alpha_{2}=\alpha$, $\beta_{1}=\beta_{2}=\beta$. Then the factor in the right hand side of (3.12) is $\frac{\kappa(\widetilde{B}_{1}^{\prime})-1}{\kappa(\widetilde{B}_{1}^{\prime})+1}$, the same as that in the steepest gradient method for a positive definite (3.5) with the eigenvalue interval $[\alpha,\beta]$, but the power is $\left[\frac{j}{2}\right]$ rather than $j$. As a result, MINRES may converge much more slowly for a highly indefinite linear system than it does for a definite one.
Before proceeding, we should notice that $r=\mathbf{0}$ when $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=(\sigma_{1},u_{1},v_{1})$. In this case we terminate JDSVD. For later use when $\tilde{u}_{1}\not=u_{1}$ and $\tilde{v}_{1}\not=v_{1}$, we next suppose any virtual right hand side $r\not=\mathbf{0}$ satisfying (2.5). Based on Theorems 3.13.2, we present the following convergence results on MINRES for (2.7).
Theorem 3.3.
For $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$ in (2.7), and suppose virtually $r\not=\mathbf{0}$ satisfying $r\perp\perp(u_{1},v_{1})$. Then (i) if $\tau>(\sigma_{\max}+\sigma_{\max,2})/2$ with $\sigma_{\max,2}$ being the second largest singular value of $A$, then $\sigma_{1}=\sigma_{\max}$ and
(3.13) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,2})/(\tau-\sigma_{\max,2})}}\right)^{j};$
(ii) if $\tau<(\sigma_{\min}+\sigma_{\min,2})/2$ with $\sigma_{\min,2}$ being the second smallest singular value of $A$, then $\sigma_{1}=\sigma_{\min}$ and
(3.14) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\left(1-\frac{2}{1+\sqrt{(\sigma_{\max}^{2}-\tau^{2})/(\sigma_{\min,2}^{2}-\tau^{2})}}\right)^{\left[\frac{j-j_{o}}{2}\right]},$
where $j_{o}=\min\{1,M-N\}$; (iii) if $(\sigma_{\min}+\sigma_{\min,2})/2<\tau<(\sigma_{\max}+\sigma_{\max,2})/2$, then $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$ and
(3.15) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau)/|\tau-\sigma_{2}|}\right)^{\left[\frac{j}{2}\right]}.$
Proof.
By (1.1), for $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$, take $\widetilde{P}_{1}=P_{1}$ and $\widetilde{Q}_{1}=Q_{1}$ in (3.1) with
(3.16) $P_{1}=[u_{2},\dots,u_{M}]\qquad\mbox{and}\qquad Q_{1}=[v_{2},\dots,v_{N}],$
and rewrite the resulting $\widetilde{B}_{1}^{\prime}=B_{1}^{\prime}$ in (3.3). Then
[…]
Therefore, the eigenvalues $\lambda_{i}$ of $B_{1}^{\prime}$ are
(3.18) ±$\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau\qquad\mbox{and}\qquad\underbrace{-\tau,\dots,-\tau}_{M-N}.$
Case (i): If $\tau>(\sigma_{\max}+\sigma_{\max,2})/2$, then clearly $\sigma_{\max}$ is the closest to $\tau$, i.e., $\sigma_{1}=\sigma_{\max}$. In this case, all the singular values but $\sigma_{1}$ of $A$ are smaller than $\tau$, and
(3.19) $\lambda_{i}\in[-(\sigma_{\max,2}+\tau),-(\tau-\sigma_{\max,2})]$
by noticing that $\sigma_{2}=\sigma_{\max,2}$. Exploiting (3.11), we obtain (3.13).
Case (ii): If $\tau<(\sigma_{\min}+\sigma_{\min,2})/2$, then $\sigma_{1}=\sigma_{\min}$, and
$\lambda_{i}\in\{-\tau\}\cup[-(\sigma_{\max}+\tau),-(\sigma_{\min,2}+\tau)]\cup[\sigma_{\min,2}-\tau,\sigma_{\max}-\tau],$
where $\{-\tau\}$ consists of an isolated point $-\tau$, which is an $(M-N)$-multiple eigenvalue of $B_{1}^{\prime}$ and disappears for $M=N$. Construct a degree $j$ polynomial $q\in\Pi_{j}$ of form
$q(\lambda)=\left(\frac{\tau+\lambda}{\tau}\right)^{j_{o}}\pi(\lambda)\qquad\mbox{with}\qquad\pi\in\Pi_{j-j_{o}},$
where $j_{o}=\min\{1,M-N\}$. Since $q(0)=\pi(0)=1$ and
$\left|\frac{\tau+\lambda}{\tau}\right|^{j_{o}}\leq\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}$
over the eigenvalues in (3.18), by Theorem 3.1 and relations (3.9)–(3.10), we obtain
$\frac{\|r_{\mathrm{in},j}\|}{\|r\|}$$\leq$ $\min_{\pi\in\Pi_{j-j_{o}}}\max_{\lambda_{i}\in\{-\tau,\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau\}}|q(\lambda_{i})|$
$\leq$ $\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\min_{\pi\in\Pi_{j-j_{o}}}\max_{i=2,\ldots,N}|\pi(\pm\sigma_{i}-\tau)|$
$\leq$ $\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\min_{\pi\in\Pi_{j-j_{o}}}\max_{\lambda\in\Delta}|\pi(\lambda)|,$
where
$\Delta=[-(\sigma_{\max}+\tau),-(\sigma_{\min,2}+\tau)]\cup[\sigma_{\min,2}-\tau,\sigma_{\max}-\tau].$
Notice that the above positive and negative intervals have the same lengths $\sigma_{\max}-\sigma_{\min,2}$. Bound (3.14) follows from applying (3.12) to the above min-max problem.
Case (iii): If $(\sigma_{\min}+\sigma_{\min,2})/2<\tau<(\sigma_{\max}+\sigma_{\max,2})/2$, denote by $\sigma_{\rm L}$ and $\sigma_{\rm R}$ respectively the singular values of $A$ closest to $\tau$ from the left and right among $\sigma_{2},\dots,\sigma_{N}$. In this case, the $\lambda_{i}$ lie in
$[-(\sigma_{\max}+\tau),-(\tau-\sigma_{\rm L})]\cup[\sigma_{\rm R}-\tau,\sigma_{\max}-\tau]$
(3.20) $\subset[-(\sigma_{\max}+\tau),-|\tau-\sigma_{2}|]\cup[|\tau-\sigma_{2}|,\sigma_{\max}+\tau].$
Notice that the two intervals in (3.20) have the same lengths $\sigma_{\max}+\tau-|\tau-\sigma_{2}|$. Then (3.14) follows from applying (3.12) to the interval union in (3.20). ∎
Remark 1.
The convergence rates of MINRES, respectively, depend on
(3.21) $\gamma_{1}=\frac{\tau-\sigma_{\max,2}}{\tau+\sigma_{\max,2}},\qquad\gamma_{2}=\sqrt{\frac{\sigma_{\min,2}^{2}-\tau^{2}}{\sigma_{\max}^{2}-\tau^{2}}},\qquad\gamma_{3}=\frac{|\tau-\sigma_{2}|}{\sigma_{\max}+\tau}.$
The better $\sigma_{\max,2}$, $\sigma_{\min,2}$ or $\sigma_{2}$ is separated from $\tau$, the smaller the convergence rate $1-\frac{2\sqrt{\gamma_{1}}}{1+\sqrt{\gamma_{1}}}$, $1-\frac{2\gamma_{2}}{1+\gamma_{2}}$ or $1-\frac{2\gamma_{3}}{1+\gamma_{3}}$ in (3.13)–(3.15) is, so that MINRES converges faster.
Remark 2.
For $\sigma_{1}=\sigma_{\max}$ or $\sigma_{\min}$, which correspond to $\tau\approx\sigma_{\max}$ or $\sigma_{\min}$, we have
$\gamma_{1}\approx\frac{\sigma_{\max}-\sigma_{\max,2}}{\sigma_{\max}+\sigma_{\max,2}}\qquad\mbox{and}\qquad\gamma_{2}\approx\sqrt{\frac{\sigma_{\min,2}^{2}-\sigma_{\min}^{2}}{\sigma_{\max}^{2}-\sigma_{\min}^{2}}}.$
Suppose $\sigma_{\max}\gg\sigma_{\min}$, $\sigma_{\max,2}\approx\sigma_{\max}$ and $\sigma_{\min,2}\approx\sigma_{\min}$. Then
$\frac{\sqrt{\gamma_{1}}}{\gamma_{2}}\approx\sqrt{\frac{\sigma_{\max}(\sigma_{\max}-\sigma_{\max,2})}{2(\sigma_{\min,2}-\sigma_{\min})(\sigma_{\min,2}+\sigma_{\min})}}\approx\sqrt{\frac{\sigma_{\max}}{2\sigma_{\min}}}>1$
if $\sigma_{\max}-\sigma_{\max,2}\approx\sigma_{\min,2}-\sigma_{\min}$, and the ratio is bigger if $\sigma_{\max}-\sigma_{\max,2}>\sigma_{\min,2}-\sigma_{\min}$ considerably, which is often true in applications. For an interior $\sigma_{1}\approx\tau$, we have
$\gamma_{3}\approx\frac{|\sigma_{1}-\sigma_{2}|}{\sigma_{\max}+\sigma_{1}},$
which is generally comparable to $\gamma_{1}$ in size. Since $\gamma_{1},\gamma_{3}<1$, we generally have $\sqrt{\gamma_{1}}>\gamma_{3}$ considerably. In summary, for $\sigma_{1}=\sigma_{\max}$, the linear system (2.7) is negative definite with $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$, and MINRES is generally faster than it is for the indefinite system (2.7) when $\sigma_{1}=\sigma_{\min}$ or an interior singular value.
Starting with Theorem 3.3, we are able to analyze the convergence of MINRES when $\tilde{u}_{1}\approx u_{1}$ and $\tilde{v}_{1}\approx v_{1}$. First, we show that $\widetilde{B}_{1}^{\prime}$ in (3.3) is orthogonally similar to $B_{1}^{\prime}$ in (3.17) with squared errors of $\tilde{u}_{1}$ and $\tilde{v}_{1}$.
Theorem 3.4.
Let $\phi=\angle(\tilde{u}_{1},u_{1})$ and $\psi=\angle(\tilde{v}_{1},v_{1})$ be the acute angles between $\tilde{u}_{1}$ and $u_{1}$ and between $\tilde{v}_{1}$ and $v_{1}$. Then $\widetilde{B}_{1}^{\prime}$ is orthogonally similar to $B_{1}^{\prime}$ with the error
(3.22) $\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|\leq\|A\|\left(\sin\phi+\sin\psi\right)^{2},$
where $Z$ is some $(M+N-2)$-by-$(M+N-2)$ orthogonal matrix; see (3.31).
Proof.
By (3.16), expand the orthonormal $\widetilde{P}_{1}$ and $\widetilde{Q}_{1}$ in (3.1) as the orthogonal sums
(3.23) $\widetilde{P}_{1}=P_{1}E_{1}+u_{1}E_{2}\qquad\mbox{and}\qquad\widetilde{Q}_{1}=Q_{1}F_{1}+v_{1}F_{2},$
where $E_{1}\in\mathbb{R}^{(M-1)\times(M-1)}$, $E_{2}\in\mathbb{R}^{1\times(M-1)}$ and $F_{1}\in\mathbb{R}^{(N-1)\times(N-1)}$, $F_{2}\in\mathbb{R}^{1\times(N-1)}$. Then we have
(3.24) $\sin\phi=\|\widetilde{P}_{1}^{T}u_{1}\|=\|E_{2}\|\qquad\mbox{and}\qquad\sin\psi=\|\widetilde{Q}_{1}^{T}v_{1}\|=\|F_{2}\|.$
Let the SVDs of the square matrices $E_{1}$ and $F_{1}$ be
(3.25) $X_{1}^{T}E_{1}X_{2}=\Sigma_{E}\qquad\mbox{and}\qquad Y_{1}^{T}F_{1}Y_{2}=\Sigma_{F}.$
Notice $\|I-\Sigma_{E}\|\leq\|I-\Sigma_{E}^{T}\Sigma_{E}\|$ as $\|\Sigma_{E}\|\leq 1$ and the diagonal $\Sigma_{E}$ is square and nonnegative definite. Making use of (3.25) and $E_{1}^{T}E_{1}+E_{2}^{T}E_{2}=I$, we obtain
(3.26) $\|I-\Sigma_{E}\|\leq\|X_{2}^{T}(I-E_{1}^{T}E_{1})X_{2}\|=\|E_{2}^{T}E_{2}\|=\|E_{2}\|^{2}=\sin^{2}\phi,$
where the last equality holds by (3.24). Likewise,
(3.27) $\|I-\Sigma_{F}\|\leq\sin^{2}\psi.$
Since $P_{1}^{T}Av_{1}=\bm{0}$ and $u_{1}^{T}AQ_{1}=\bm{0}$, it is known from (3.17) and (3.23) that $\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}=E_{1}^{T}\Sigma_{1}^{\prime}F_{1}+\sigma_{1}E_{2}^{T}F_{2}$. Exploiting (3.25), we have
(3.28) $\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}=X_{2}\Sigma_{E}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}\Sigma_{F}Y_{2}^{T}+\sigma_{1}E_{2}^{T}F_{2}.$
Define
(3.29) $\widehat{\Sigma}_{1}^{\prime}=X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}=X_{2}\Sigma_{1}^{\prime\prime}Y_{2}^{T}\quad\mbox{with}\quad\Sigma_{1}^{\prime\prime}=X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}.$
Subtracting $\widehat{\Sigma}_{1}^{\prime}$ from both hand sides of (3.28) and taking $2$-norms, we obtain
(3.30) $\|\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}\!-\widehat{\Sigma}_{1}^{\prime}\|$ $=$ $\|X_{2}(\Sigma_{E}\Sigma_{1}^{\prime\prime}\Sigma_{F}-\Sigma_{1}^{\prime\prime})Y_{2}^{T}+\sigma_{1}E_{2}^{T}F_{2}\|$
$=$ $\|X_{2}\left((\Sigma_{E}-I)\Sigma_{1}^{\prime\prime}\Sigma_{F}+\Sigma_{1}^{\prime\prime}(\Sigma_{F}-I)\right)Y_{2}^{T}+\sigma_{1}E_{2}^{T}F_{2}\|$
$\leq$ $\|\Sigma_{E}-I\|\|\Sigma_{1}^{\prime\prime}\|+\|\Sigma_{1}^{\prime\prime}\|\|\Sigma_{F}-I\|+\sigma_{1}\|E_{2}\|\|F_{2}\|$
$\leq$ $\|A\|(\sin^{2}\phi+\sin^{2}\psi+\sin\phi\sin\psi)$
$\leq$ $\|A\|(\sin\phi+\sin\psi)^{2},$
where the second inequality holds by making use of (3.24), (3.26), (3.27) and $\|\Sigma_{1}^{\prime\prime}\|=\|\Sigma_{1}^{\prime}\|\leq\|A\|$, $\sigma_{1}\leq\|A\|$. Define the orthogonal matrix
(3.31) $Z=\begin{bmatrix}X_{1}X_{2}^{T}&\\ &Y_{1}Y_{2}^{T}\end{bmatrix}\in\mathbb{R}^{(M+N-2)\times(M+N-2)}.$
By the definitions of $\widetilde{B}_{1}^{\prime}$ and $B_{1}^{\prime}$ in (3.3) and (3.17) as well as that of $\widehat{\Sigma}_{1}^{\prime}$ in (3.29), a manipulation shows
$\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|=\|\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}-X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}\|=\|\widetilde{P}_{1}^{T}\!A\widetilde{Q}_{1}\!-\widehat{\Sigma}_{1}^{\prime}\|.$
Applying (3.30) to the above establishes (3.22). ∎
With the help of Theorems 3.33.4, we can establish the convergence results on MINRES for the correction equation (3.2), i.e., (2.7).
Theorem 3.5.
With the notations of Theorems 3.4, suppose that $\tilde{u}_{1}$ and $\tilde{v}_{1}$ satisfy
(3.32) $\delta:=\|A\|(\sin\phi+\sin\psi)^{2}<|\sigma_{2}-\tau|.$
(i) If $\sigma_{1}=\sigma_{\max}$, then
(3.33) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,2}+\delta)/(\tau-\sigma_{\max,2}-\delta)}}\right)^{j};$
(ii) if $\sigma_{1}\!=\!\sigma_{\min}$ and writing $j_{o}\!=\!\min\{1,M\!-\!N\}$, then
(3.34) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}+\delta}{\tau}\right)^{j_{o}}\left(1-\frac{2}{1+\sqrt{\frac{(\sigma_{\max}+\delta)^{2}-\tau^{2}}{(\sigma_{\min,2}-\delta)^{2}-\tau^{2}}}}\right)^{\left[\frac{j-j_{o}}{2}\right]};$
(iii) if $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$, then
(3.35) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau+\delta)/(|\sigma_{2}-\tau|-\delta)}\right)^{\left[\frac{j}{2}\right]}.$
Proof.
Since the matrix $Z$ in (3.31) is orthogonal, the eigenvalues $\lambda_{i}$ of $Z^{T}B_{1}^{\prime}Z$ are those of $B_{1}^{\prime}$, i.e., the ones in (3.18). Label the eigenvalues $\lambda_{i}$ and $\tilde{\lambda}_{i}$ of $B_{1}^{\prime}$ and $\widetilde{B}_{1}^{\prime}$ in the ascending or descending order. By the eigenvalue perturbation theory (cf. Theorem 10.3.1 of [21]), we obtain
(3.36) $|\tilde{\lambda}_{i}-\lambda_{i}|\leq\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|\leq\delta,$
where the last inequality holds by (3.22) and (3.32).
(i) For $\sigma_{1}=\sigma_{\max}$, since the eigenvalues $\lambda_{i}$ of $Z^{T}B_{1}^{\prime}Z$ lie in the interval in (3.19), bound (3.36) shows that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
(3.37) $[-(\sigma_{\max,2}+\tau)-\delta,-(\tau-\sigma_{\max,2})+\delta].$
Note that $\sigma_{\max,2}=\sigma_{2}$. Condition (3.32) indicates $\tau-\sigma_{\max,2}>\delta$. Then taking $\alpha=\tau-\sigma_{\max,2}-\delta$ and $\beta=\tau+\sigma_{\max,2}+\delta$ in (3.11) of Theorem 3.2 proves (3.33).
(ii) Since that $\widetilde{B}_{1}^{\prime}$ always has an $(M-N)$-multiple eigenvalue $\tilde{\lambda}=-\tau$ for $M>N$, for $\sigma_{1}=\sigma_{\min}$, (3.18) and (3.36) indicate that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
$\{-\tau\}\cup[-(\sigma_{\max}+\tau)-\delta,-(\sigma_{\min,2}+\tau)+\delta]\cup[\sigma_{\min,2}-\tau-\delta,\sigma_{\max}-\tau+\delta].$
Notice that $\sigma_{\min,2}=\sigma_{2}$ and (3.32) mean that $\sigma_{\min,2}-\tau>\delta$, and observe that the above positive and negative intervals have the same lengths $\sigma_{\max}-\sigma_{\min,2}+2\delta$. Following the proof of item (ii) of Theorem 3.3, we obtain bound (3.34).
(iii) For $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$, since the eigenvalues of $Z^{T}B_{1}^{\prime}Z$ lie in the interval of (3.20), bound (3.36) means that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
(3.38) $[-(\sigma_{\max}+\tau)-\delta,-|\sigma_{2}-\tau|+\delta]\cup[|\sigma_{2}-\tau|-\delta,\sigma_{\max}-\tau+\delta]$
$\subset[-(\sigma_{\max}+\tau)-\delta,-|\sigma_{2}-\tau|+\delta]\cup[|\sigma_{2}-\tau|-\delta,\sigma_{\max}+\tau+\delta]$ ,
Since the two intervals in (3.38) have the equal length $\sigma_{\max}+\tau-|\sigma_{2}-\tau|+2\delta$, applying the proof of item (ii) of Theorem 3.2 to this interval union establishes (3.35). ∎
Remark 3.
It can be verified that the factors in (3.33)–(3.35) are approximations to $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ in (3.21) with the relative errors $\mathcal{O}(\frac{\delta}{\tau-\sigma_{\max,2}})$, $\mathcal{O}(\frac{\delta}{\sigma_{\min,2}-\tau})$ or $\mathcal{O}(\frac{\delta}{|\sigma_{2}-\tau|})$, respectively. Therefore, for $\frac{\delta}{\tau-\sigma_{\max,2}}$, $\frac{\delta}{\sigma_{\min,2}-\tau}$ or $\frac{\delta}{|\sigma_{2}-\tau|}$ fairly small, Remarks 12 on Theorem 3.3 apply to Theorem 3.5.
4 A preconditioned correction equation, MINRES, and practical issues
Recall (3.16)–(3.18), and write $\widetilde{W}_{1}=W_{1}$ in (3.1) for $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$. Then the normalized eigenvectors of $\widetilde{B}_{1}=B_{1}=W_{1}B_{1}^{\prime}W_{1}^{T}$ defined in (3.3) associated with the eigenvalues ±$\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau$ and $-\tau$ are $\frac{1}{\sqrt{2}}[u_{i}^{T},\pm v_{i}^{T}]^{T},\ i=2,\ldots,N$ and $[u^{T}_{i},\mathbf{0}]^{T}$, $i=N+1,N+2,\ldots,M$, respectively; see (1.1). Therefore, those smallest eigenvalues in magnitude of $B_{1}$ correspond to the singular values of $A$ closest to $\tau$. Notice that for $\tilde{u}_{1}\not=u_{1}$ and $\tilde{v}_{1}\not=v_{1}$, since $\widetilde{B}_{1}\rightarrow B_{1}$ as $\tilde{u}_{1}\rightarrow u_{1}$ and $\tilde{v}_{1}\rightarrow v_{1}$, the eigenvalues and eigenvectors (rigorously speaking, the eigenspaces associated with each of the well isolated clusters of eigenvalues) of $\widetilde{B}_{1}$ converge to those of $B_{1}$ correspondingly. As a result, for $\tilde{u}_{1}\rightarrow u_{1}$ and $\tilde{v}_{1}\rightarrow v_{1}$, if $B_{1}$ has clustered smallest eigenvalues in magnitude well separated from the others, then $\widetilde{B}_{1}$ has the same number of clustered smallest eigenvalues in magnitude well isolated with the others.
The preceding convergence results show that MINRES may converge very slowly when $\sigma_{1}$ is clustered with some other singular values of $A$, which corresponds to the case that the clustered smallest eigenvalues in magnitude of $\widetilde{B}_{1}$ are well separated from the others. In this case, it is vital to precondition the correction equation (2.7) in order to accelerate MINRES. Theoretically, this can be realized by annihilating the effects of those smallest eigenvalues in magnitude of $\widetilde{B}_{1}$. Strikingly, as JDSVD proceeds, the searching subspaces $\mathcal{U}$ and $\mathcal{V}$ generally contain increasingly more accurate approximations to the singular vectors of $A$ associated with several singular values closest to $\tau$. Based on the last paragraph, this remarkable property indicates that JDSVD provides rich information on the eigenvectors of $B_{1}$ and thus those of $\widetilde{B}_{1}$ associated with those smallest eigenvalues in magnitude. It, therefore, makes us possible to fully exploit such information to precondition the correction equation effectively so as to solve it by MINRES more efficiently, which leads to our inner preconditioned JDSVD (IPJDSVD) method. We will show that such preconditioning is equivalent to forming a preconditioned correction equation, that makes full use of all the approximate left and right singular vectors associated with the clustered singular values nearest to $\tau$. If the clustered singular values closest to $\tau$ are well isolated with the other singular values of $A$, those smallest eigenvalues in magnitude of the preconditioned coefficient matrix are well separated from the other eigenvalues, so that the convergence of MINRES is improved substantially.
Section 4.1 is devoted to a detailed derivation of the preconditioned correction equation; section 4.2 presents convergence results on MINRES for solving it; section 4.3 settles down the practical issue of how to adaptively form the preconditioned correction equation at each outer iteration.
Suppose that there are $m$ clustered singular values of $A$ closest to $\tau$:
(4.1) $|\sigma_{1}-\tau|\lesssim\dots\lesssim|\sigma_{m}-\tau|\ll|\sigma_{m+1}-\tau|\leq\dots\leq|\sigma_{N}-\tau|.$
From (1.1), we denote by
[…]
Such $(\Sigma_{m},U_{m},V_{m})$ is a partial SVD of $A$ with the $m$ singular values closest to $\tau$.
At the $k$th outer iteration of JDSVD, suppose that $k>m$. Recall from section 2 that $(\theta_{i},c_{i},d_{i})$, $i=1,\dots,k$ are the singular triplets of $H$ in (2.2), labeled as $|\theta_{1}-\tau|\leq|\theta_{2}-\tau|\leq\dots\leq|\theta_{k}-\tau|$. We compute the $k$ Ritz approximations
(4.3) $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})=(\theta_{i},\widetilde{U}c_{i},\widetilde{V}d_{i}),\qquad i=1,\dots,k,$
and take $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ to approximate the desired $(\sigma_{1},u_{1},v_{1})$. Denote
[…]
Then $(\Theta_{m},\widetilde{U}_{m},\widetilde{V}_{m})$ is an approximation to $(\Sigma_{m},U_{m},V_{m})$.
4.1 A preconditioned correction equation
By (2.1) and (2.4), from (4.4) we have
[…]
(4.6) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r,$
where we have exploited the fact that $(s,t)\perp\perp(\tilde{u}_{1},\tilde{v}_{1})$ and the vectors $\tilde{u}_{1}$ and $\tilde{v}_{1}$ are the first columns of $\widetilde{U}_{m}$ and $\widetilde{V}_{m}$, respectively. Writing
$\begin{bmatrix}s\\ t\end{bmatrix}=\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}+\begin{bmatrix}\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}$
and substituting this relation into (4.6), we obtain
(4.7) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}\!\!\!\!&\\ &\!\!\!\!I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}\!\!\!\!&\\ &\!\!\!\!I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r-r_{\mathrm{tail}}$
with
(4.8) $r_{\mathrm{tail}}$ $=$ $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}$
$=$ $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}&\\ &\widetilde{V}_{m}\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}^{T}s\\ \widetilde{V}_{m}^{T}t\end{bmatrix}$
$=$ $\begin{bmatrix}&A\widetilde{V}_{m}-\widetilde{U}_{m}\Theta_{m}\\ A^{T}\widetilde{U}_{m}-\widetilde{V}_{m}\Theta_{m}\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}^{T}s\\ \widetilde{V}_{m}^{T}t\end{bmatrix},$
where the last equality comes from $\Theta_{m}=\widetilde{U}_{m}^{T}A\widetilde{V}_{m}$.
From (2.7), the first elements of $\widetilde{U}_{m}^{T}s$ and $\widetilde{V}_{m}^{T}t$ are $\tilde{u}_{1}^{T}s=0$ and $\tilde{v}_{1}^{T}t=0$. Denote $\widetilde{U}_{m}^{\prime}=[\tilde{u}_{2},\dots,\tilde{u}_{m}]$, $\widetilde{V}_{m}^{\prime}=[\tilde{v}_{2},\dots,\tilde{v}_{m}]$ and $\Theta_{m}^{\prime}=[\theta_{2},\dots,\theta_{m}]$. Then (4.8) becomes
(4.9) $r_{\mathrm{tail}}=\begin{bmatrix}&\!\!\!\!A\widetilde{V}_{m}^{\prime}-\widetilde{U}_{m}^{\prime}\Theta_{m}^{\prime}\\ A^{T}\widetilde{U}_{m}^{\prime}-\widetilde{V}_{m}^{\prime}\Theta_{m}^{\prime}&\end{bmatrix}\begin{bmatrix}(\widetilde{U}_{m}^{\prime})^{T}s\\ (\widetilde{V}_{m}^{\prime})^{T}t\end{bmatrix}=\begin{bmatrix}&\!\!\!\!R_{1}\\ R_{2}&\end{bmatrix}\begin{bmatrix}(\widetilde{U}_{m}^{\prime})^{T}s\\ (\widetilde{V}_{m}^{\prime})^{T}t\end{bmatrix},$
where $R_{1}=A\widetilde{V}_{m}^{\prime}-\widetilde{U}_{m}^{\prime}\Theta_{m}^{\prime}$ and $R_{2}=A^{T}\widetilde{U}_{m}^{\prime}-\widetilde{V}_{m}^{\prime}\Theta_{m}^{\prime}$, whose $(i-1)$-th columns are respectively the upper and lower parts of the residual $r^{(i)}=r(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ of $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ defined by (2.4), $i=2,\dots,m$. Keep in mind the three properties: (i) the sizes of $\|s\|$ and $\|t\|$ are of the same orders as the errors of $\tilde{u}_{i}$ and $\tilde{v}_{i}$, respectively; (ii) $\|r^{(i)}\|$ is of the same order as the errors of $\tilde{u}_{i},\tilde{v}_{i}$, $i=1,\dots,m$; (iii) $(\widetilde{U}_{m}^{\prime})^{T}s$ and $(\widetilde{V}_{m}^{\prime})^{T}t$ are no more than $s$ and $t$ in size. Therefore, whenever $\|r^{(i)}\|,i=2,\ldots,m$ are comparable to $\|r\|$, that is, $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i}),i=2,\dots,m$ have comparable accuracy to $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$, one has $\|r_{\rm tail}\|=\mathcal{O}(\|r\|^{2})$ and thus $r+r_{\rm tail}\approx r$ in (4.7) once $\|r\|$ becomes fairly small.
Omitting $r_{\mathrm{tail}}$ in (4.7), we have now derived a preconditioned correction equation
(4.10) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r,$
whose solution $(s,t)\perp\perp(\widetilde{U}_{m},\widetilde{V}_{m})$ has the same quality as the solution $(s,t)$ to (2.7) provided that $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=2,\dots,m$ have comparable accuracy to $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$.
With an approximate solution $(\tilde{s}^{T},\tilde{t}^{T})^{T}$ of (4.10) available, we orthonormalize $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ to obtain the expansion vectors $u_{+}$ and $v_{+}$, and expand the subspaces in a regular way until $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ converges. This forms our basic IPJDSVD. Importantly, the above derivation and analysis indicate that the solution to the preconditioned correction equation expands the searching subspaces as effectively as the solution to the standard correction equation (2.7) does. In other words, IPJDSVD does not spoil the subsequent outer iterations of JDSVD.
4.2 Convergence analysis on MINRES for solving (4.10)
Let $\widetilde{P}_{m}$ and $\widetilde{Q}_{m}$ make $[\widetilde{U}_{m},\widetilde{P}_{m}]$ and $[\widetilde{V}_{m},\widetilde{Q}_{m}]$ orthogonal, and define the column orthonormal matrix
(4.11) $\widetilde{W}_{m}=\begin{bmatrix}\widetilde{P}_{m}&\\ &\widetilde{Q}_{m}\end{bmatrix}\in\mathbb{R}^{(M+N)\times(M+N-2m)}.$
Then the coefficient matrix in (4.10) can be written as
(4.12) $\widetilde{B}_{m}=\widetilde{W}_{m}\widetilde{B}_{m}^{\prime}\widetilde{W}_{m}^{T}\quad\mbox{with}\quad\widetilde{B}_{m}^{\prime}=\begin{bmatrix}-\tau I_{M-m}&\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}\\ \widetilde{Q}_{m}^{T}A^{T}\widetilde{P}_{m}&-\tau I_{N-m}\end{bmatrix}.$
Denote by $\tilde{w}_{j}$ the $j$th approximate solution computed by MINRES for (4.10) with $\tilde{w}_{0}=\bm{0}$, $j=1,2,\ldots,M+N-2m$. Then the corresponding inner residuals are
(4.13) $\qquad r_{\mathrm{in},j}=\widetilde{B}_{m}\tilde{w}_{j}+r,\qquad\qquad j=0,1,\ldots,M+N-2m.$
Using a similar proof to that of Theorem 3.1, we can obtain the following results.
Theorem 4.1.
With $\widetilde{B}_{m}^{\prime}$ defined in (4.12), the solution to the preconditioned correction equation (4.10) is $w=\widetilde{W}_{m}z$ where $z=\widetilde{W}_{m}^{T}w\in\mathbb{R}^{M+N-2m}$ solves
(4.14) $\widetilde{B}_{m}^{\prime}z=-r^{\prime}\qquad\mbox{with}\qquad r^{\prime}=\widetilde{W}_{m}^{T}r.$
If MINRES is applied to (4.14) with the initial guess $\tilde{z}_{0}=\bm{0}$, then the $j$th approximate solution $\tilde{z}_{j}=\widetilde{W}_{m}^{T}\tilde{w}_{j}$ with $\tilde{w}_{j}=\widetilde{W}_{m}\tilde{z}_{j}$ and
(4.15) $\|r_{\mathrm{in},j}\|=\|r^{\prime}_{\mathrm{in},j}\|=\|\widetilde{B}_{m}^{\prime}\tilde{z}_{j}+r^{\prime}\|,\ j=1,2,\ldots,M+N-2m.$
Theorem 4.1 establishes the equivalence of MINRES for solving (4.10) and (4.14). Notice that $r=\mathbf{0}$ when $\widetilde{U}_{m}=U_{m}$ and $\widetilde{V}_{m}=V_{m}$. In this case, we can extend Theorem 3.3 to MINRES for solving (4.10) by supposing any virtual right hand side $r\not=\mathbf{0}$ that satisfies (4.5).
Theorem 4.2.
Suppose that $\mathcal{R}(\widetilde{U}_{m})=\mathcal{R}(U_{m})$ and $\mathcal{R}(\widetilde{V}_{m})=\mathcal{R}(V_{m})$ in (4.10) with $U_{m}$ and $V_{m}$ in (4.2) and $\mathcal{R}(\cdot)$ the column space of a matrix. Then (i) if $\tau>(\sigma_{\max}+\sigma_{\max,m+1})/2$ with $\sigma_{\max,m+1}$ the $(m+1)$th largest singular value of $A$, then
(4.16) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,m+1})/(\tau-\sigma_{\max,m+1})}}\right)^{j};$
(ii) if $\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$ with $\sigma_{\min,m+1}$ the $(m+1)$th smallest singular value of $A$ and writing $j_{o}=\min\{1,M-N\}$, then
(4.17) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\left(1-\frac{2}{1+\sqrt{(\sigma_{\max}^{2}-\tau^{2})/(\sigma_{\min,m+1}^{2}-\tau^{2})}}\right)^{\left[\frac{j-j_{o}}{2}\right]};$
(iii) if $(\sigma_{\min}+\sigma_{\min,m+1})/2<\tau<(\sigma_{\max}+\sigma_{\max,m+1})/2$, then
(4.18) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau)/|\sigma_{m+1}-\tau|}\right)^{\left[\frac{j}{2}\right]}$
with $\sigma_{m+1}$ the $(m+1)$th closest singular value to $\tau$ in terms of the labeling rule (1.2).
Proof.
Under the hypothesis, we take $\widetilde{P}_{m}=P_{m}$ and $\widetilde{Q}_{m}=Q_{m}$ in (4.11) with
(4.19) $P_{m}=[u_{m+1},\dots,u_{M}]\qquad\mbox{and}\qquad Q_{m}=[v_{m+1},\dots,v_{N}].$
In this case, the matrix $\widetilde{B}_{m}^{\prime}$ defined in (4.12) is simplified as
(4.20) $B_{m}^{\prime}=\begin{bmatrix}-\tau I_{M-m}&P_{m}^{T}AQ_{m}\\ Q_{m}^{T}A^{T}P_{m}&-\tau I_{N-m}\end{bmatrix}=\begin{bmatrix}-\tau I_{M-m}&\Sigma_{m}^{\prime}\\ (\Sigma_{m}^{\prime})^{T}&-\tau I_{N-m}\end{bmatrix},$
where
[…]
Therefore, the eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ are
(4.22) ±$\pm\sigma_{m+1}-\tau,\dots,\pm\sigma_{N}-\tau\qquad\mbox{and}\qquad\underbrace{-\tau,\dots,-\tau}_{M-N}.$
(i) If $\tau>(\sigma_{\max}+\sigma_{\max,m+1})/2$, then from (4.1) the $m$ singular values $\sigma_{1}$, $\dots$, $\sigma_{m}$ closest to $\tau$ are the $m$ largest ones $\sigma_{\max}$, $\sigma_{\max,2}$, $\dots$, $\sigma_{\max,m}$ of $A$ and the remaining singular values $\sigma_{m+1},\ldots,\sigma_{N}$ are smaller than $\tau$, among which $\sigma_{m+1}=\sigma_{\max,m+1}$. Therefore, the eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ in (4.22) are negative and lie in
$[-(\sigma_{\max,m+1}+\tau),-(\tau-\sigma_{\max,m+1})].$
Applying (3.11) to MINRES for (4.10) leads to (4.16).
(ii) If $\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$, then the $m$ desired singular values are the $m$ smallest ones $\sigma_{\min}$, $\sigma_{\min,2}$, $\dots$, $\sigma_{\min,m}$ of $A$. In this case, $\sigma_{m+1}=\sigma_{\min,m+1}$. The eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ in (4.22) lie in
$\{-\tau\}\cup[-(\sigma_{\max}+\tau),-(\sigma_{\min,m+1}+\tau)]\cup[\sigma_{\min,m+1}-\tau,\sigma_{\max}-\tau],$
where the negative and positive intervals have the equal length $\sigma_{\max}-\sigma_{\min,m+1}$. Using a similar proof approach to that of (3.14), we obtain (4.17).
(iii) For $(\sigma_{\min}+\sigma_{\min,m+1})/2<\tau<(\sigma_{\max}+\sigma_{\max,m+1})/2$, the $\lambda_{i}$ in (4.22) lie in
(4.23) $[-(\sigma_{\max}+\tau),-|\tau-\sigma_{m+1}|]\cup[|\tau-\sigma_{m+1}|,\sigma_{\max}-\tau]$
$\subset[-(\sigma_{\max}+\tau),-|\tau-\sigma_{m+1}|]\cup[|\tau-\sigma_{m+1}|,\sigma_{\max}+\tau]$
Notice that the two intervals in (4.23) have the same length $\sigma_{\max}+\tau-|\tau-\sigma_{m+1}|$. Applying (3.12) to MINRES for (4.10) with this interval union yields (4.18). ∎
Remark 4.
Note that items (i)–(iii) of Theorem 4.2 correspond to $\sigma_{1}=\sigma_{\max},\sigma_{\min}$ and $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$. The convergence rates of MINRES for (4.10) thus depend on
(4.24) $\gamma_{4}=\frac{\tau-\sigma_{\max,m+1}}{\tau+\sigma_{\max,m+1}},\qquad\gamma_{5}=\sqrt{\frac{\sigma_{\min,m+1}^{2}-\tau^{2}}{\sigma_{\max}^{2}-\tau^{2}}},\qquad\gamma_{6}=\frac{|\sigma_{m+1}-\tau|}{\sigma_{\max}+\tau},$
which are considerably bigger than $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ in (3.21) whenever $\tau-\sigma_{\max,m+1}\gg\tau-\sigma_{\max,2}$, $\sigma_{\min,m+1}-\tau\gg\sigma_{\min,2}-\tau$ and $|\sigma_{m+1}-\tau|\gg|\sigma_{2}-\tau|$, that is, the $m$ clustered singular values $\sigma_{1},\ldots,\sigma_{m}$ are well separated from the others. As a result, MINRES for (4.10) will converge much faster than it does for (2.7) correspondingly.
In practical computations, $\mathcal{R}(\widetilde{U}_{m})\neq\mathcal{R}(U_{m})$ and $\mathcal{R}(\widetilde{V}_{m})\neq\mathcal{R}(V_{m})$. Denote by
(4.25) $\Phi=\angle\left(\mathcal{R}(\widetilde{U}_{m}),\mathcal{R}(U_{m})\right)\quad\mbox{and}\quad\Psi=\angle\left(\mathcal{R}(\widetilde{V}_{m}),\mathcal{R}(V_{m})\right)$
the principal or canonical angle matrices between $\mathcal{R}(\widetilde{U}_{m})$ and $\mathcal{R}(U_{m})$ and between $\mathcal{R}(\widetilde{V}_{m})$ and $\mathcal{R}(V_{m})$, respectively (cf. [4, p. 329]). For $\widetilde{B}_{m}^{\prime}$ in (4.12) and $B_{m}^{\prime}$ in (4.20), we have the following extension of Theorem 3.4.
Theorem 4.3.
$\widetilde{B}_{m}^{\prime}$ is orthogonally similar to $B_{m}^{\prime}$ with the error
(4.26) $\|\widetilde{B}_{m}^{\prime}-Z^{T}B_{m}^{\prime}Z\|\leq\|A\|\left(\|\sin\Phi\|+\|\sin\Psi\|\right)^{2},$
where $Z$ is some orthogonal matrix in $\mathbb{R}^{(M+N-2m)\times(M+N-2m)}$; see (4.33).
Proof.
By (4.19), let us decompose the orthonormal $\widetilde{P}_{m}$ and $\widetilde{Q}_{m}$, defined in the beginning of section 4.2, as the orthogonal sums
(4.27) $\widetilde{P}_{m}=P_{m}E_{1}+U_{m}E_{2}\qquad\mbox{and}\qquad\widetilde{Q}_{m}=Q_{m}F_{1}+V_{m}F_{2}$
with $E_{1}\in\mathbb{R}^{(M-m)\times(M-m)}$, $E_{2}\in\mathbb{R}^{m\times(M-m)}$, $F_{1}\in\mathbb{R}^{(N-m)\times(N-m)}$, $F_{2}\in\mathbb{R}^{m\times(N-m)}$. Therefore,
(4.28) $\|\sin\Phi\|=\|\widetilde{P}_{m}^{T}U_{m}\|=\|E_{2}\|\qquad\mbox{and}\qquad\|\sin\Psi\|=\|\widetilde{Q}_{m}^{T}V_{m}\|=\|F_{2}\|.$
Let the SVDs of $E_{1}$ and $F_{1}$ be
(4.29) $E_{1}=X_{1}\Sigma_{E}X_{2}^{T}\qquad\mbox{and}\qquad F_{1}=Y_{1}\Sigma_{F}Y_{2}^{T}.$
Using the same approach to proving (3.26), we obtain
(4.30) $\|I-\Sigma_{E}\|\leq\|\sin\Phi\|^{2}\qquad\mbox{and}\qquad\|I-\Sigma_{F}\|\leq\|\sin\Psi\|^{2}.$
By the definitions of $P_{m},Q_{m}$ and $U_{m},V_{m}$, we have $U_{m}^{T}AV_{m}=\Sigma_{m}$, $P_{m}^{T}AQ_{m}=\Sigma_{m}^{\prime}$ and $P_{m}^{T}AV_{m}=\bm{0}$, $U_{m}^{T}AQ_{m}=\bm{0}$, where $\Sigma_{m}$ and $\Sigma_{m}^{\prime}$ are defined in (4.2) and (4.21). Therefore, exploiting (4.27) and (4.29), we have
(4.31) $\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}=E_{1}^{T}\Sigma_{m}^{\prime}F_{1}+E_{2}\Sigma_{m}F_{2}=X_{2}\Sigma_{E}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}Y_{2}^{T}+E_{2}\Sigma_{m}F_{2}.$
Note that $X_{i},Y_{i},i=1,2$ are orthogonal and $\|\Sigma_{m}^{\prime}\|\leq\|A\|$, $\|\Sigma_{m}\|\leq\|A\|$. Exploiting (4.28), (4.30) and (4.31), we obtain
$\|E_{2}\Sigma_{m}F_{2}\|\leq\|A\|\|\sin\Phi\|\|\sin\Psi\|$
and
(4.32) $\|\!\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}\!-\!X_{\!2}X_{1}^{T}\Sigma_{m}^{\prime}Y_{\!1}Y_{\!2}^{T}\|$ $=$ $\|\Sigma_{E}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}-X_{1}^{T}\!\Sigma_{m}^{\prime}Y_{\!1}+X_{2}^{T}E_{2}\Sigma_{m}F_{2}Y_{2}\|$
$\leq$ $\|\!(\Sigma_{E}\!-\!I)X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}\!+\!X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}(\Sigma_{F}\!-\!I)\!\|\!+\!\|\!E_{2}\Sigma_{m}F_{2}\!\|$
$\leq$ $\|\Sigma_{E}-I\|\|\Sigma_{m}^{\prime}\|+\|\Sigma_{m}^{\prime}\|\|\Sigma_{F}-I\|+\|E_{2}\|\|\Sigma_{m}\|\|F_{2}\|$
$\leq$ $\|A\|(\|\sin\Phi\|^{2}+\|\sin\Psi\|^{2}+\|\sin\Phi\|\|\sin\Psi\|)$
$\leq$ $\|A\|(\|\sin\Phi\|+\|\sin\Psi\|)^{2}.$
Denote the orthogonal matrix
(4.33) $Z=\begin{bmatrix}X_{1}X_{2}^{T}&\\ &Y_{1}Y_{2}^{T}\end{bmatrix}\in\mathbb{R}^{(M+N-2m)\times(M+N-2m)}.$
By the definitions of $\widetilde{B}_{m}^{\prime}$ in (4.12) and $B_{m}^{\prime}$ in (4.20), we obtain
$\|\widetilde{B}_{m}^{\prime}-Z^{T}B_{m}^{\prime}Z\|=\|\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}-X_{2}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}Y_{2}^{T}\|.$
Then (4.26) follows from applying (4.32) to the above. ∎
Making use of Theorems 4.24.3, we can establish the following convergence results on MINRES for solving (4.10). The proof is analogous to that of Theorem 3.5.
Theorem 4.4.
With the notations of Theorem 4.3, assume that
(4.34) $\delta:=\|A\|(\|\sin\Phi\|+\|\sin\Psi\|)^{2}<|\sigma_{m+1}-\tau|.$
(i) If $\tau>(\sigma_{\max}+\sigma_{\max,m+1})/2$, then
(4.35) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,m+1}+\delta)/(\tau-\sigma_{\max,m+1}-\delta)}}\right)^{j};$
(ii) If $\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$ and writing $j_{o}=\min\{1,M-N\}$, then
(4.36) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}+\delta}{\tau}\right)^{j_{o}}\left(1-\frac{2}{1+\sqrt{\frac{(\sigma_{\max}+\delta)^{2}-\tau^{2}}{(\sigma_{\min,m+1}-\delta)^{2}-\tau^{2}}}}\right)^{\left[\frac{j-j_{o}}{2}\right]}.$
(iii) If $(\sigma_{\min}+\sigma_{\min,m+1})/2<\tau<(\sigma_{\max}+\sigma_{\max,m+1})/2$, then
(4.37) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau+\delta)/(|\sigma_{m+1}-\tau|-\delta)}\right)^{\left[\frac{j}{2}\right]}.$
The factors in (4.35)–(4.37) are approximations to $\gamma_{4},\gamma_{5}$ and $\gamma_{6}$ in (4.24) with the relative errors $\mathcal{O}(\frac{\delta}{\tau-\sigma_{\max,m+1}})$, $\mathcal{O}(\frac{\delta}{\sigma_{\min,m+1}-\tau})$ and $\mathcal{O}(\frac{\delta}{|\sigma_{m+1}-\tau|})$, respectively. This indicates that the convergence behavior of MINRES for (4.10) resembles that in the ideal case as $(\Theta_{m},\widetilde{U}_{m},\widetilde{V}_{m})$ is converging. If $\sigma_{1}$ is close to $\sigma_{2}$ but well separated from $\sigma_{m+1}$, MINRES for the preconditioned correction equation (4.10) can converge much faster than it does for the standard correction equation (2.7).
4.3 Selection strategy of approximate singular triplets in inner preconditioning
The previous results and analysis have shown that it is much better to use MINRES to solve the preconditioned (4.10) rather than (2.7) when there are totally $m$ singular values of $A$ clustered at $\tau$ and their approximate singular vectors already have some accuracy. However, the number $m$ is unknown in practice. We need to set up necessary criteria so as to propose a practical IPJDSVD algorithm.
Quantitatively, we will claim a singular value $\sigma_{i}$ clustered at $\tau$ if
(4.38) $\frac{|\sigma_{i}-\tau|}{\max\{\sigma_{i},1\}}\leq\textit{pretol}_{1}$
for a sufficiently small $\textit{pretol}_{1}$. Theorems 4.34.4 indicate that if the singular vectors with the $m$ clustered singular values are computed with some accuracy, say $\|\sin\Phi\|+\|\sin\Psi\|\leq\textit{pretol}_{2}$ for an appropriately small $\textit{pretol}_{2}$, then we benefit very much from MINRES for solving the preconditioned (4.10).
During each outer iteration, for the currently computed approximate singular triplet $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=2,\dots,k$, if $\theta_{i}$ and the residual $r^{(i)}=r(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ in (2.4) satisfy
(4.39) $|\theta_{i}-\tau|\leq\max\{\theta_{i},1\}\cdot\textit{pretol}_{1}\qquad\mbox{and}\qquad\|r^{(i)}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{pretol}_{2},$
then we accept $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ as a reasonably good approximate singular triplet of $A$ and claim $\theta_{i}$ clustered at $\tau$, where $\|A\|_{\mathrm{e}}=\sqrt{\|A\|_{1}\cdot\|A\|_{\infty}}$ is a rough estimate of $\|A\|=\sigma_{\max}$ with $\|A\|_{1}$ and $\|A\|_{\infty}$ the $1$- and $\infty$-norms of $A$, respectively. If $A$ is accessible, we calculate $\|A\|_{\mathrm{e}}$ directly; if $A$ is implicit and only products of vectors with $A$ and $A^{T}$ can be computed, we compute $\|A\|_{1}=\max\limits_{1\leq i\leq N}\{\|Ae_{i,N}\|_{1}\}$ and $\|A\|_{\infty}=\max\limits_{1\leq i\leq M}\{\|A^{T}e_{i,M}\|_{1}\}$ with $e_{i,j}$ and the $i$th column of the identity matrix $I_{j}$, $j=N,M$. We take the $\tilde{u}_{i},\tilde{v}_{i}$ to form
(4.40) $\widetilde{U}_{\tilde{m}}=[\tilde{u}_{1},\tilde{u}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]\qquad\mbox{and}\qquad\widetilde{V}_{\tilde{m}}=[\tilde{v}_{1},\tilde{v}_{i_{2}},\dots,\tilde{v}_{i_{\tilde{m}}}],$
where the indices $i_{2},\dots,i_{\tilde{m}}\in\{2,\dots,k\}$ are such that (4.39) is fulfilled, and replace $\widetilde{U}_{m}$ and $\widetilde{V}_{m}$ by them in (4.10).
For $k$ small, there may be no sufficiently good approximate singular triplet that satisfies (4.39). In this case, $\tilde{m}=1$ in (4.40), and (4.10) degenerates to (2.7). However, for $m>1$, as the dimension $k$ increases, more and more computed approximate singular triplets start to converge and fulfill (4.39), so that $\tilde{m}>1$ pairs of approximate singular vectors participate (4.10). Then the inner preconditioning starts to show its strength. In computations, $\tilde{m}$ in (4.40) is determined dynamically according to (4.39). $\tilde{m}$ increases from 1 to $m$, and we have $\tilde{m}=m$ ultimately for $k$ sufficiently large.
5 A thick-restart IPJDSVD algorithm with deflation and purgation
For practical purpose, we must limit the subspace dimension $k\leq k_{\max}$. If JDSVD and IPJDSVD do not yet converge until $k=k_{\max}$, we must restart them. Furthermore, if we are required to compute $\ell>1$ singular triplets of $A$ corresponding to the $\ell$ singular values $\sigma_{1},\ldots,\sigma_{\ell}$ closest to the target $\tau$ (cf. (1.2)), an appropriate deflation is necessary. In addition, a certain purgation is crucial and may greatly reduce the outer iterations for computing the second to the $\ell$th singular triplets.
5.1 A new thick-restart
If IPJDSVD does not yet converge for computing the desired $\ell$ singular triplets for a given $k_{\max}$, we modify a commonly used thick restart, initially proposed in [25] and later popularized in the JDSVD and JDGSVD type methods [10, 11, 12, 29]. In what follows we propose a more effective thick-restart IPJDSVD than the standard thick-restart JDSVD. Our thick-restart retains some $k_{\mathrm{new}}$-dimensional restarting subspaces with
(5.1) $k_{\mathrm{new}}=\max\{k_{\min},\tilde{m}\}$
where $k_{\min}>1$ is a user-prescribed minimal dimension of the searching subspaces and $\tilde{m}$ is the number of approximate singular triplets of $A$ that are selected to participate in the preconditioned correction equation (4.10) in the current expansion phase.
Specifically, assume that $\tilde{m}$ approximate singular triplets $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i}),i=1,i_{2},\dots$, $i_{\tilde{m}}$ of $A$ with $i_{2},\dots,i_{\tilde{m}}\in\{2,\dots,k\}$ join the preconditioned (4.10) before restart. If $\tilde{m}\leq k_{\min}$, we simply restore the standard thick-restart scheme, which forms the orthonormal basis matrices of the new $k_{\min}$ dimensional $\mathcal{U}$ and $\mathcal{V}$ by
(5.2) $\widetilde{U}=[\tilde{u}_{1},\tilde{u}_{2},\dots,\tilde{u}_{k_{\min}}],\qquad\widetilde{V}=[\tilde{v}_{1},\tilde{v}_{2},\dots,\tilde{v}_{k_{\min}}],$
and updates the projection matrix of $A$ with respect to these two subspaces by
[…]
If $\tilde{m}>k_{\min}$, we form the orthonormal basis matrices of the new $\tilde{m}$ dimensional $\mathcal{U}$ and $\mathcal{V}$ and the corresponding projection matrix of $A$ by
[…]
with $\widetilde{U}_{\tilde{m}}$ and $\widetilde{V}_{\tilde{m}}$ defined by (4.40). In this way, the $k_{\rm new}$-dimensional restarting subspaces contain approximate left and right singular subspaces associated with all available approximate singular values clustered at $\tau$. Whenever $m<k_{\max}$ considerably, the new thick-restart IPJDSVD is expected to substantially outperform the standard thick-restart JDSVD in terms of the outer iterations.
5.2 Deflation and purgation
Suppose that we are required to compute the $\ell>1$ singular triplets of $A$ corresponding to the $\ell$ singular values closest to the target $\tau$. Then IPJDSVD combined with the deflation advocated in [10] meet this demand perfectly. Concretely, assume that $j<\ell$ approximate singular triplets $\left(\theta_{i,\mathrm{c}},u_{i,\mathrm{c}},v_{i,\mathrm{c}}\right)$ have already converged to $(\sigma_{i},u_{i},v_{i})$, $i=1,\dots,j$ with the residual $r^{(i,\mathrm{c})}=r(\theta_{i,\mathrm{c}},u_{i,\mathrm{c}},v_{i,\mathrm{c}})$ defined by (2.4) satisfying
(5.5) $\|r^{(i,\mathrm{c})}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{tol},\qquad i=1,\dots,j$
for a user-prescribed stopping tolerance $\textit{tol}>0$, where $\|A\|_{\rm e}$ is defined as in (4.39). Denote
$\Sigma_{\mathrm{c}}=\mathrm{diag}\{\theta_{1,\mathrm{c}},\dots,\theta_{j,\mathrm{c}}\},\qquad U_{\mathrm{c}}=[u_{1,\mathrm{c}},\dots,u_{j,\mathrm{c}}],\qquad V_{\mathrm{c}}=[v_{1,\mathrm{c}},\dots,v_{j,\mathrm{c}}].$
[…]
If $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ does not converge, we take $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=1,i_{2},\dots,i_{\tilde{m}}$ that meet (4.39) to form $\widetilde{U}_{\tilde{m}}$ and $\widetilde{V}_{\tilde{m}}$ in (4.40), as described in section 4.3. Then we set $U_{\mathrm{p}}=[U_{\mathrm{c}},\widetilde{U}_{\tilde{m}}]$ and $V_{\mathrm{p}}=[V_{\mathrm{c}},\widetilde{V}_{\tilde{m}}]$, and use MINRES to approximately solve the new preconditioned correction equation
(5.6) $\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}\!\!&\\ &\!\!I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}\!\!&\\ &\!\!I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=r_{\mathrm{p}}$
for $(s,t)\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$, where
$r_{\mathrm{p}}:=-\begin{bmatrix}I-U_{\mathrm{c}}U_{\mathrm{c}}^{T}&\\ &I-V_{\mathrm{c}}V_{\mathrm{c}}^{T}\end{bmatrix}r$
with $r=r(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ the residual of $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ defined by (2.4). We compute an approximate solution $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$ and terminate MINRES when the inner residual $r_{\mathrm{in}}$ in (4.13) satisfies
(5.7) $\|r_{\mathrm{in}}\|\leq\|r\|\cdot\min\{{\rho\tilde{\varepsilon}},0.01\},$
where $\tilde{\varepsilon}\in[10^{-4},10^{-3}]$ is a user-prescribed tolerance and $\rho$ is a constant that is estimated by using all the approximate singular values computed at the current cycle; see [10] for details. We use the approximate solution $\tilde{s}$ and $\tilde{t}$ to expand the subspaces $\mathcal{U}$ and $\mathcal{V}$. Since $(\widetilde{U},\widetilde{V})\perp\perp(U_{\mathrm{c}},V_{\mathrm{c}})$ and $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{c}},V_{\mathrm{c}})$, the expanded $\mathcal{U}$ and $\mathcal{V}$ are automatically orthogonal to $\mathcal{R}(U_{\mathrm{c}})$ and $\mathcal{R}(V_{\mathrm{c}})$, respectively. IPJDSVD then proceeds in a regular way until $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ converges to $(\sigma_{j+1},u_{j+1},v_{j+1})$. We denote $\left(\theta_{j+1,\mathrm{c}},u_{j+1,\mathrm{c}},v_{j+1,\mathrm{c}}\right)=(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$, add it to the converged $\left(\Sigma_{\mathrm{c}},U_{\mathrm{c}},V_{\mathrm{c}}\right)$, and set $j=j+1$. Proceed in such a way until all the $\ell$ singular triplets are found.
In the deflated IPJDSVD method described above, we now have $1\leq\tilde{m}\leq m-j$ for $j<m$ since we have already computed $j$ converged singular triplets and the deflated matrix has only $m-j$ clustered singular values closest to $\tau$. For $j\geq m$, the preconditioned correction equations degenerate to the standard ones, and non-restarted IPJDSVD reduces to the standard JDSVD.
For the purgation issue, we may benefit much from making full use of the current $\mathcal{U}$ and $\mathcal{V}$ when starting to compute the next singular triplet $(\sigma_{j+1},u_{j+1},v_{j+1})$, and, compared with the computation of $(\sigma_{1},u_{1},v_{1})$, save outer iterations considerably when computing $(\sigma_{i},u_{i},v_{i}),\ i=2,\ldots,\ell$. The purgation approach [10] can be straightforwardly introduced into the thick-restart IPJDSVD.
Algorithm 1 Thick-restart IPJDSVD with deflation and purgation for the target $\tau$.
1:  Initialization: Set $k=1$, $k_{\mathrm{c}}=0$, $\Sigma_{\mathrm{c}}=[\ ],U_{\mathrm{c}}=[\ ]$, and $V_{\mathrm{c}}=[\ ]$. Let $\widetilde{U}=[\ ]$, $\widetilde{V}=[\ ]$, $H=[\ ]$, and choose normalized starting vectors $u_{+}=u_{0}$ and $v_{+}=v_{0}$.
2:  Calculate $\|A\|_{e}=\sqrt{\|A\|_{1}\cdot\|A\|_{\infty}}$.
3:  while $k\geq 0$ do
4:      Set $\widetilde{U}=[\widetilde{U},u_{+}]$ and $\widetilde{V}=[\widetilde{V},v_{+}]$, and update $H=\widetilde{U}^{T}A\widetilde{V}$.
5:      Compute the singular triplets $(\theta_{i},c_{i},d_{i}),i=1,\dots,k$ of $H$ with $\theta_{i}$ ordered increasingly by their distances from $\tau$. Form the approximate singular triplets $(\theta_{i},\tilde{u}_{i}\!=\widetilde{U}c_{i},\tilde{v}_{i}\!=\widetilde{V}d_{i})$ and the residuals $r^{(i)}\!=\!\begin{bmatrix}\begin{smallmatrix}A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i}\\ A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i}\end{smallmatrix}\end{bmatrix}$, $i\!=\!1,\dots,k$.
6:     if $\|r^{(1)}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{tol}$ then
[…]
8:        if $k_{\mathrm{c}}=\ell$ then return $(\Sigma_{\mathrm{c}},U_{\mathrm{c}},V_{\mathrm{c}})$ and stop. fi
9:         Perform the purgation, set $k=k-1$, and go to step 5.
10:     end if
11:     Determine the $\tilde{m}-1$ indices $i=i_{2},\dots,i_{\tilde{m}}$ according to the condition in (4.39) and form $\widetilde{U}_{\tilde{m}}=[\tilde{u}_{1},\tilde{u}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]$ and $\widetilde{V}_{\tilde{m}}=[\tilde{v}_{1},\tilde{v}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]$.
12:     Set $U_{\mathrm{p}}=[U_{\mathrm{c}},\widetilde{U}_{\tilde{m}}]$, $V_{\mathrm{p}}=[V_{\mathrm{c}},\widetilde{V}_{\tilde{m}}]$ and $r_{\mathrm{p}}=-\begin{bmatrix}\begin{smallmatrix}I-U_{\mathrm{c}}U_{\mathrm{c}}^{T}&\\ &I-V_{\mathrm{c}}V_{\mathrm{c}}^{T}\end{smallmatrix}\end{bmatrix}r^{(1)}$, and solve
$\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}&\\ &I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}&\\ &I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=r_{\mathrm{p}}$
for the approximate solution $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$ with the inner residual $r_{\mathrm{in}}$ satisfying the stopping criterion (5.7).
13:     if $k=k_{\max}$ then perform the thick-restart and reset $k=k_{\mathrm{new}}$ in (5.1). fi
14:      Orthonormalize $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ to obtain the expansion vectors $u_{+}$ and $v_{+}$, respectively, and set $k=k+1$.
15:  end while
5.3 Thick-restart IPJDSVD algorithm with deflation and purgation
Algorithm 1 sketches our thick-restart IPJDSVD with deflation and purgation. It requires the devices to form the matrix-vector multiplications with $A$ and $A^{T}$, the target $\tau$, the number $\ell$ of desired singular triplets, and the outer stopping tolerance tol. It produces a converged partial SVD $(\Sigma_{c},U_{c},V_{c})$ that satisfies
(5.8) $\sqrt{\|AV_{c}-U_{c}\Sigma_{c}\|_{\mathrm{F}}^{2}+\|A^{T}U_{c}-V_{c}\Sigma_{c}\|_{\mathrm{F}}^{2}}\leq\sqrt{\ell}\|A\|_{\mathrm{e}}\cdot\textit{tol},$
where $\|\cdot\|_{\rm F}$ is the F-norm of a matrix. Other optional parameters are normalized starting vectors $u_{0}\in\mathbb{R}^{M}$ and $v_{0}\in\mathbb{R}^{N}$, the maximum and minimum dimensions $k_{\max}$ and $k_{\min}$ of the searching subspaces, the accuracy requirement $\tilde{\varepsilon}$ for expansion vectors in (5.7), and the preconditioning parameters $\textit{pretol}_{1}$ and $\textit{pretol}_{2}$ in (4.39). By defaults, we set $tol=10^{-8}$, $u_{0}=\frac{1}{\sqrt{M}}[1,\dots,1]$, $v_{0}=\frac{1}{\sqrt{N}}[1,\dots,1]$, $k_{\max}=30$, $k_{\min}=3$, $\tilde{\varepsilon}=10^{-4}$, $\textit{pretol}_{1}=0.05$ and $\textit{pretol}_{2}=0.01$.
When going back to step 5 from step 9, the SVD of the updated projection matrix $H$ is known explicitly, so are the approximate singular triplets and corresponding residuals. As a matter of fact, the orthonormal bases of $\mathcal{U}$ and $\mathcal{V}$ after purgation and the corresponding projection matrix of $A$ are
(5.9) $\widetilde{U}=[\tilde{u}_{2},\dots,\tilde{u}_{k}],\qquad\widetilde{V}=[\tilde{v}_{2},\dots,\tilde{v}_{k}],\qquad H=\mathrm{diag}\{\theta_{2},\dots,\theta_{k}\}.$
Therefore, IPJDSVD only needs to reorder all the approximate singular triplets, reassign $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1}):=(\theta_{2},\tilde{u}_{2},\tilde{v}_{2})$, and test its convergence.
6 Numerical experiments
We report numerical experiments to illustrate the considerable superiority of IPJDSVD to JDSVD. All the experiments were performed on a 13th Gen Intel(R) Core(TM) i9-13900KF CPU 3.00 GHz with 64 GB main memory and 24 cores using the Matlab R2022b under the Microsoft Windows 11 Pro 64-bit system with the machine precision $\epsilon=2.22\times 10^{-16}$.
For both algorithms we took the default parameters as described in section 5.3, except for JDSVD, where $\textit{pretol}_{1}=0$ and $\textit{pretol}_{2}=0$. For (2.7), (4.10) and (5.6), we took zero vector as the initial guesses and used the Matlab function minres to solve them until the relative residual norms satisfied (5.7).
In all the tables, we denote by $I_{\mathrm{out}}$ and $I_{\mathrm{in}}$ the total outer and inner iterations used by each algorithm to achieve the convergence, and by $T_{\mathrm{cpu}}$ the total CPU time in seconds counted by the Matlab commands tic and toc. Notice that solving correction equations dominates the overall efficiency of IPJDSVD and JDSVD and their cost at each outer iteration. Here $I_{\mathrm{out}}$ denotes the total number of correction equations solved, which equals the actual outer iterations minus $\ell$ because one does not solve correction equations in the final outer iteration when computing each singular triplet.
Table 1: Properties of the test matrices.
$A$ $M$ $N$ nnz $\sigma_{\max}$ $\sigma_{\min}$
Alemdar 6245 6245 42581 69.5 3.30e-3
bcsstk33 8738 8738 591904 75.8 2.18e-95
cat_ears_3_4 5226 13271 39592 5.49 3.06e-1
epb2 25228 25228 175027 3.34 1.28e-3
relat8 345688 12347 1334038 18.8 0
Rucci1   1977885   109900   7791168 7.07 1.04e-3
dw8192 8192 8192 41746 110 2.88e-5
hep-th 8361 8361 31502 40.5 2.32e-21
ca-HepTh 9877 9877 51971 31.0 2.22e-18
California 9664 9664 16150 21.6 0
inlet 11730 11730 328323 6.57 1.69e-6
garon2 13535 13535 373235 44.9 7.32e-7
e40r0100 17281 17281 553562 13.1 8.68e-8
tomographic1 73159 59498 647495 6.98 0
degme 185501 659415 8127528   2.24e+3 4.13
[…]
Experiment 6.1.
We compute the $10$ singular triplets of $A=\mathrm{Alemdar}$ with the singular values closest to $\tau=67.2$, which are the largest ones and clustered.
Refer to captionRefer to caption
Figure 1: Computing $10$ singular triplets of $A=\mathrm{Alemdar}$ with $\tau=67.2$.
Both IPJDSVD and JDSVD succeeded in computing the $10$ desired singular triplets of $A$ in the correct order. Figure 1 (left) draws the outer iterations used by IPJDSVD and JDSVD to compute each singular triplet, and Figure 1 (right) depicts the average inner iterations per outer iteration of IPJDSVD and JDSVD to compute each desired singular triplet. It is seen from the figure that IPJDSVD used the same outer iterations for the first, second, fourth, sixth and eighth singular triplets as JDSVD, but consumed fewer outer iterations than JDSVD for the fifth, seventh, ninth and last ones, showing the effectiveness of the new thick-restart. Regarding the overall efficiency, IPJDSVD, on average, saved $104$ inner iterations for MINRES to solve the preconditioned correction equation at each outer iteration. IPJDSVD totally consumed $54$ outer iterations, $7433$ inner iterations and $1.83$ seconds, while JDSVD used $59$ outer iterations, $12472$ inner iterations and $2.53$ seconds. IPJDSVD outperforms JDSVD considerably, and reduced $8.47\%$ of the outer iterations, $40.40\%$ of the inner iterations, and $27.76\%$ of the CPU time.
Experiment 6.2.
[…]
Refer to captionRefer to caption
[…]
From Figure 2, we observe that for the first six singular triplets of $A$, IPJDSVD used almost the same outer iterations as JDSVD did; but for the last four desired singular triplets of $A$, IPJDSVD was much faster and used substantially fewer outer iterations, illustrating that the new thick-restart can play a crucial role. More importantly, for each desired singular triplet, the preconditioned MINRES was much faster than the unpreconditioned one.
Precisely, IPJDSVD consumed $73$ outer iterations, $7051$ inner iterations and $3.74$ seconds, while JDSVD used $87$ outer iterations, $12225$ inner iterations and $5.72$ seconds. The former outmatched the latter considerably by reducing $16.09\%$ of the outer iterations, $42.32\%$ of the inner iterations, and $34.61\%$ of the CPU time.
Experiment 6.3.
We compute $10$ singular triplets of four test matrices in Table 1 with given targets $\tau$’s, where the desired singular values are not very interior.
Table 2: Results on Experment 6.3.
$A$ $\tau$ Algorithm: IPJDSVD Algorithm: JDSVD
$I_{\mathrm{out}}$ $I_{\mathrm{in}}$ $T_{\rm cpu}$ $I_{\mathrm{out}}$ $I_{\mathrm{in}}$ $T_{\rm cpu}$
cat_ears_3_4 4.9 74 6627 2.19 84 11219 2.76
epb2 3.3 53 5176 3.68 61 9688 5.09
relat8 16 64 7510 83.2 70 12533 110
Rucci1 6.7 55 5302 331 63 8039 415
Table 2 lists the results. IPJDSVD and JDSVD succeeded in computing the desired singular triplets of all the test matrices. Regarding the outer convergence, we observe from Table 2 that, for all the test matrices, IPJDSVD converged faster than JDSVD, confirming the higher effectiveness of the new thick-restart. Meanwhile, for all the test matrices, IPJDSVD reduced more than $34\%$ of the total inner iterations and more than $20\%$ of the CPU time by JDSVD.
Obviously, for these matrices, IPJDSVD outperforms JDSVD considerably.
Experiment 6.4.
We compute $\ell=10$ very interior and clustered singular triplets of the nine matrices in Table 1 with given $\tau$’s.
Table 3: Results on Experiment 6.4.
$A$ $\tau$ Algorithm: IPJDSVD Algorithm: JDSVD
$I_{\mathrm{out}}$ $I_{\mathrm{in}}$ $T_{\rm cpu}$ $I_{\mathrm{out}}$ $I_{\mathrm{in}}$ $T_{\rm cpu}$
dw8192 109.7 72 4845 2.06 171 11469 3.02
hep-th 11 63 20253 7.34 60 31334 10.3
ca-HepTh 8 65 22019 8.65 61 37117 12.8
California 6 67 30252 9.00 67 46918 12.2
inlet 5.6 75 23565 10.3 85 42909 16.2
garon2 23 79 47190 21.5 91 81969 34.0
e40r0100 13 72 82225 51.1 66 116456 60.8
tomographic1 5.6 57 11227 31.4 63 19629 47.7
degme 28.6 69 18102 565 61 29744 689
[…]
7 Conclusions
We have analyzed the convergence of MINRES for the correction equations in JDSVD. Based on the results obtained, we have proposed an inner preconditioning procedure for the correction equation and derived the preconditioned correction equation, whose solution is proved to have the same accuracy as that to the standard correction equation. We have extended the convergence results on MINRES for the standard correction equation to the preconditoned MINRES, and shown that MINRES for solving the preconditioned correction equation can converge much faster than it does for solving the standard correction equation.
We have presented a practical strategy that dynamically selects approximate singular triplets from current available ones and forms the preconditioned correction equation. Finally, we have developed a new thick-restart IPJDSVD algorithm with deflation and purgation to compute several singular triplets. Numerical experiments have illustrated that the new thick-restart IPJDSVD algorithm outperforms the standard thick-restart JDSVD algorithm considerably, in terms of inner and outer iterations as well as CPU time.
Declarations
The two authors declare that they have no financial interests, and they read and approved the final manuscript.