### v2

Abstract
In a Jacobi–Davidson (JD) type method for singular value decomposition (SVD) problems, called JDSVD, a large symmetric and generally indefinite correction equation is solved iteratively at each outer iteration, which constitutes the inner iterations and dominates the overall efficiency of JDSVD. In this paper, a convergence analysis is made on the minimal residual (MINRES) method for the correction equation. Motivated by the results obtained, at each outer iteration a new correction equation is derived that extracts useful information from current subspaces to construct effective preconditioners for the correction equation and is proven to retain the same convergence of outer iterations of JDSVD. The resulting method is called inner preconditioned JDSVD (IPJDSVD) method; it is also a new JDSVD method, and any viable preconditioner for the correction equations in JDSVD is straightforwardly applicable to those in IPJDSVD. Convergence results show that MINRES for the new correction equation can converge much faster when there is a cluster of singular values closest to a given target. A new thick-restart IPJDSVD algorithm with deflation and purgation is proposed that simultaneously accelerates the outer and inner convergence of the standard thick-restart JDSVD and computes several singular triplets. Numerical experiments justify the theory and illustrate the considerable superiority of IPJDSVD to JDSVD, and demonstrate that a similar two-stage IPJDSVD algorithm substantially outperforms the most advanced PRIMME_SVDS software nowadays for computing the smallest singular triplets.
keywords:
singular value decomposition, singular value, singular vector, Jacobi–Davidson method, outer iteration, inner iteration, correction equation, MINRES, convergence, preconditioning
{AMS}
65F15, 15A18, 65F10
1 Introduction
Let $A\in\mathbb{R}^{M\times N}$ be large and possibly sparse with $M\geq N$, and its singular value decomposition (SVD) be
(1) $A=U\Sigma V^{T},$
where the superscript $T$ denotes the transpose of a matrix, $U=[u_{1},\dots,u_{M}]\in\mathbb{R}^{M\times M}$ and $V=[v_{1},\dots,v_{N}]\in\mathbb{R}^{N\times N}$ are orthogonal, and $\Sigma=\operatorname{diag}\{\sigma_{1},\dots,\sigma_{N}\}\in\mathbb{R}^{M\times N}$ with $\sigma_{i}\geq 0,i=1,\dots,N$. The $(\sigma_{i},u_{i},v_{i})$, $i=1,\dots,N$ are the singular triplets of $A$ with $\sigma_{i}$ the singular values and $u_{i},v_{i}$ the associated left and right singular vectors.
Given a target $\tau\geq 0$ that is not equal to any singular value of $A$, assume that the singular values of $A$ are labeled by
(2) $|\sigma_{1}-\tau|<\dots<|\sigma_{\ell}-\tau|<|\sigma_{\ell+1}-\tau|\leq\dots\leq|\sigma_{N}-\tau|.$
We are interested in the $\ell$ singular triplets $(\sigma_{i},u_{i},v_{i})$, $i=1\dots,\ell$. Specifically, if $\tau$ is inside $(\sigma_{\min},\sigma_{\max})$ but neither close to $\sigma_{\min}$ nor to $\sigma_{\max}$, then $(\sigma_{i},u_{i},v_{i})$, $i=1\dots,\ell$ are called interior singular triplets of $A$; otherwise they are called extreme, i.e., the largest or smallest, singular triplets. Particularly, $\tau=0$ corresponds to the smallest singular triplets and $\tau$ closer to $\sigma_{\max}$ than to the second largest singular value corresponds to the largest singular triplets.
For the concerning large SVD problem, only projection methods can be used. Over the years, projection methods [3, 20, 21, 25, 26] for large eigenvalue problems have been adapted to the computation of a partial SVD of $A$. Given left and right searching subspaces, there are four classes of projection methods for computing approximate singular triplets of $A$, which are based on the standard extraction [1, 5, 6, 7, 13, 28], the harmonic extraction [6, 7, 8, 12, 17, 19], the refined extraction [7, 8, 9, 13, 27, 28], and the refined harmonic extraction [9, 14, 27, 28], respectively. For these methods, a common core task is the construction of left and right searching subspaces so that they contain as much information on the desired singular vectors as possible. Besides a very recent approach that is used in the FEAST SVDsolvers [15, 16] for computing the singular triplets with all the singular values in a given interval, there are mainly two types of approaches for this. The first approach is the Lanczos bidiagonalization (LBD) process [1, 3]. It reduces $A$ to a sequence of small bidiagonal matrices and generates two sets of orthonormal bases of the left and right Krylov subspaces formed by $AA^{T}$ and $A^{T}A$ with relevant starting vectors. Such searching subspaces generally favor the extreme left and right singular vectors, and the distances between these singular vectors and the corresponding Krylov subspaces generally tend to zero fast as the LBD process proceeds. Combined with appropriate extraction approaches, state-of-the-art LBD type algorithms [13, 14, 17, 18, 22] can compute the extreme singular triplets of $A$ efficiently if the desired singular values are not many and are well separated. Unfortunately, the smallest singular values from applications are often clustered, so that the LBD type methods may work poorly.
The second approach is that used in Jacobi–Davidson (JD) type methods proposed by Hochstenbach [6] for the large SVD computation, referred to as JDSVD method. Wu and Stathopoulos [28] present a hybrid two-stage JDSVD method, called PHSVD. At the first stage, PHSVD transforms the SVD problem of $A$ into the mathematically equivalent eigenproblem of the cross-product matrix $A^{T}A$ and computes the desired singular values and right singular vectors to best achievable accuracy by a chosen projection method efficiently. Then if further accuracy is required, it switches to the second stage that solves the eigenproblem of the augmented $\begin{bmatrix}\begin{smallmatrix}\bm{0}&A\\ A^{T}&\bm{0}\end{smallmatrix}\end{bmatrix}$ by a chosen projection method using the approximate singular vectors obtained in the first stage as initial guesses. Wu, Romero and Stathopoulos [27] develop the PRIMME__\__SVDS software programmed in C language; it implements the PHSVD method with the refined extraction in the second stage and uses JDQMR by default, while the method in the first stage is taken as the default ‘DYNAMIC’ that dynamically switches between JDQMR and the GD+k method to minimize the time, where QMR in JDQMR means that a symmetric QMR (sQMR) is used to solve the correction equations [23]. A great merit of the hybrid PRIMME__\__SVDS is that, regarding the outer iterations, JDQMR or GD+k in the first stage is generally much faster than it is in the second stage and, in the meantime, sQMR is generally faster than it is in the second stage, which is true when the smallest singular values are desired. An advantage of sQMR over the MINRES method [4, 21] is that sQMR allows a symmetric indefinite preconditioner and uses the same matrix-vector products at each iteration as MINRES; without preconditioner, sQMR is mathematically equivalent to MINRES [23].
In the sequel, we always assume that PRIMME__\__SVDS uses JDQMR in the second stage. Then at each step of JDSVD [6] and the second stage of PRIMME__\__SVDS, one must solve a certain correction equation that involves $\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ and the current approximate singular triplet, and uses its solution to expand the searching subspaces. Whenever the current approximate singular value starts to converge with fair accuracy, it will replace $\tau$ in the correction equation so that its solution provides richer information on the desired left and right singular vectors. Such kind of correction equations is approximately solved iteratively, called the inner iterations, while the outer iterations compute approximate singular triplets with respect to the searching subspaces. The least solution accuracy requirement on the correction equations and convergence speed of the inner iterations critically affect the outer convergence and dominate the overall efficiency of JDSVD, respectively. The least solution accuracy means that the outer iterations of the resulting inexact JDSVD methods behave as if all the correction equations were solved exactly. The authors in [9] have figured out such accuracy requirement and proven that it generally suffices to solve all the correction equations with the relative errors $\varepsilon\in[10^{-4},10^{-2}]$ of approximate solutions.
As will be seen, the correction equations may be highly indefinite and/or ill conditioned, which is definitely true when the desired singular values are the smallest ones and/or clustered. As a result, it is generally very hard to iteratively solve them using Krylov solvers. To this end, for the overall efficiency of a JDSVD-type method, effectively preconditioning correction equations plays a crucial role. For a general large linear system $Bx=z$, the preconditioning is to transform it into a mathematically equivalent one that is easier to be solved by a suitable Krylov solver, e.g., the conjugate gradient (CG) method, the minimal residual (MINRES) method, the generalized MINRES (GMRES) method, and the biconjugate gradient stabilized (BiCGstab) method [4, 21]. Among others, two major classes of popular algebraic preconditioning techniques are incomplete LU factorization (ILU)-based preconditioners $LU\approx B$ and sparse approximate inverse (SAI) preconditioners $C\approx B^{-1}$ [21].
Since the coefficient matrices in the correction equations involve $\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ and are generally symmetric indefinite, the ILU and SAI preconditioning procedures encounter severe difficulties and are generally not appropriate choices in several aspects. For instance, (i) constructing an SAI preconditioner is costly. In a JDSVD-type method with adaptively changing shifts, the coefficient matrix in the correction equation changes at each outer iteration, which requires that one construct a new SAI preconditioner from scratch. (ii) ILU preconditioners are generally ineffective and have intrinsic numerical instability for indefinite symmetric or highly nonsymmetric systems. (iii) The symmetry of correction equations enables us to fully exploit the merits of MINRES, but ILU and SAI preconditioners destroy the symmetry, so that we have to use a much inferior nonsymmetric solver GMRES or BiCGstab to solve the preconditioned unsymmetric correction equations. Although sQMR allows an indefinite preconditioner and may have convergence behavior similar to MINRES, it indeed has irregular convergence and exhibits plateaus [23, 28], impairing the overall efficiency of JDQMR.
For symmetric positive (or negative) definite and indefinite linear systems, the CG and MINRES are preferences among Krylov iterative solvers, respectively. Setting aside initial vectors, their convergence is uniquely determined by the eigenvalues of an underlying coefficient matrix. The methods converge fast when the eigenvalues are clustered, and they may converge very slowly when the coefficient matrix has relatively small eigenvalues in magnitude [4]. For such a linear system, we seek to preserve the symmetry of the preconditioned linear system and, meanwhile, shrink the eigenvalue inclusion interval(s) of the preconditioned coefficient matrix. Starting with this perspective, for the correction equation in JDSVD, we will see that an ideal preconditioning procedure is to annihilate the $m$ clustered smallest eigenvalues in magnitude of the coefficient matrix that are well separated from the other eigenvalues, so that the eigenvalue distribution of the preconditioned coefficient matrix is favorable and MINRES converges fast. This will turn out to amount to deriving a preconditioned correction equation and forming a new correction equation at each outer iteration, in which the orthogonal projector is formed by the left and right singular vectors $u_{i}$ and $v_{i}$ associated with the $m$ clustered singular values $\sigma_{i},i=1,2,\ldots,m$ rather than only $\sigma_{1}$. Stacking these vectors, we obtain an $m$-dimensional eigenspace of the coefficient matrix in the correction equation associated with its $m$ clustered smallest eigenvalues in magnitude, which will eliminate the effects of these eigenvalues on MINRES and thus accelerate MINRES substantially.
We have observed from extensive numerical experiments that, as the outer iterations of JDSVD proceed, the current subspaces provide increasingly better approximations to the left and right singular vectors $u_{i}$ and $v_{i}$, $i=1,\ldots,m$ corresponding to the $m$ clustered singular values. These approximations form a good approximate eigenspace of the coefficient matrix corresponding to its $m$ clustered smallest eigenvalues in magnitude. We will fully exploit such rich information to derive a new symmetric correction equation, which is equivalent to preconditioning the original one in JDSVD and eliminates the effects of these smallest eigenvalues as much as possible, so that MINRES can converge much faster. The resulting method is called the inner-preconditioned JDSVD (IPJDSVD) method. Importantly, from another perspective, IPJDSVD can be considered to be a new JDSVD method with a new correction equation solved at each outer iteration, so that any applicable preconditioner to the correction equations in JDSVD can be straightforwardly used for the new correction equations in IPJDSVD.
In this paper, we first make a detailed convergence analysis on MINRES for the correction equation in JDSVD for computing one singular triplet $(\sigma_{1},u_{1},v_{1})$. Then we consider the case that there are $m$ clustered singular values closest to the target $\tau$, and show that MINRES may converge very slowly. For the preconditioned correction equation, we prove an extremely important property: its solution has the same quality as the solution to the standard correction equation when expanding the current subspaces; that is, the solution to the preconditioned correction equation expands the subspaces as effectively as the solution to the standard one and, therefore, does not spoil the subsequent outer iterations. Then we establish convergence results on MINRES for the preconditioned correction equation, and prove that it can converge substantially faster than it does for the standard one. In computations, we shall set up some necessary criteria and propose an adaptive approach to pick up appropriate approximate singular triplets to form the preconditioned correction equation. Importantly and practically, we propose a new thick-restart IPJDSVD algorithm with deflation and purgation that computes $\ell$ singular triplets of $A$. Whenever $m>k_{\min}$, the initial subspace dimension of the standard thick-restart, the new thick-restart is more effective than the standard one and saves outer iterations.
Since the goal of this paper is to propose an effective preconditioning procedure for the correction equation in JDSVD and a new thick-restart strategy, we shall take the standard extraction-based JDSVD method as example and the shifts in all the correction equations as the fixed $\tau$. However, the theoretical results and analysis on the correction equations and preconditioning are adaptable to the harmonic, refined and refined harmonic extraction-based JDSVD methods and their adaptively-changing-shift variants [9], so are the corresponding thick-restart algorithms.
In Section 2 we review the standard extraction-based JDSVD method. In Section 3 we make a convergence analysis on MINRES for the correction equation in JDSVD. In Section 4 we derive the preconditioned correction equation, and extend the results and analysis in Section 3 to the preconditioned correction equation. We consider how to dynamically form a preconditioned correction equation in computations. In Section 5 we propose the thick-restart IPJDSVD with deflation and purgation to compute $\ell$ singular triplets of $A$. Numerical experiments are reported in Section 6 to illustrate the high performance of the thick-restart IPJDSVD. In the meantime, to compare our algorithm with the currently most advanced PRIMME__\__SVDS software [27] in C language, in the same spirit we have also developed a hybrid two-stage thick-restarted IPJDSVD algorithm, called IPJDSVD_HYBRID: in the first stage, a similar IPJDSVD method computes the desired largest or smallest singular triplets based on the eigenproblem of $A^{T}A$ (or $AA^{T}$ if $m<n$) with the same outer stopping tolerance as that in the first stage of PRIMME_SVDS, and it then switches to the second stage whenever necessary, in which IPJDSVD is based on the eigenproblem of $\begin{bmatrix}\begin{smallmatrix}\mathbf{0}&A\\ A^{T}&\mathbf{0}\end{smallmatrix}\end{bmatrix}$ using the approximate singular triplets computed at the first stage as initial guesses, as done and detailed in this paper, and further improves approximate singular triplets until it converges for a reasonably prescribed stopping tolerance. We demonstrate that IPJDSVD_HYBRID substantially outperforms PRIMME_SVDS when computing the simple smallest singular triplets and is at least competitive with the latter when computing the simple largest ones. Section 7 concludes the paper.
Throughout this paper, denote by $\|X\|$ the 2-norm of a vector or matrix $X$, by $\kappa(X)=\sigma_{\max}(X)/\sigma_{\min}(X)$ the condition number of $X$ with $\sigma_{\max}(X)$ and $\sigma_{\min}(X)$ the largest and smallest singular values of $X$, respectively, and by $I_{i}$ and $\bm{0}_{i,j}$ the identity matrix of order $i$ and zero matrix of size $i\times j$, respectively. The subscripts of identity and zero matrices are omitted whenever their sizes are clear from the context.
2 The basic JDSVD method
We review the basic standard extraction-based JDSVD method [6] that computes the singular triplet $(\sigma_{1},u_{1},v_{1})$ of $A$.
At the $k$th outer iteration, given two $k$-dimensional left and right searching subspaces $\mathcal{U}\subset\mathbb{R}^{M}$ and $\mathcal{V}\subset\mathbb{R}^{N}$, the standard extraction approach seeks for scalars $\theta_{i}>0$ and unit-length vectors $\tilde{u}_{i}\in\mathcal{U}$ and $\tilde{v}_{i}\in\mathcal{V}$ that satisfy the double-orthogonality
(3) $\qquad A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i}\perp\mathcal{U}\quad\mbox{and}\quad A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i}\perp\mathcal{V},\qquad i=1,\dots,k.$
The $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ are the Ritz approximations of $A$ with respect to $\mathcal{U}$ and $\mathcal{V}$, and (3) is denoted as $(A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i},A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i})\perp\perp(\mathcal{U},\mathcal{V}),\ i=1,\dots,k.$ Label $|\theta_{1}-\tau|\leq|\theta_{2}-\tau|\leq\dots\leq|\theta_{k}-\tau|$. The method takes $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ as an approximation to $(\sigma_{1},u_{1},v_{1})$.
Let $\widetilde{U}\in\mathbb{R}^{M\times k}$ and $\widetilde{V}\in\mathbb{R}^{N\times k}$ be orthonormal basis matrices of $\mathcal{U}$ and $\mathcal{V}$. Then the left and right Ritz vectors $\tilde{u}_{i}=\widetilde{U}c_{i}$ and $\tilde{v}_{i}=\widetilde{V}d_{i}$ with unit-length $c_{i}\in\mathbb{R}^{k}$ and $d_{i}\in\mathbb{R}^{k}$. Define $H=\widetilde{U}^{T}A\widetilde{V}$. Then (3) is equivalent to
(4) $\qquad Hd_{i}=\theta_{i}c_{i}\qquad\mbox{and}\qquad H^{T}c_{i}=\theta_{i}d_{i},\qquad i=1,\dots,k.$
This means that the $(\theta_{i},c_{i},d_{i})$ are the singular triplets of $H$ and
(5) $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=(\theta_{1},\widetilde{U}c_{1},\widetilde{V}d_{1}),$
whose residual is defined as
(6) $r=r(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=\begin{bmatrix}A\tilde{v}_{1}-\theta_{1}\tilde{u}_{1}\\ A^{T}\tilde{u}_{1}-\theta_{1}\tilde{v}_{1}\end{bmatrix}.$
By (3), we particularly have
(7) $r\perp\perp(\tilde{u}_{1},\tilde{v}_{1}).$
$r=\bm{0}$ if and only if $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ is an exact singular triplet of $A$. If
(8) $\|r\|\leq\|A\|\cdot\textit{tol}$
for a user prescribed tolerance $\textit{tol}>0$, we claim that $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ converges.
If $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ does not yet converge, JDSVD expands $\mathcal{U}$ and $\mathcal{V}$ in the following way: It first solves the large symmetric and generally indefinite correction equation
(9) $\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r$
for $(s,t)\perp\perp(\tilde{u}_{1},\tilde{v}_{1})$ approximately using, e.g., the preferable MINRES method, and then orthonormalizes $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ to obtain $u_{+}$ and $v_{+}$. The columns of
$\widetilde{U}_{\operatorname{new}}=[\widetilde{U},u_{+}]\qquad\mbox{and}\qquad\widetilde{V}_{\operatorname{new}}=[\widetilde{V},v_{+}]$
are used to construct the expanded $(k+1)$-dimensional left and right subspaces
$\mathcal{U}_{\operatorname{new}}=\operatorname{span}\{\widetilde{U}_{\operatorname{new}}\}=\operatorname{span}\{\widetilde{U},\tilde{s}\}\quad\mbox{and}\quad\mathcal{V}_{\operatorname{new}}=\operatorname{span}\{\widetilde{V}_{\operatorname{new}}\}=\operatorname{span}\{\widetilde{V},\tilde{t}\}.$
JDSVD then computes a new approximation to $(\sigma_{1},u_{1},v_{1})$ with respect to $\mathcal{U}_{\operatorname{new}}$ and $\mathcal{V}_{\operatorname{new}}$.
3 Convergence analysis on MINRES for solving (9)
Observe that the coefficient matrix of system (9) is the restriction of $B=\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ to the subspace
$(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}:=\{(x,y)|x\in\mathbb{R}^{M},y\in\mathbb{R}^{N},x\perp\tilde{u}_{1},y\perp\tilde{v}_{1}\},$
and maps the subspace to itself. Let $\widetilde{P}_{1}\in\mathbb{R}^{M\times(M-1)}$ and $\widetilde{Q}_{1}\in\mathbb{R}^{N\times(N-1)}$ be such that the matrices $[\tilde{u}_{1},\widetilde{P}_{1}]$ and $[\tilde{v}_{1},\widetilde{Q}_{1}]$ are orthogonal. Then the columns of
(10) $\widetilde{W}_{1}=\begin{bmatrix}\widetilde{P}_{1}&\\ &\widetilde{Q}_{1}\end{bmatrix}\in\mathbb{R}^{(M+N)\times(M+N-2)}$
form an orthonormal basis of $(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$, and the orthogonal projector in (9) is
$\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}=\begin{bmatrix}\widetilde{P}_{1}\widetilde{P}_{1}^{T}&\\ &\widetilde{Q}_{1}\widetilde{Q}_{1}^{T}\end{bmatrix}=\widetilde{W}_{1}\widetilde{W}_{1}^{T}.$
Denote by $w=(s^{T},t^{T})^{T}$. Then the correction equation (9) is
(11) $\widetilde{B}_{1}w=-r,$
where
(12) $\widetilde{B}_{1}=\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\quad\ \mbox{with}\quad\ \widetilde{B}_{1}^{\prime}=\widetilde{W}_{1}^{T}B\widetilde{W}_{1}=\begin{bmatrix}-\tau I_{M-1}&\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}\\ \widetilde{Q}_{1}^{T}A^{T}\widetilde{P}_{1}&-\tau I_{N-1}\end{bmatrix}.$
Let $\tilde{w}_{j}=(\tilde{s}_{j}^{T},\tilde{t}_{j}^{T})^{T}$ be the approximate solution of (11) computed by MINRES at the $j$th iteration. Then the residual of $\tilde{w}_{j}$ is
(13) $r_{\mathrm{in},j}=\widetilde{B}_{1}\tilde{w}_{j}+r.$
We now present an equivalent form of (9) that will be used later.
Theorem 3.1.
Let $\widetilde{B}_{1}^{\prime}$ be defined as in (12). Then the solution $w=\widetilde{W}_{1}z$ to equation (9), where $z\in\mathbb{R}^{M+N-2}$ is the solution to the linear system
(14) $\widetilde{B}_{1}^{\prime}z=-r^{\prime}\qquad\mbox{with}\qquad r^{\prime}=\widetilde{W}_{1}^{T}r.$
If MINRES is applied to (11) and (14) with the initial guesses $\tilde{w}_{0}$ and $\tilde{z}_{0}$ related by $\widetilde{w}_{0}=\widetilde{W}_{1}\tilde{z}_{0}$, the $j$th approximate solutions $\tilde{w}_{j}$ and $\tilde{z}_{j}$ obtained by it satisfy $\tilde{w}_{j}=\widetilde{W}_{1}\tilde{z}_{j}$, $j=1,2,\ldots,M+N-2$. Write $r_{\mathrm{in},j}^{\prime}=\widetilde{B}_{1}^{\prime}\tilde{z}_{j}+r^{\prime}$. Then the residual norms
(15) $\|r_{\mathrm{in},j}\|=\|r^{\prime}_{\mathrm{in},j}\|,\qquad j=0,1,\ldots,M+N-2.$
Proof 3.2.
By (7) and (9), since $r\in(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$ and $w\in(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$, we have
$r=\widetilde{W}_{1}\widetilde{W}_{1}^{T}r=\widetilde{W}_{1}r^{\prime}\qquad\mbox{and}\qquad w=\widetilde{W}_{1}\widetilde{W}_{1}^{T}w=\widetilde{W}_{1}z$
with $r^{\prime}=\widetilde{W}_{1}^{T}r$ and $z=\widetilde{W}_{1}^{T}w$. Substituting them and (12) into (11) yields
(16) $\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\widetilde{W}_{1}z=\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}z=-\widetilde{W}_{1}r^{\prime},$
premultiplying which by $\widetilde{W}_{1}^{T}$ leads to (14).
Write the $j$-dimensional Krylov subspace
$\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})=\mathrm{span}\{r_{\mathrm{in},0},\widetilde{B}_{1}r_{\mathrm{in},0},\widetilde{B}_{1}^{2}r_{\mathrm{in},0},\dots,\widetilde{B}_{1}^{j-1}r_{\mathrm{in},0}\}.$
Since $\tilde{w}_{j}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})$ satisfies
(17) $\|r_{\mathrm{in},j}\|=\|\widetilde{B}_{1}\tilde{w}_{j}+r\|=\min_{\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})}\|\widetilde{B}_{1}\tilde{w}+r\|,$
by definition it is straightforward to verify from (12) and (13) that
$\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})=\mathcal{K}_{j}(\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T},\widetilde{W}_{1}r^{\prime}_{\mathrm{in},0})=\widetilde{W}_{1}\mathcal{K}_{j}(\widetilde{B}_{1}^{\prime},r_{\mathrm{in},0}^{\prime});$
that is, for any $\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})$, there exists $\tilde{z}\in\mathcal{K}_{j}(\widetilde{B}_{1}^{\prime},r_{\mathrm{in},0}^{\prime})$ such that $\tilde{w}=\widetilde{W}_{1}\tilde{z}$ and thus $\tilde{z}=\widetilde{W}_{1}^{T}\tilde{w}$, $j=1,2,\ldots,M+N-2$. Therefore, from (17) we obtain
$\|r_{\mathrm{in},j}\|=\min_{\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})}\|\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\tilde{w}+\widetilde{W}_{1}r^{\prime}\|=\min_{\tilde{z}\in\mathcal{K}_{j}(\widetilde{B}_{1}^{\prime},r_{\mathrm{in},0}^{\prime})}\|\widetilde{B}_{1}^{\prime}\tilde{z}+r^{\prime}\|=\|r_{\mathrm{in},j}^{\prime}\|,$
which proves (15) and $\tilde{w}_{j}=\widetilde{W}_{1}\tilde{z}_{j}$.
Theorem 3.1 shows that the convergence of MINRES for (11) and (14) is identical. Without loss of generality, take $\tilde{w}_{0}=\bm{0}$ and $\tilde{z}_{0}=\bm{0}$. Then
(18) $r_{\mathrm{in},0}=r\qquad\mbox{and}\qquad r_{\mathrm{in},0}^{\prime}=r^{\prime}.$
Denote by $\tilde{\lambda}_{i},i=1,\dots,M+N-2$ the eigenvalues of $\widetilde{B}_{1}^{\prime}$ in (12). It is known from [4] that the residual norm of the approximate solution $\tilde{z}_{j}$ of (14) satisfies
(19) $\frac{\|r_{\mathrm{in},j}^{\prime}\|}{\|r^{\prime}\|}\leq\min_{\pi\in\Pi_{j}}\max_{1\leq i\leq M+N-2}|\pi(\tilde{\lambda}_{i})|,$
where $\Pi_{j}$ denotes the set of polynomials of degree no more than $j$ and $\pi(0)=1$ for $\pi\in\Pi_{j}$. Keep in mind (15). Then there are the following estimates.
Theorem 3.3 ([4, Chapter 3]).
(i) If the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in the interval $[-\beta,-\alpha]$ or $[\alpha,\beta]$ with $0<\alpha<\beta$, then
(20) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\tfrac{2}{1+\sqrt{\beta}/\sqrt{\alpha}}\right)^{j};$
(ii) If the $\tilde{\lambda}_{i}\in[-\beta_{1},-\alpha_{1}]\cup[\alpha_{2},\beta_{2}]$ with $\alpha_{1},\alpha_{2},\beta_{1},\beta_{2}>0$ and $\beta_{2}-\alpha_{2}=\beta_{1}-\alpha_{1}$, then
(21) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{\beta_{1}\beta_{2}}/\sqrt{\alpha_{1}\alpha_{2}}}\right)^{\left[\frac{j}{2}\right]},$
where $\left[\frac{j}{2}\right]$ denotes the integer part of $\frac{j}{2}$.
This theorem indicates that the smaller $\beta-\alpha$ or $\beta_{1}-\alpha_{1}$ is, the closer $\beta/\alpha$ or $(\beta_{1}\beta_{2})/(\alpha_{1}\alpha_{2})$ is to one and thus the faster MINRES converges. Bound (20) is the same as that for the CG method for solving the same linear system [4], where the convergence factor is $\frac{\sqrt{\kappa(\widetilde{B}_{1}^{\prime})}-1}{\sqrt{\kappa(\widetilde{B}_{1}^{\prime})}+1}$ with $\kappa(\widetilde{B}_{1}^{\prime})=\beta/\alpha$. Suppose that $[-\beta_{1},-\alpha_{1}]$ and $[\alpha_{2},\beta_{2}]$ are symmetric with respect to the origin, i.e., $\alpha_{1}=\alpha_{2}=\alpha$, $\beta_{1}=\beta_{2}=\beta$. Then the factor in the right-hand side of (21) is $\frac{\kappa(\widetilde{B}_{1}^{\prime})-1}{\kappa(\widetilde{B}_{1}^{\prime})+1}$, the same as that for the steepest gradient method applied to a positive definite (14) with the eigenvalue interval $[\alpha,\beta]$, but the power is $\left[\frac{j}{2}\right]$ rather than $j$. As a result, MINRES may converge much more slowly for a highly indefinite linear system than it does for a definite one.
Before proceeding, we should notice that $r=\mathbf{0}$ when $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=(\sigma_{1},u_{1},v_{1})$. In this case we terminate JDSVD. For later use, we next suppose an arbitrary virtual right-hand side $r\not=\mathbf{0}$ satisfying (7) when $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$. Based on Theorems 3.13.3, we present the following convergence results on MINRES for (9).
Theorem 3.4.
For $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$ in (9), and suppose virtually $r\not=\mathbf{0}$ satisfying $r\perp\perp(u_{1},v_{1})$. Then (i) if $\tau>(\sigma_{\max}+\sigma_{\max,2})/2$ with $\sigma_{\max,2}$ being the second largest singular value of $A$, then $\sigma_{1}=\sigma_{\max}$ and
(22) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,2})/(\tau-\sigma_{\max,2})}}\right)^{j};$
(ii) if $\sigma_{\min}\leq\tau<(\sigma_{\min}+\sigma_{\min,2})/2$ with $\sigma_{\min,2}$ being the second smallest singular value of $A$, then $\sigma_{1}=\sigma_{\min}$ and
(23) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\Bigg{(}1-\frac{2}{1+\sqrt{(\sigma_{\max}^{2}-\tau^{2})/(\sigma_{\min,2}^{2}-\tau^{2})}}\Bigg{)}^{\left[\frac{j-j_{o}}{2}\right]},$
where $j_{o}=\min\{1,M-N\}$; (iii) if $(\sigma_{\min}+\sigma_{\min,2})/2<\tau<(\sigma_{\max}+\sigma_{\max,2})/2$, then $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$ and
(24) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau)/|\tau-\sigma_{2}|}\right)^{\left[\frac{j}{2}\right]}.$
Proof 3.5.
By (1), for $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$, take $\widetilde{P}_{1}=P_{1}$ and $\widetilde{Q}_{1}=Q_{1}$ in (10) with
(25) $P_{1}=[u_{2},\dots,u_{M}]\qquad\mbox{and}\qquad Q_{1}=[v_{2},\dots,v_{N}],$
and rewrite the resulting $\widetilde{B}_{1}^{\prime}=B_{1}^{\prime}$ in (12). Then
(26) $B_{1}^{\prime}=\begin{bmatrix}-\tau I_{M-1}&\Sigma_{1}^{\prime}\\ {\Sigma_{1}^{\prime}}^{T}&-\tau I_{N-1}\end{bmatrix}\qquad\mbox{with}\qquad\Sigma_{1}^{\prime}=P_{1}^{T}AQ_{1}=\operatorname{diag}\{\sigma_{2},\dots,\sigma_{N}\}.$
Therefore, the eigenvalues $\lambda_{i}$ of $B_{1}^{\prime}$ are
(27) ±$\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau\qquad\mbox{and}\qquad\underbrace{-\tau,\dots,-\tau}_{M-N}.$
Case (i): In this case, $\sigma_{\max}$ is the closest to $\tau$, i.e., $\sigma_{1}=\sigma_{\max}$, all the singular values but $\sigma_{1}$ of $A$ are smaller than $\tau$, and
(28) $\lambda_{i}\in[-(\sigma_{\max,2}+\tau),-(\tau-\sigma_{\max,2})]$
by noticing that $\sigma_{2}=\sigma_{\max,2}$. Exploiting (20), we obtain (22).
Case (ii): In this case, we have $\sigma_{1}=\sigma_{\min}$ and
$\lambda_{i}\in\{-\tau\}\cup[-(\sigma_{\max}+\tau),-(\sigma_{\min,2}+\tau)]\cup[\sigma_{\min,2}-\tau,\sigma_{\max}-\tau],$
where $\{-\tau\}$ consists of an isolated point $-\tau$, which is an $(M-N)$-multiple eigenvalue of $B_{1}^{\prime}$ and disappears for $M=N$. Construct a degree $j$ polynomial $q\in\Pi_{j}$ of form
$q(\lambda)=\left(\tfrac{\tau+\lambda}{\tau}\right)^{j_{o}}\pi(\lambda)\qquad\mbox{with}\qquad\pi\in\Pi_{j-j_{o}},$
where $j_{o}=\min\{1,M-N\}$. Since $q(0)=\pi(0)=1$ and $\left|\frac{\tau+\lambda}{\tau}\right|^{j_{o}}\leq\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}$ over the eigenvalues in (27), by Theorem 3.1 and relations (18)–(19), we obtain
$\frac{\|r_{\mathrm{in},j}\|}{\|r\|}$$\leq$ $\min_{\pi\in\Pi_{j-j_{o}}}\max_{\lambda_{i}\in\{-\tau,\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau\}}|q(\lambda_{i})|$
$\leq$ $\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\min_{\pi\in\Pi_{j-j_{o}}}\max_{i=2,\ldots,N}|\pi(\pm\sigma_{i}-\tau)|$
$\leq$ $\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\min_{\pi\in\Pi_{j-j_{o}}}\max_{\lambda\in\Delta}|\pi(\lambda)|,$
where
$\Delta=[-(\sigma_{\max}+\tau),-(\sigma_{\min,2}+\tau)]\cup[\sigma_{\min,2}-\tau,\sigma_{\max}-\tau].$
Notice that the above positive and negative intervals have the same lengths $\sigma_{\max}-\sigma_{\min,2}$. Bound (23) follows from applying (21) to the above min-max problem.
Case (iii): In this case, denote by $\sigma_{\rm L}$ and $\sigma_{\rm R}$ the singular values of $A$ closest to $\tau$ from the left and right among $\sigma_{2},\dots,\sigma_{N}$, respectively. Then the $\lambda_{i}$ lie in
$[-(\sigma_{\max}+\tau),-(\tau-\sigma_{\rm L})]\cup[\sigma_{\rm R}-\tau,\sigma_{\max}-\tau]$
(29) $\subset[-(\sigma_{\max}+\tau),-|\tau-\sigma_{2}|]\cup[|\tau-\sigma_{2}|,\sigma_{\max}+\tau].$
Notice that the two intervals in (29) have the same lengths $\sigma_{\max}+\tau-|\tau-\sigma_{2}|$. Then (23) follows from applying (21) to the interval union in (29).
Remark 3.6.
The convergence rates of MINRES respectively depend on
(30) $\gamma_{1}=\frac{\tau-\sigma_{\max,2}}{\tau+\sigma_{\max,2}},\qquad\gamma_{2}=\frac{\sigma_{\min,2}^{2}-\tau^{2}}{\sigma_{\max}^{2}-\tau^{2}},\qquad\gamma_{3}=\frac{|\tau-\sigma_{2}|}{\sigma_{\max}+\tau}.$
The better $\sigma_{\max,2}$, $\sigma_{\min,2}$ or $\sigma_{2}$ is separated from $\tau$, the smaller $1-\frac{2\sqrt{\gamma_{1}}}{1+\sqrt{\gamma_{1}}}$, $1-\frac{2\sqrt{\gamma_{2}}}{1+\sqrt{\gamma_{2}}}$ or $1-\frac{2\gamma_{3}}{1+\gamma_{3}}$ in (22)–(24) is, causing that MINRES converges faster.
Remark 3.7.
For $\sigma_{1}=\sigma_{\max}$ or $\sigma_{\min}$, which corresponds to $\tau\approx\sigma_{\max}$ or $\sigma_{\min}$, we have
$\gamma_{1}\approx\frac{\sigma_{\max}-\sigma_{\max,2}}{\sigma_{\max}+\sigma_{\max,2}}\qquad\mbox{or}\qquad\gamma_{2}\approx\frac{\sigma_{\min,2}^{2}-\sigma_{\min}^{2}}{\sigma_{\max}^{2}-\sigma_{\min}^{2}}.$
Suppose $\sigma_{\max,2}\approx\sigma_{\max}$ and $\sigma_{\min,2}\approx\sigma_{\min}$ roughly. Then
$\frac{\gamma_{1}}{\gamma_{2}}\approx\frac{\sigma_{\max}(\sigma_{\max}-\sigma_{\max,2})}{\sigma_{\min}(\sigma_{\min,2}-\sigma_{\min})}\gg 1$
if $\sigma_{\max}-\sigma_{\max,2}\approx\sigma_{\min,2}-\sigma_{\min}$ and $\sigma_{\max}\gg\sigma_{\min}$, and the ratio is even bigger if $\sigma_{\max}-\sigma_{\max,2}>\sigma_{\min,2}-\sigma_{\min}$ considerably, a typical case in applications. As a consequence, MINRES is generally much faster for $\sigma_{1}=\sigma_{\max}$ than it is for $\sigma_{1}=\sigma_{\min}$.
Starting with Theorem 3.4, we are able to analyze the convergence of MINRES when $\tilde{u}_{1}\approx u_{1}$ and $\tilde{v}_{1}\approx v_{1}$. First, we show that $\widetilde{B}_{1}^{\prime}$ in (12) is orthogonally similar to $B_{1}^{\prime}$ in (26) with squared errors of $\tilde{u}_{1}$ and $\tilde{v}_{1}$.
Theorem 3.8.
Let $\phi=\angle(\tilde{u}_{1},u_{1})$ and $\psi=\angle(\tilde{v}_{1},v_{1})$ be the acute angles between $\tilde{u}_{1}$ and $u_{1}$ and between $\tilde{v}_{1}$ and $v_{1}$. Then $\widetilde{B}_{1}^{\prime}$ is orthogonally similar to $B_{1}^{\prime}$ with the error
(31) $\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|\leq\|A\|\left(\sin\phi+\sin\psi\right)^{2},$
where $Z$ is some $(M+N-2)$-by-$(M+N-2)$ orthogonal matrix; see (39).
Proof 3.9.
By (25), expand the orthonormal $\widetilde{P}_{1}$ and $\widetilde{Q}_{1}$ in (10) as the orthogonal sums
(32) $\widetilde{P}_{1}=P_{1}E_{1}+u_{1}E_{2}\qquad\mbox{and}\qquad\widetilde{Q}_{1}=Q_{1}F_{1}+v_{1}F_{2},$
where $E_{1}\in\mathbb{R}^{(M-1)\times(M-1)}$, $E_{2}\in\mathbb{R}^{1\times(M-1)}$ and $F_{1}\in\mathbb{R}^{(N-1)\times(N-1)}$, $F_{2}\in\mathbb{R}^{1\times(N-1)}$. Then
(33) $\sin\phi=\|\widetilde{P}_{1}^{T}u_{1}\|=\|E_{2}\|\qquad\mbox{and}\qquad\sin\psi=\|\widetilde{Q}_{1}^{T}v_{1}\|=\|F_{2}\|.$
Let the SVDs of $E_{1}$ and $F_{1}$ be
(34) $X_{1}^{T}E_{1}X_{2}=\Sigma_{E}\qquad\mbox{and}\qquad Y_{1}^{T}F_{1}Y_{2}=\Sigma_{F}.$
Notice $\|I-\Sigma_{E}\|\leq\|I-\Sigma_{E}^{T}\Sigma_{E}\|$ as $\|\Sigma_{E}\|\leq 1$ and the diagonal $\Sigma_{E}$ is square and nonnegative definite. Making use of (34) and $E_{1}^{T}E_{1}+E_{2}^{T}E_{2}=I$, we obtain
(35) $\|I-\Sigma_{E}\|\leq\|X_{2}^{T}(I-E_{1}^{T}E_{1})X_{2}\|=\|E_{2}^{T}E_{2}\|=\|E_{2}\|^{2}=\sin^{2}\phi,$
where the last equality holds by (33). Likewise,
(36) $\|I-\Sigma_{F}\|\leq\sin^{2}\psi.$
Since $P_{1}^{T}Av_{1}=\bm{0}$ and $u_{1}^{T}AQ_{1}=\bm{0}$, it is known from (26) and (32) that $\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}=E_{1}^{T}\Sigma_{1}^{\prime}F_{1}+\sigma_{1}E_{2}^{T}F_{2}$. Exploiting (34), we have
(37) $\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}=X_{2}\Sigma_{E}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}\Sigma_{F}Y_{2}^{T}+\sigma_{1}E_{2}^{T}F_{2}.$
Note that $X_{i},Y_{i}$, $i=1,2$ are orthogonal matrices. Subtracting $X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}$ from both hand sides of (37) and taking $2$-norms, we obtain
(38) $\|\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}\!-\!X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}\|$ $=$ $\|\Sigma_{E}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}\Sigma_{F}-X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}+\sigma_{1}X_{2}^{T}E_{2}^{T}F_{2}Y_{2}\|$
$\leq$ $\|\!(\Sigma_{E}\!-\!I)X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}\Sigma_{F}\!+\!X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}(\Sigma_{F}\!-\!I)\!\|\!+\!\sigma_{1}\|E_{2}^{T}\!F_{2}\|$
$\leq$ $\|\Sigma_{E}-I\|\|\Sigma_{1}^{\prime}\|+\|\Sigma_{1}^{\prime}\|\|\Sigma_{F}-I\|+\sigma_{1}\|E_{2}\|\|F_{2}\|$
$\leq$ $\|A\|(\sin^{2}\phi+\sin^{2}\psi+\sin\phi\sin\psi)$
$\leq$ $\|A\|(\sin\phi+\sin\psi)^{2},$
where the third inequality holds by making use of (33), (35), (36) and $\|\Sigma_{1}^{\prime\prime}\|=\|\Sigma_{1}^{\prime}\|\leq\|A\|$, $\sigma_{1}\leq\|A\|$. Define the orthogonal matrix
(39) $Z=\begin{bmatrix}X_{1}X_{2}^{T}&\\ &Y_{1}Y_{2}^{T}\end{bmatrix}\in\mathbb{R}^{(M+N-2)\times(M+N-2)}.$
By the definitions of $\widetilde{B}_{1}^{\prime}$ and $B_{1}^{\prime}$ in (12) and (26), a manipulation shows
$\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|=\|\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}-X_{2}X_{1}^{T}\Sigma_{1}^{\prime}Y_{1}Y_{2}^{T}\|.$
Applying (38) to the above establishes (31).
With the help of Theorems 3.4 and 3.8, we can establish the convergence results on MINRES for the correction equation (11), i.e., (9).
Theorem 3.10.
With the notations of Theorem 3.8, suppose that $\tilde{u}_{1}$ and $\tilde{v}_{1}$ satisfy
(40) $\delta:=\|A\|(\sin\phi+\sin\psi)^{2}<|\sigma_{2}-\tau|.$
(i) If $\sigma_{1}=\sigma_{\max}$, then
(41) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,2}+\delta)/(\tau-\sigma_{\max,2}-\delta)}}\right)^{j};$
(ii) if $\sigma_{1}\!=\!\sigma_{\min}$ and writing $j_{o}\!=\!\min\{1,M\!-\!N\}$, then
(42) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}+\delta}{\tau}\right)^{j_{o}}\Bigg{(}1-\frac{2}{1+\sqrt{\frac{(\sigma_{\max}+\delta)^{2}-\tau^{2}}{(\sigma_{\min,2}-\delta)^{2}-\tau^{2}}}}\Bigg{)}^{\left[\frac{j-j_{o}}{2}\right]};$
(iii) if $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$, then
(43) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau+\delta)/(|\sigma_{2}-\tau|-\delta)}\right)^{\left[\frac{j}{2}\right]}.$
Proof 3.11.
Since the matrix $Z$ in (39) is orthogonal, the eigenvalues $\lambda_{i}$ of $Z^{T}B_{1}^{\prime}Z$ are those of $B_{1}^{\prime}$, i.e., the ones in (27). Label the eigenvalues $\lambda_{i}$ and $\tilde{\lambda}_{i}$ of $B_{1}^{\prime}$ and $\widetilde{B}_{1}^{\prime}$ in the ascending or descending order. By the eigenvalue perturbation theory (cf. Theorem 10.3.1 of [20]), we obtain
(44) $|\tilde{\lambda}_{i}-\lambda_{i}|\leq\|\widetilde{B}_{1}^{\prime}-Z^{T}B_{1}^{\prime}Z\|\leq\delta,$
where the last inequality holds by (31) and (40).
(i) For $\sigma_{1}=\sigma_{\max}$, since the eigenvalues $\lambda_{i}$ of $Z^{T}B_{1}^{\prime}Z$ lie in the interval in (28), bound (44) shows that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
(45) $[-(\sigma_{\max,2}+\tau)-\delta,-(\tau-\sigma_{\max,2})+\delta].$
Note that $\sigma_{\max,2}=\sigma_{2}$. Condition (40) indicates $\tau-\sigma_{\max,2}>\delta$. Then taking $\alpha=\tau-\sigma_{\max,2}-\delta$ and $\beta=\tau+\sigma_{\max,2}+\delta$ in (20) of Theorem 3.3 proves (41).
(ii) Since $\widetilde{B}_{1}^{\prime}$ has an $(M-N)$-multiple eigenvalue $\tilde{\lambda}=-\tau$ for $M>N$, for $\sigma_{1}=\sigma_{\min}$, (27) and (44) indicate that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
$\{-\tau\}\cup[-(\sigma_{\max}+\tau)-\delta,-(\sigma_{\min,2}+\tau)+\delta]\cup[\sigma_{\min,2}-\tau-\delta,\sigma_{\max}-\tau+\delta].$
Notice that $\sigma_{\min,2}=\sigma_{2}$ and (40) mean that $\sigma_{\min,2}-\tau>\delta$, and observe that the above positive and negative intervals have the same lengths $\sigma_{\max}-\sigma_{\min,2}+2\delta$. Following the proof of item (ii) of Theorem 3.4, we obtain bound (42).
(iii) For $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$, since the eigenvalues of $Z^{T}B_{1}^{\prime}Z$ lie in the interval defined by (29), bound (44) means that the eigenvalues $\tilde{\lambda}_{i}$ of $\widetilde{B}_{1}^{\prime}$ lie in
$[-(\sigma_{\max}+\tau)-\delta,-|\sigma_{2}-\tau|+\delta]$$\cup$ $[|\sigma_{2}-\tau|-\delta,\sigma_{\max}-\tau+\delta]$
(46) $\subset\ [-(\sigma_{\max}+\tau)-\delta,-|\sigma_{2}-\tau|+\delta]$$\cup$ $[|\sigma_{2}-\tau|-\delta,\sigma_{\max}+\tau+\delta],$
Since the two intervals in (46) have the equal length $\sigma_{\max}+\tau-|\sigma_{2}-\tau|+2\delta$, applying the proof of item (iii) of Theorem 3.4 to this interval union establishes (43).
Remark 3.12.
It can be verified that the factors in (41)–(43) are approximations to $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ in (30) with the relative errors $\mathcal{O}(\frac{\delta}{\tau-\sigma_{\max,2}})$, $\mathcal{O}(\frac{\delta}{\sigma_{\min,2}-\tau})$ or $\mathcal{O}(\frac{\delta}{|\sigma_{2}-\tau|})$, respectively. Therefore, for $\delta$ reasonably small, Remarks 3.6 and 3.7 apply to Theorem 3.10.
4 A preconditioned correction equation, MINRES, and practical issues
Recall (25)–(27), and write $\widetilde{W}_{1}=W_{1}$ in (10) for $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$. Then the normalized eigenvectors of $\widetilde{B}_{1}=B_{1}=W_{1}B_{1}^{\prime}W_{1}^{T}$, defined in (12), associated with the eigenvalues ±$\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau$ and $-\tau$ are $\frac{1}{\sqrt{2}}[u_{i}^{T},\pm v_{i}^{T}]^{T},\ i=2,\ldots,N$ and $[u^{T}_{i},\mathbf{0}]^{T}$, $i=N+1,N+2,\ldots,M$, respectively; see (1). Therefore, those smallest eigenvalues in magnitude of $B_{1}$ correspond to the singular values of $A$ closest to $\tau$. Notice that for $\tilde{u}_{1}\not=u_{1}$ and $\tilde{v}_{1}\not=v_{1}$, since $\widetilde{B}_{1}\rightarrow B_{1}$ as $\tilde{u}_{1}\rightarrow u_{1}$ and $\tilde{v}_{1}\rightarrow v_{1}$, the eigenvalues and eigenvectors (rigorously speaking, the eigenspaces associated with each of the well isolated clusters of eigenvalues) of $\widetilde{B}_{1}$ converge to those of $B_{1}$ correspondingly. As a result, for $\tilde{u}_{1}\rightarrow u_{1}$ and $\tilde{v}_{1}\rightarrow v_{1}$, if the clustered smallest eigenvalues of $B_{1}$ are well separated from the others, then $\widetilde{B}_{1}$ has the same number of clustered smallest eigenvalues in magnitude.
The preceding convergence results show that MINRES may converge very slowly when $\sigma_{1}$ is clustered with some other singular values of $A$, which corresponds to the case that the clustered smallest eigenvalues in magnitude of $\widetilde{B}_{1}$ are well separated from the others. In this case, it is vital to precondition the correction equation (9) in order to accelerate MINRES. Theoretically, this can be realized by annihilating the effects of those smallest eigenvalues in magnitude of $\widetilde{B}_{1}$. Strikingly, as JDSVD proceeds, the subspaces $\mathcal{U}$ and $\mathcal{V}$ generally contain increasingly more accurate approximations to the singular vectors of $A$ associated with several singular values closest to $\tau$. Based on the last paragraph, this property indicates that JDSVD provides rich information on the eigenvectors of $B_{1}$ and thus those of $\widetilde{B}_{1}$ associated with the smallest eigenvalues in magnitude. It, therefore, makes us possible to exploit such information to precondition the correction equation effectively so as to solve it by MINRES more efficiently, which leads to our inner preconditioned JDSVD (IPJDSVD) method. We will show that such preconditioning is equivalent to forming a preconditioned correction equation that makes full use of all the approximate left and right singular vectors associated with the clustered singular values nearest to $\tau$. If the clustered singular values closest to $\tau$ are well isolated with the other singular values of $A$, those smallest eigenvalues in magnitude of the preconditioned coefficient matrix are well separated from the other eigenvalues, too, so that the convergence of MINRES is improved substantially.
Subsection 4.1 is devoted to a detailed derivation of the preconditioned correction equation; Section 4.2 presents convergence results on MINRES for solving it; Section 4.3 settles down the practical issue of how to adaptively form the preconditioned correction equation at each outer iteration.
Suppose that there are $m$ clustered singular values of $A$ closest to $\tau$:
(47) $|\sigma_{1}-\tau|\lesssim\dots\lesssim|\sigma_{m}-\tau|\ll|\sigma_{m+1}-\tau|\leq\dots\leq|\sigma_{N}-\tau|.$
From (1), we denote by
(48) $\Sigma_{m}=\operatorname{diag}\{\sigma_{1},\dots,\sigma_{m}\},\qquad U_{m}=[u_{1},\dots,u_{m}],\qquad V_{m}=[v_{1},\dots,v_{m}].$
Such $(\Sigma_{m},U_{m},V_{m})$ is a partial SVD of $A$ with the $m$ singular values closest to $\tau$.
At the $k$th outer iteration of JDSVD, suppose that $k>m$. Recall from Section 2 that $(\theta_{i},c_{i},d_{i})$, $i=1,\dots,k$ are the singular triplets of $H$ in (4), labeled as $|\theta_{1}-\tau|\leq|\theta_{2}-\tau|\leq\dots\leq|\theta_{k}-\tau|$. We compute the $k$ Ritz approximations
(49) $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})=(\theta_{i},\widetilde{U}c_{i},\widetilde{V}d_{i}),\qquad i=1,\dots,k,$
and take $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ to approximate the desired $(\sigma_{1},u_{1},v_{1})$. Denote
(50) $\Theta_{m}=\operatorname{diag}\{\theta_{1},\dots,\theta_{m}\},\qquad\widetilde{U}_{m}=[\tilde{u}_{1},\dots,\tilde{u}_{m}],\qquad\widetilde{V}_{m}=[\tilde{v}_{1},\dots,\tilde{v}_{m}].$
Then $(\Theta_{m},\widetilde{U}_{m},\widetilde{V}_{m})$ is an approximation to $(\Sigma_{m},U_{m},V_{m})$.
4.1 A preconditioned correction equation
By (3) and (6), from (49)–(50) we have
(51) $r\perp\perp\left(\widetilde{U}_{m},\widetilde{V}_{m}\right)\mbox{,\ \ i.e.,\ \ }r=\operatorname{diag}\{I-\widetilde{U}_{m}\widetilde{U}_{m}^{T},I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\}r.$
Therefore, premultiplying (9) by $\operatorname{diag}\{I-\widetilde{U}_{m}\widetilde{U}_{m}^{T},I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\}$ yields
(52) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r,$
where we have exploited the fact that $(s,t)\perp\perp(\tilde{u}_{1},\tilde{v}_{1})$ and the vectors $\tilde{u}_{1}$ and $\tilde{v}_{1}$ are the first columns of $\widetilde{U}_{m}$ and $\widetilde{V}_{m}$, respectively. Writing
$\begin{bmatrix}s\\ t\end{bmatrix}=\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}+\begin{bmatrix}\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}$
and substituting this relation into (52), we obtain
(53) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}\!\!\!\!&\\ &\!\!\!\!I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}\!\!\!\!&\\ &\!\!\!\!I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r-r_{\mathrm{tail}},$
where, by $\Theta_{m}=\widetilde{U}_{m}^{T}A\widetilde{V}_{m}$,
(54) $r_{\mathrm{tail}}$ $=$ $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}$
$=$ $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}&\\ &\widetilde{V}_{m}\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}^{T}s\\ \widetilde{V}_{m}^{T}t\end{bmatrix}$
$=$ $\begin{bmatrix}&A\widetilde{V}_{m}-\widetilde{U}_{m}\Theta_{m}\\ A^{T}\widetilde{U}_{m}-\widetilde{V}_{m}\Theta_{m}\end{bmatrix}\begin{bmatrix}\widetilde{U}_{m}^{T}s\\ \widetilde{V}_{m}^{T}t\end{bmatrix}.$
From (9), the first elements of $\widetilde{U}_{m}^{T}s$ and $\widetilde{V}_{m}^{T}t$ are $\tilde{u}_{1}^{T}s=0$ and $\tilde{v}_{1}^{T}t=0$. Denote $\widetilde{U}_{m}^{\prime}=[\tilde{u}_{2},\dots,\tilde{u}_{m}]$, $\widetilde{V}_{m}^{\prime}=[\tilde{v}_{2},\dots,\tilde{v}_{m}]$ and $\Theta_{m}^{\prime}=[\theta_{2},\dots,\theta_{m}]$. Then (54) becomes
(55) $r_{\mathrm{tail}}=\begin{bmatrix}&\!\!\!\!A\widetilde{V}_{m}^{\prime}-\widetilde{U}_{m}^{\prime}\Theta_{m}^{\prime}\\ A^{T}\widetilde{U}_{m}^{\prime}-\widetilde{V}_{m}^{\prime}\Theta_{m}^{\prime}&\end{bmatrix}\begin{bmatrix}(\widetilde{U}_{m}^{\prime})^{T}s\\ (\widetilde{V}_{m}^{\prime})^{T}t\end{bmatrix}=\begin{bmatrix}&\!\!\!\!R_{1}\\ R_{2}&\end{bmatrix}\begin{bmatrix}(\widetilde{U}_{m}^{\prime})^{T}s\\ (\widetilde{V}_{m}^{\prime})^{T}t\end{bmatrix},$
where $R_{1}=A\widetilde{V}_{m}^{\prime}-\widetilde{U}_{m}^{\prime}\Theta_{m}^{\prime}$ and $R_{2}=A^{T}\widetilde{U}_{m}^{\prime}-\widetilde{V}_{m}^{\prime}\Theta_{m}^{\prime}$, whose $(i-1)$-th columns are respectively the upper and lower parts of the residual $r^{(i)}=r(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ of $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ defined by (6), $i=2,\dots,m$. Keep in mind the three properties: (i) $\|s\|$ and $\|t\|$ are of the same orders as the errors of $\tilde{u}_{i}$ and $\tilde{v}_{i}$, respectively; (ii) $\|r^{(i)}\|$ is of the same order as the errors of $\tilde{u}_{i},\tilde{v}_{i}$, $i=1,\dots,m$; (iii) $(\widetilde{U}_{m}^{\prime})^{T}s$ and $(\widetilde{V}_{m}^{\prime})^{T}t$ are no more than $s$ and $t$ in size. Therefore, whenever $\|r^{(i)}\|,i=2,\ldots,m$ are comparable to $\|r\|$, one has $\|r_{\rm tail}\|=\mathcal{O}(\|r\|^{2})$ and thus $r+r_{\rm tail}\approx r$ in (53) once $\|r\|$ becomes fairly small.
Omitting $r_{\mathrm{tail}}$ in (53), we have now derived a preconditioned correction equation:
(56) $\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\widetilde{U}_{m}\widetilde{U}_{m}^{T}&\\ &I-\widetilde{V}_{m}\widetilde{V}_{m}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r,$
whose solution $(s,t)\perp\perp(\widetilde{U}_{m},\widetilde{V}_{m})$ has the same quality as the solution $(s,t)$ to (9) provided that $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=2,\dots,m$ have comparable accuracy to $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$. The resulting method is the basic IPJDSVD, and can also be regarded as a new JDSVD method with the correction equation (56) solved at each outer iteration.
With an approximate solution $(\tilde{s}^{T},\tilde{t}^{T})^{T}$ of (56) available, we orthonormalize $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ to obtain $u_{+}$ and $v_{+}$, and expand the subspaces in a regular way. The above derivation and analysis indicate that the solution to equation (56) expands the searching subspaces as effectively as the solution to the standard correction equation (9) does.
4.2 Convergence analysis on MINRES for solving (56)
Let $\widetilde{P}_{m}$ and $\widetilde{Q}_{m}$ make $[\widetilde{U}_{m},\widetilde{P}_{m}]$ and $[\widetilde{V}_{m},\widetilde{Q}_{m}]$ orthogonal, and define the column orthonormal matrix
(57) $\widetilde{W}_{m}=\begin{bmatrix}\widetilde{P}_{m}&\\ &\widetilde{Q}_{m}\end{bmatrix}\in\mathbb{R}^{(M+N)\times(M+N-2m)}.$
Then the coefficient matrix in (56) can be written as
(58) $\widetilde{B}_{m}=\widetilde{W}_{m}\widetilde{B}_{m}^{\prime}\widetilde{W}_{m}^{T}\quad\mbox{with}\quad\widetilde{B}_{m}^{\prime}=\begin{bmatrix}-\tau I_{M-m}&\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}\\ \widetilde{Q}_{m}^{T}A^{T}\widetilde{P}_{m}&-\tau I_{N-m}\end{bmatrix}.$
Denote by $\tilde{w}_{j}$ the $j$th approximate solution computed by MINRES for (56) with $\tilde{w}_{0}=\bm{0}$, $j=1,2,\ldots,M+N-2m$. Then the corresponding inner residuals are
(59) $\qquad r_{\mathrm{in},j}=\widetilde{B}_{m}\tilde{w}_{j}+r,\qquad\qquad j=0,1,\ldots,M+N-2m.$
Using a similar proof to that of Theorem 3.1, we can obtain the following results.
Theorem 4.1.
With $\widetilde{B}_{m}^{\prime}$ defined in (58), the solution to the preconditioned correction equation (56) is $w=\widetilde{W}_{m}z$ where $z=\widetilde{W}_{m}^{T}w\in\mathbb{R}^{M+N-2m}$ solves
(60) $\widetilde{B}_{m}^{\prime}z=-r^{\prime}\qquad\mbox{with}\qquad r^{\prime}=\widetilde{W}_{m}^{T}r.$
If MINRES is applied to (60) with the initial guess $\tilde{z}_{0}=\bm{0}$, then the $j$th approximate solution $\tilde{z}_{j}=\widetilde{W}_{m}^{T}\tilde{w}_{j}$ with $\tilde{w}_{j}=\widetilde{W}_{m}\tilde{z}_{j}$ and
(61) $\|r_{\mathrm{in},j}\|=\|r^{\prime}_{\mathrm{in},j}\|=\|\widetilde{B}_{m}^{\prime}\tilde{z}_{j}+r^{\prime}\|,\ j=1,2,\ldots,M+N-2m.$
Theorem 4.1 establishes the equivalence of MINRES for solving (56) and (60). Notice that $r=\mathbf{0}$ when $\widetilde{U}_{m}=U_{m}$ and $\widetilde{V}_{m}=V_{m}$. In this case, we can extend Theorem 3.4 to MINRES for solving (56) by supposing any virtual right-hand side $r\not=\mathbf{0}$ that satisfies (51).
Theorem 4.2.
Suppose that $\mathcal{R}(\widetilde{U}_{m})=\mathcal{R}(U_{m})$ and $\mathcal{R}(\widetilde{V}_{m})=\mathcal{R}(V_{m})$ in (56) with $U_{m}$ and $V_{m}$ in (48) and $\mathcal{R}(\cdot)$ the column space of a matrix. Then (i) if $\tau>(\sigma_{\max}+\sigma_{\max,m+1})/2$ with $\sigma_{\max,m+1}$ the $(m+1)$th largest singular value of $A$, then
(62) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\Bigg{(}1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,m+1})/(\tau-\sigma_{\max,m+1})}}\Bigg{)}^{j};$
(ii) if $\sigma_{\min}\leq\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$ with $\sigma_{\min,m+1}$ the $(m+1)$th smallest singular value of $A$ and writing $j_{o}=\min\{1,M-N\}$, then
(63) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\Bigg{(}1-\frac{2}{1+\sqrt{(\sigma_{\max}^{2}-\tau^{2})/(\sigma_{\min,m+1}^{2}-\tau^{2})}}\Bigg{)}^{\left[\frac{j-j_{o}}{2}\right]};$
(iii) if $(\sigma_{\min}+\sigma_{\min,m+1})/2<\tau<(\sigma_{\max}+\sigma_{\max,m+1})/2$, then
(64) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau)/|\sigma_{m+1}-\tau|}\right)^{\left[\frac{j}{2}\right]}$
with $\sigma_{m+1}$ the $(m+1)$th closest singular value to $\tau$ in terms of the labeling rule (2).
Proof 4.3.
Under the hypothesis, we take $\widetilde{P}_{m}=P_{m}$ and $\widetilde{Q}_{m}=Q_{m}$ in (57) with
(65) $P_{m}=[u_{m+1},\dots,u_{M}]\qquad\mbox{and}\qquad Q_{m}=[v_{m+1},\dots,v_{N}].$
In this case, the matrix $\widetilde{B}_{m}^{\prime}$ defined in (58) reduces to
(66) $B_{m}^{\prime}=\begin{bmatrix}-\tau I_{M-m}&P_{m}^{T}AQ_{m}\\ Q_{m}^{T}A^{T}P_{m}&-\tau I_{N-m}\end{bmatrix}=\begin{bmatrix}-\tau I_{M-m}&\Sigma_{m}^{\prime}\\ (\Sigma_{m}^{\prime})^{T}&-\tau I_{N-m}\end{bmatrix},$
where
(67) $\Sigma_{m}^{\prime}=P_{m}^{T}AQ_{m}=\operatorname{diag}\{\sigma_{m+1},\dots,\sigma_{N}\}\in\mathbb{R}^{(M-m)\times(N-m)}.$
Therefore, the eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ are
(68) ±$\pm\sigma_{m+1}-\tau,\dots,\pm\sigma_{N}-\tau\qquad\mbox{and}\qquad\underbrace{-\tau,\dots,-\tau}_{M-N}.$
(i) In this case, from (47) the $m$ singular values $\sigma_{1}$, $\dots$, $\sigma_{m}$ closest to $\tau$ are the $m$ largest ones $\sigma_{\max}$, $\sigma_{\max,2}$, $\dots$, $\sigma_{\max,m}$ of $A$, and the remaining singular values $\sigma_{m+1},\ldots,\sigma_{N}$ are smaller than $\tau$, among which $\sigma_{m+1}=\sigma_{\max,m+1}$. Therefore, the eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ in (68) are negative and lie in
$[-(\sigma_{\max,m+1}+\tau),-(\tau-\sigma_{\max,m+1})].$
Applying (20) to MINRES for (56) leads to (62).
(ii) In this case, the $m$ desired singular values are the $m$ smallest ones $\sigma_{\min}$, $\sigma_{\min,2}$, $\dots$, $\sigma_{\min,m}$ of $A$, and $\sigma_{m+1}=\sigma_{\min,m+1}$. The eigenvalues $\lambda_{i}$ of $B_{m}^{\prime}$ in (68) lie in
$\{-\tau\}\cup[-(\sigma_{\max}+\tau),-(\sigma_{\min,m+1}+\tau)]\cup[\sigma_{\min,m+1}-\tau,\sigma_{\max}-\tau],$
where the negative and positive intervals have the equal length $\sigma_{\max}-\sigma_{\min,m+1}$. Using a similar proof approach to that of (23), we obtain (63).
(iii) In this case, the $\lambda_{i}$ in (68) lie in
$[-(\sigma_{\max}+\tau),-|\tau-\sigma_{m+1}|]$$\cup$ $[|\tau-\sigma_{m+1}|,\sigma_{\max}-\tau]$
(69) $\subset\ [-(\sigma_{\max}+\tau),-|\tau-\sigma_{m+1}|]$$\cup$ $[|\tau-\sigma_{m+1}|,\sigma_{\max}+\tau].$
Notice that the two intervals in (69) have the same length $\sigma_{\max}+\tau-|\tau-\sigma_{m+1}|$. Applying (21) to MINRES for (56) with this interval union yields (64).
Remark 4.4.
Note that items (i)–(iii) of Theorem 4.2 correspond to $\sigma_{1}=\sigma_{\max},\sigma_{\min}$ and $\sigma_{1}\not=\sigma_{\max},\sigma_{\min}$. The convergence rates of MINRES for (56) thus depend on
(70) $\gamma_{4}=\frac{\tau-\sigma_{\max,m+1}}{\tau+\sigma_{\max,m+1}},\qquad\gamma_{5}=\frac{\sigma_{\min,m+1}^{2}-\tau^{2}}{\sigma_{\max}^{2}-\tau^{2}},\qquad\gamma_{6}=\frac{|\sigma_{m+1}-\tau|}{\sigma_{\max}+\tau},$
which are considerably bigger than $\gamma_{1},\gamma_{2}$ and $\gamma_{3}$ in (30) whenever $\tau-\sigma_{\max,m+1}\gg\tau-\sigma_{\max,2}$, $\sigma_{\min,m+1}-\tau\gg\sigma_{\min,2}-\tau$ and $|\sigma_{m+1}-\tau|\gg|\sigma_{2}-\tau|$, that is, the $m$ clustered singular values $\sigma_{1},\ldots,\sigma_{m}$ are well separated from the others. As a result, MINRES for (56) will converge much faster than it does for (9) correspondingly.
In computations, $\mathcal{R}(\widetilde{U}_{m})\neq\mathcal{R}(U_{m})$ and $\mathcal{R}(\widetilde{V}_{m})\neq\mathcal{R}(V_{m})$. Denote by
(71) $\Phi=\angle\left(\mathcal{R}(\widetilde{U}_{m}),\mathcal{R}(U_{m})\right)\quad\mbox{and}\quad\Psi=\angle\left(\mathcal{R}(\widetilde{V}_{m}),\mathcal{R}(V_{m})\right)$
the canonical angle matrices between $\mathcal{R}(\widetilde{U}_{m})$ and $\mathcal{R}(U_{m})$ and between $\mathcal{R}(\widetilde{V}_{m})$ and $\mathcal{R}(V_{m})$, respectively (cf. [3, p. 329]). For $\widetilde{B}_{m}^{\prime}$ in (58) and $B_{m}^{\prime}$ in (66), we have the following extension of Theorem 3.8.
Theorem 4.5.
$\widetilde{B}_{m}^{\prime}$ is orthogonally similar to $B_{m}^{\prime}$ with the error
(72) $\|\widetilde{B}_{m}^{\prime}-Z^{T}B_{m}^{\prime}Z\|\leq\|A\|\left(\|\sin\Phi\|+\|\sin\Psi\|\right)^{2},$
where $Z$ is some orthogonal matrix in $\mathbb{R}^{(M+N-2m)\times(M+N-2m)}$; see (79).
Proof 4.6.
By (65), let us decompose the orthonormal $\widetilde{P}_{m}$ and $\widetilde{Q}_{m}$, defined in the beginning of Section 4.2, as the orthogonal sums
(73) $\widetilde{P}_{m}=P_{m}E_{1}+U_{m}E_{2}\qquad\mbox{and}\qquad\widetilde{Q}_{m}=Q_{m}F_{1}+V_{m}F_{2}$
with $E_{1}\in\mathbb{R}^{(M-m)\times(M-m)}$, $E_{2}\in\mathbb{R}^{m\times(M-m)}$, $F_{1}\in\mathbb{R}^{(N-m)\times(N-m)}$, $F_{2}\in\mathbb{R}^{m\times(N-m)}$. Therefore,
(74) $\|\sin\Phi\|=\|\widetilde{P}_{m}^{T}U_{m}\|=\|E_{2}\|\qquad\mbox{and}\qquad\|\sin\Psi\|=\|\widetilde{Q}_{m}^{T}V_{m}\|=\|F_{2}\|.$
Let the SVDs of $E_{1}$ and $F_{1}$ be
(75) $E_{1}=X_{1}\Sigma_{E}X_{2}^{T}\qquad\mbox{and}\qquad F_{1}=Y_{1}\Sigma_{F}Y_{2}^{T}.$
Using the same approach to proving (35), we obtain
(76) $\|I-\Sigma_{E}\|\leq\|\sin\Phi\|^{2}\qquad\mbox{and}\qquad\|I-\Sigma_{F}\|\leq\|\sin\Psi\|^{2}.$
By the definitions of $P_{m},Q_{m}$ and $U_{m},V_{m}$, we have $U_{m}^{T}AV_{m}=\Sigma_{m}$, $P_{m}^{T}AQ_{m}=\Sigma_{m}^{\prime}$ and $P_{m}^{T}AV_{m}=\bm{0}$, $U_{m}^{T}AQ_{m}=\bm{0}$, where $\Sigma_{m}$ and $\Sigma_{m}^{\prime}$ are defined in (48) and (67). Therefore, exploiting (73) and (75), we have
(77) $\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}=E_{1}^{T}\Sigma_{m}^{\prime}F_{1}+E_{2}\Sigma_{m}F_{2}=X_{2}\Sigma_{E}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}Y_{2}^{T}+E_{2}\Sigma_{m}F_{2}.$
Note that $X_{i},Y_{i},i=1,2$ are orthogonal and $\|\Sigma_{m}^{\prime}\|\leq\|A\|$, $\|\Sigma_{m}\|\leq\|A\|$. Exploiting (74), (76) and (77), we obtain
(78) $\|\!\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}\!-\!X_{\!2}X_{1}^{T}\Sigma_{m}^{\prime}Y_{\!1}Y_{\!2}^{T}\|$ $=$ $\|\Sigma_{E}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}-X_{1}^{T}\!\Sigma_{m}^{\prime}Y_{\!1}+X_{2}^{T}E_{2}\Sigma_{m}F_{2}Y_{2}\|$
$\leq$ $\|\!(\Sigma_{E}\!-\!I)X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}\Sigma_{F}\!+\!X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}(\Sigma_{F}\!-\!I)\!\|\!+\!\|\!E_{2}\Sigma_{m}F_{2}\!\|$
$\leq$ $\|\Sigma_{E}-I\|\|\Sigma_{m}^{\prime}\|+\|\Sigma_{m}^{\prime}\|\|\Sigma_{F}-I\|+\|E_{2}\|\|\Sigma_{m}\|\|F_{2}\|$
$\leq$ $\|A\|(\|\sin\Phi\|^{2}+\|\sin\Psi\|^{2}+\|\sin\Phi\|\|\sin\Psi\|)$
$\leq$ $\|A\|(\|\sin\Phi\|+\|\sin\Psi\|)^{2}.$
Denote the orthogonal matrix
(79) $Z=\begin{bmatrix}X_{1}X_{2}^{T}&\\ &Y_{1}Y_{2}^{T}\end{bmatrix}\in\mathbb{R}^{(M+N-2m)\times(M+N-2m)}.$
By the definitions of $\widetilde{B}_{m}^{\prime}$ in (58) and $B_{m}^{\prime}$ in (66), we obtain
$\|\widetilde{B}_{m}^{\prime}-Z^{T}B_{m}^{\prime}Z\|=\|\widetilde{P}_{m}^{T}A\widetilde{Q}_{m}-X_{2}X_{1}^{T}\Sigma_{m}^{\prime}Y_{1}Y_{2}^{T}\|.$
Then (72) follows from applying (78) to the above.
Making use of Theorems 4.2 and 4.5, we can establish the following convergence results on MINRES for solving (56). The proof is analogous to that of Theorem 3.10.
Theorem 4.7.
With the notations of Theorem 4.5, assume that
(80) $\delta:=\|A\|(\|\sin\Phi\|+\|\sin\Psi\|)^{2}<|\sigma_{m+1}-\tau|.$
(i) If $\tau>(\sigma_{\max}+\sigma_{\max,m+1})/2$, then
(81) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\Bigg{(}1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,m+1}+\delta)/(\tau-\sigma_{\max,m+1}-\delta)}}\Bigg{)}^{j};$
(ii) If $\sigma_{\min}\leq\tau<(\sigma_{\min}+\sigma_{\min,m+1})/2$ and writing $j_{o}=\min\{1,M-N\}$, then
(82) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}+\delta}{\tau}\right)^{j_{o}}\Bigg{(}1-\frac{2}{1+\sqrt{\frac{(\sigma_{\max}+\delta)^{2}-\tau^{2}}{(\sigma_{\min,m+1}-\delta)^{2}-\tau^{2}}}}\Bigg{)}^{\left[\frac{j-j_{o}}{2}\right]}.$
(iii) If $(\sigma_{\min}+\sigma_{\min,m+1})/2<\tau<(\sigma_{\max}+\sigma_{\max,m+1})/2$, then
(83) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau+\delta)/(|\sigma_{m+1}-\tau|-\delta)}\right)^{\left[\frac{j}{2}\right]}.$
The factors in (81)–(83) are approximations to $\gamma_{4},\gamma_{5}$ and $\gamma_{6}$ in (70) with the relative errors $\mathcal{O}(\frac{\delta}{\tau-\sigma_{\max,m+1}})$, $\mathcal{O}(\frac{\delta}{\sigma_{\min,m+1}-\tau})$ and $\mathcal{O}(\frac{\delta}{|\sigma_{m+1}-\tau|})$, respectively. This indicates that the convergence behavior of MINRES for (56) resembles that in the ideal case as $(\Theta_{m},\widetilde{U}_{m},\widetilde{V}_{m})$ converges. If $\sigma_{1}$ is close to $\sigma_{2}$ but well separated from $\sigma_{m+1}$, MINRES for the preconditioned correction equation (56) can converge much faster than it does for the standard correction equation (9).
4.3 Selection strategy of approximate singular triplets in inner preconditioning
The previous results and analysis have shown that it is much better to use MINRES to solve the preconditioned (56) rather than (9) when there are totally $m$ singular values of $A$ clustered at $\tau$ and their approximate singular vectors already have some accuracy. However, the number $m$ is practically unknown. We need to set up necessary criteria so as to propose a practical IPJDSVD algorithm.
Quantitatively, we will claim a singular value $\sigma_{i}$ clustered at $\tau$ if
(84) $\frac{|\sigma_{i}-\tau|}{\max\{\sigma_{i},1\}}\leq\textit{pretol}_{1}$
for a reasonably small $\textit{pretol}_{1}$. Theorems 4.5 and 4.7 indicate that if the singular vectors with the $m$ clustered singular values are computed with some accuracy, say $\|\sin\Phi\|+\|\sin\Psi\|\leq\textit{pretol}_{2}$ for a reasonably small $\textit{pretol}_{2}$, then we can benefit very much from MINRES for solving the preconditioned (56).
During each outer iteration, for the currently computed approximate singular triplet $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=2,\dots,k$, if $\theta_{i}$ and the residual $r^{(i)}=r(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ in (6) satisfy
(85) $|\theta_{i}-\tau|\leq\max\{\theta_{i},1\}\cdot\textit{pretol}_{1}\qquad\mbox{and}\qquad\|r^{(i)}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{pretol}_{2},$
then we claim $\theta_{i}$ clustered at $\tau$ and accept $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ as a reasonably good approximate singular triplet of $A$, where $\|A\|_{\mathrm{e}}=\sqrt{\|A\|_{1}\cdot\|A\|_{\infty}}$ is a rough estimate of $\|A\|=\sigma_{\max}$ with $\|A\|_{1}$ and $\|A\|_{\infty}$ the $1$- and $\infty$-norms of $A$, respectively. We then take the $\tilde{u}_{i},\tilde{v}_{i}$ to form
(86) $\widetilde{U}_{\tilde{m}}=[\tilde{u}_{1},\tilde{u}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]\qquad\mbox{and}\qquad\widetilde{V}_{\tilde{m}}=[\tilde{v}_{1},\tilde{v}_{i_{2}},\dots,\tilde{v}_{i_{\tilde{m}}}],$
where the indices $i_{2},\dots,i_{\tilde{m}}\in\{2,\dots,k\}$ are such that (85) is fulfilled, and replace $\widetilde{U}_{m}$ and $\widetilde{V}_{m}$ by them in (56).
For $k$ small, there may be no sufficiently good approximate singular triplet that satisfies (85). In this case, $\tilde{m}=1$ in (86), and (56) degenerates to (9). However, for $m>1$, as the dimension $k$ increases, more and more computed approximate singular triplets start to converge and fulfill (85), so that $\tilde{m}>1$ pairs of approximate singular vectors participate (56). Then the inner preconditioning starts to show its strength. In computations, $\tilde{m}$ in (86) is determined dynamically according to (85). $\tilde{m}$ increases from 1 to $m$, and we have $\tilde{m}=m$ ultimately for $k$ sufficiently large.
5 A thick-restart IPJDSVD algorithm with deflation and purgation
For practical purpose, we must limit the subspace dimension $k\leq k_{\max}$. If JDSVD and IPJDSVD do not yet converge until $k=k_{\max}$, we must restart them. Furthermore, if we are required to compute $\ell>1$ singular triplets of $A$ corresponding to the $\ell$ singular values $\sigma_{1},\ldots,\sigma_{\ell}$ closest to the target $\tau$ (cf. (2)), an appropriate deflation is necessary. Besides, it appears that a certain purgation is crucial and may greatly reduce the outer iterations for computing the second to the $\ell$th singular triplets.
5.1 A new thick-restart
If IPJDSVD does not yet converge for computing the desired $\ell$ singular triplets for a given $k_{\max}$, we modify a commonly used thick restart, initially proposed in [24] and later popularized in the JDSVD and JDGSVD type methods [9, 10, 11, 28], and propose a more effective thick-restart IPJDSVD than the standard thick-restart JDSVD. Our thick-restart retains some $k_{\operatorname{new}}$-dimensional restarting subspaces with
(87) $k_{\operatorname{new}}=\max\{k_{\min},\tilde{m}\}$
where $k_{\min}>1$ is a user-prescribed minimal dimension of the searching subspaces and $\tilde{m}$ is the number of approximate singular triplets of $A$ that are selected to participate in the preconditioned correction equation (56) in the current expansion phase.
Specifically, assume that $\tilde{m}$ approximate singular triplets $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i}),i=1,i_{2},\dots$, $i_{\tilde{m}}$ of $A$ with $i_{2},\dots,i_{\tilde{m}}\in\{2,\dots,k\}$ join the preconditioned (56) before restart. If $\tilde{m}\leq k_{\min}$, we simply restore the standard thick-restart scheme, which forms the orthonormal basis matrices of the new $k_{\min}$-dimensional subspaces $\mathcal{U}$ and $\mathcal{V}$ by
(88) $\widetilde{U}=[\tilde{u}_{1},\tilde{u}_{2},\dots,\tilde{u}_{k_{\min}}],\qquad\widetilde{V}=[\tilde{v}_{1},\tilde{v}_{2},\dots,\tilde{v}_{k_{\min}}],$
and updates the projection matrix of $A$ with respect to these two subspaces by
(89) $H=\widetilde{U}^{T}A\widetilde{V}=\operatorname{diag}\{\theta_{1},\theta_{2},\dots,\theta_{k_{\min}}\}.$
If $\tilde{m}>k_{\min}$, we form the orthonormal basis matrices of the new $\tilde{m}$-dimensional $\mathcal{U}$ and $\mathcal{V}$ and the corresponding projection matrix of $A$ by
(90) $\widetilde{U}=\widetilde{U}_{\tilde{m}},\qquad\widetilde{V}=\widetilde{V}_{\tilde{m}},\qquad H=\widetilde{U}_{\tilde{m}}^{T}A\widetilde{V}_{\tilde{m}}=\operatorname{diag}\{\theta_{1},\theta_{i_{2}},\dots,\theta_{i_{\tilde{m}}}\}$
with $\widetilde{U}_{\tilde{m}}$ and $\widetilde{V}_{\tilde{m}}$ defined by (86). In this way, the $k_{\rm new}$-dimensional restarting subspaces contain approximate left and right singular subspaces associated with all available approximate singular values clustered at $\tau$. Whenever $m<k_{\max}$ considerably, the new thick-restart IPJDSVD is expected to substantially outperform the standard thick-restart JDSVD in terms of the outer iterations.
5.2 Deflation and purgation
Suppose that we are required to compute the $\ell>1$ singular triplets of $A$ corresponding to the $\ell$ singular values closest to the target $\tau$. Then IPJDSVD combined with the deflation advocated in [9] meets this demand perfectly. Concretely, assume that $j<\ell$ approximate singular triplets $\left(\theta_{i,\mathrm{c}},u_{i,\mathrm{c}},v_{i,\mathrm{c}}\right)$ have already converged to $(\sigma_{i},u_{i},v_{i})$, $i=1,\dots,j$ with the residual $r^{(i,\mathrm{c})}=r(\theta_{i,\mathrm{c}},u_{i,\mathrm{c}},v_{i,\mathrm{c}})$ defined by (6) satisfying
(91) $\|r^{(i,\mathrm{c})}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{tol},\qquad i=1,\dots,j$
for a user-prescribed tolerance $\textit{tol}>0$, where $\|A\|_{\rm e}$ is defined as in (85). Denote
$\Sigma_{\mathrm{c}}=\mathrm{diag}\{\theta_{1,\mathrm{c}},\dots,\theta_{j,\mathrm{c}}\},\qquad U_{\mathrm{c}}=[u_{1,\mathrm{c}},\dots,u_{j,\mathrm{c}}],\qquad V_{\mathrm{c}}=[v_{1,\mathrm{c}},\dots,v_{j,\mathrm{c}}].$
Then $\left(\Sigma_{\mathrm{c}},U_{\mathrm{c}},V_{\mathrm{c}}\right)$ is a converged approximation to the partial SVD $(\Sigma_{j},U_{j},V_{j})$ of $A$ in (48). Suppose that the next singular triplet $(\sigma_{j+1},u_{j+1},v_{j+1})$ of $A$ is desired. At the extraction phase, the deflation in [9] requires that IPJDSVD implements a chosen extraction approach, e.g., the standard extraction, with respect to the $k$-dimensional $\mathcal{U}$ and $\mathcal{V}$ that are orthogonal to $\mathcal{R}(U_{\mathrm{c}})$ and $\mathcal{R}(V_{\mathrm{c}})$, respectively; we obtain approximate singular triplets $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=1,\dots,k$ as described in Section 2, and take $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ as the approximation to $(\sigma_{j+1},u_{j+1},v_{j+1})$.
If $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ does not converge, we take $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$, $i=1,i_{2},\dots,i_{\tilde{m}}$ that meet (85) to form $\widetilde{U}_{\tilde{m}}$ and $\widetilde{V}_{\tilde{m}}$ in (86). Then we set $U_{\mathrm{p}}=[U_{\mathrm{c}},\widetilde{U}_{\tilde{m}}]$ and $V_{\mathrm{p}}=[V_{\mathrm{c}},\widetilde{V}_{\tilde{m}}]$, and use MINRES to approximately solve the new preconditioned correction equation
(92) $\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}\!\!&\\ &\!\!I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}\!\!&\\ &\!\!I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=r_{\mathrm{p}}$
for $(s,t)\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$, where $r_{\mathrm{p}}=-\begin{bmatrix}I-U_{\mathrm{c}}U_{\mathrm{c}}^{T}\!\!\!&\\ &\!\!\!I-V_{\mathrm{c}}V_{\mathrm{c}}^{T}\end{bmatrix}r$ with $r=r(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ the residual of $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ defined by (6). We compute an approximate solution $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$ and terminate MINRES when the inner residual $r_{\mathrm{in}}$ in (59) satisfies
(93) $\|r_{\mathrm{in}}\|\leq\|r\|\cdot\min\{{\rho\tilde{\varepsilon}},0.1\},$
where $\tilde{\varepsilon}\in[10^{-4},10^{-3}]$ is a user-prescribed tolerance and $\rho$ is a constant that is estimated by using all the approximate singular values computed at the current cycle; see [9] for details. We use the approximate solution $\tilde{s}$ and $\tilde{t}$ to expand the subspaces $\mathcal{U}$ and $\mathcal{V}$. Since $(\widetilde{U},\widetilde{V})\perp\perp(U_{\mathrm{c}},V_{\mathrm{c}})$ and $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{c}},V_{\mathrm{c}})$, the expanded $\mathcal{U}$ and $\mathcal{V}$ are automatically orthogonal to $\mathcal{R}(U_{\mathrm{c}})$ and $\mathcal{R}(V_{\mathrm{c}})$, respectively. IPJDSVD then proceeds in a regular way until convergence. We then assign $\left(\theta_{j+1,\mathrm{c}},u_{j+1,\mathrm{c}},v_{j+1,\mathrm{c}}\right):=(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$, add it to the converged $\left(\Sigma_{\mathrm{c}},U_{\mathrm{c}},V_{\mathrm{c}}\right)$, and set $j=j+1$. Proceed in such a way until all the $\ell$ singular triplets are found.
In the deflated IPJDSVD method described above, we now have $1\leq\tilde{m}\leq m-j$ for $j<m$ since we have already computed $j$ converged singular triplets and the deflated matrix has only $m-j$ clustered singular values near $\tau$. For $j\geq m$, the preconditioned correction equations degenerate to the standard ones, and IPJDSVD reduces to JDSVD.
For the purgation issue, we may benefit much from making full use of the current $\mathcal{U}$ and $\mathcal{V}$ when starting to compute the next singular triplet $(\sigma_{j+1},u_{j+1},v_{j+1})$, and, compared to the computation of $(\sigma_{1},u_{1},v_{1})$, save outer iterations considerably when computing $(\sigma_{i},u_{i},v_{i}),\ i=2,\ldots,\ell$. The purgation approach [9] can be straightforwardly introduced into the thick-restart IPJDSVD.
Algorithm 1 Thick-restart IPJDSVD with deflation and purgation for the target $\tau$.
1:  Initialization: Set $k=1$, $k_{\mathrm{c}}=0$, $\Sigma_{\mathrm{c}}=[\ ],U_{\mathrm{c}}=[\ ]$, and $V_{\mathrm{c}}=[\ ]$. Let $\widetilde{U}=[\ ]$, $\widetilde{V}=[\ ]$, $H=[\ ]$, and choose normalized starting vectors $u_{+}=u_{0}$ and $v_{+}=v_{0}$.
2:  Calculate $\|A\|_{e}=\sqrt{\|A\|_{1}\cdot\|A\|_{\infty}}$.
3:  while $k\geq 0$ do
4:      Set $\widetilde{U}=[\widetilde{U},u_{+}]$ and $\widetilde{V}=[\widetilde{V},v_{+}]$, and update $H=\widetilde{U}^{T}A\widetilde{V}$.
5:      Compute the singular triplets $(\theta_{i},c_{i},d_{i}),i=1,\dots,k$ of $H$ with $\theta_{i}$ ordered increasingly by their distances from $\tau$. Form the approximate singular triplets $(\theta_{i},\tilde{u}_{i}\!=\widetilde{U}c_{i},\tilde{v}_{i}\!=\widetilde{V}d_{i})$ and the residuals $r^{(i)}\!=\!\begin{bmatrix}\begin{smallmatrix}A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i}\\ A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i}\end{smallmatrix}\end{bmatrix}$, $i\!=\!1,\dots,k$.
6:     if $\|r^{(1)}\|\leq\|A\|_{\mathrm{e}}\cdot\textit{tol}$ then
7:        Update $\Sigma_{\mathrm{c}}=\operatorname{diag}\{\Sigma_{\mathrm{c}},\theta_{1}\},U_{\mathrm{c}}=[U_{\mathrm{c}},\tilde{u}_{1}]$, $V_{\mathrm{c}}=[V_{\mathrm{c}},\tilde{v}_{1}]$, and set $k_{c}=k_{c}+1$.
8:        if $k_{\mathrm{c}}=\ell$ then return $(\Sigma_{\mathrm{c}},U_{\mathrm{c}},V_{\mathrm{c}})$ and stop. fi
9:         Perform the purgation, set $k=k-1$, and go to step 5.
10:     end if
11:     Determine the $\tilde{m}-1$ indices $i=i_{2},\dots,i_{\tilde{m}}$ according to the condition in (85), and form $\widetilde{U}_{\tilde{m}}=[\tilde{u}_{1},\tilde{u}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]$ and $\widetilde{V}_{\tilde{m}}=[\tilde{v}_{1},\tilde{v}_{i_{2}},\dots,\tilde{u}_{i_{\tilde{m}}}]$.
12:     Set $U_{\mathrm{p}}=[U_{\mathrm{c}},\widetilde{U}_{\tilde{m}}]$, $V_{\mathrm{p}}=[V_{\mathrm{c}},\widetilde{V}_{\tilde{m}}]$ and $r_{\mathrm{p}}=-\begin{bmatrix}\begin{smallmatrix}I-U_{\mathrm{c}}U_{\mathrm{c}}^{T}&\\ &I-V_{\mathrm{c}}V_{\mathrm{c}}^{T}\end{smallmatrix}\end{bmatrix}r^{(1)}$, and solve
$\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}&\\ &I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-U_{\mathrm{p}}U_{\mathrm{p}}^{T}&\\ &I-V_{\mathrm{p}}V_{\mathrm{p}}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=r_{\mathrm{p}}$
for the approximate solution $(\tilde{s},\tilde{t})\perp\perp(U_{\mathrm{p}},V_{\mathrm{p}})$ with the inner residual norm $\|r_{\mathrm{in}}\|$ satisfying the stopping criterion (93).
13:     if $k=k_{\max}$ then perform the thick-restart and reset $k=k_{\operatorname{new}}$ in (87). fi
14:      Orthonormalize $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ to obtain the expansion vectors $u_{+}$ and $v_{+}$, respectively, and set $k=k+1$.
15:  end while
5.3 Thick-restart IPJDSVD algorithm with deflation and purgation
Algorithm 1 sketches our thick-restart IPJDSVD with deflation and purgation. It requires the devices to form the matrix-vector multiplications with $A$ and $A^{T}$, the target $\tau$, the number $\ell$ of desired singular triplets, and the outer stopping tolerance tol. It produces a converged partial SVD $(\Sigma_{c},U_{c},V_{c})$ that satisfies
(94) $\sqrt{\|AV_{c}-U_{c}\Sigma_{c}\|_{\mathrm{F}}^{2}+\|A^{T}U_{c}-V_{c}\Sigma_{c}\|_{\mathrm{F}}^{2}}\leq\sqrt{\ell}\|A\|_{\mathrm{e}}\cdot\textit{tol},$
where $\|\cdot\|_{\rm F}$ is the F-norm of a matrix. Other optional parameters are normalized starting vectors $u_{0}\in\mathbb{R}^{M}$ and $v_{0}\in\mathbb{R}^{N}$, the maximum and minimum dimensions $k_{\max}$ and $k_{\min}$ of the searching subspaces, the accuracy requirement $\tilde{\varepsilon}$ for expansion vectors in (93), and the preconditioning parameters $\textit{pretol}_{1}$ and $\textit{pretol}_{2}$ in (85). By defaults, we set $tol=10^{-8}$, $u_{0}=\frac{1}{\sqrt{M}}[1,\dots,1]$, $v_{0}=\frac{1}{\sqrt{N}}[1,\dots,1]$, $k_{\max}=30$, $k_{\min}=3$, $\tilde{\varepsilon}=10^{-4}$, $\textit{pretol}_{1}=0.05$ and $\textit{pretol}_{2}=0.01$.
When going back to step 5 from step 9, the SVD of the updated projection matrix $H$ is known explicitly, so are the approximate singular triplets and corresponding residuals. As a matter of fact, the orthonormal bases of $\mathcal{U}$ and $\mathcal{V}$ after purgation and the corresponding projection matrix of $A$ are
(95) $\widetilde{U}=[\tilde{u}_{2},\dots,\tilde{u}_{k}],\qquad\widetilde{V}=[\tilde{v}_{2},\dots,\tilde{v}_{k}],\qquad H=\mathrm{diag}\{\theta_{2},\dots,\theta_{k}\}.$
Therefore, IPJDSVD only needs to reorder all the approximate singular triplets, reassign $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1}):=(\theta_{2},\tilde{u}_{2},\tilde{v}_{2})$, and test its convergence.
6 Numerical experiments
We report numerical experiments to illustrate the considerable superiority of IPJDSVD to JDSVD. We have also developed a hybrid two-stage IPJDSVD algorithm in Matlab, which is based on the same motivation and is along the same line as that of the state-of-the-art two-stage software package PRIMME_SVDS in C language developed in [27] for accurately computing several extreme singular triplets. In the first stage of IPJDSVD_HYBRID we perform a similar IPJDSVD method on the eigenproblem of the cross product matrix $A^{T}A$ or that of $AA^{T}$ if $m<n$ with the same outer stopping tolerance as that of PRIMME_SVDS, where $A$ has full column or row rank; a convergence analysis of IPJDSVD on such eigenproblem can be made similarly and is adaptable from the current context, and we omit it due to space. We will illustrate that IPJDSVD_HYBRID is at least competitive with and can outperform PRIMME_SVDS considerably. All the experiments were implemented on a 12th Gen Intel(R) Core(TM) i7-12700K CPU 3.60GHz with 32 GB main memory and 12 cores using the Matlab R2023b under the Ubuntu 22.04.4 LTS 64 bit system with the machine precision $\epsilon=2.22\times 10^{-16}$.
Table 1: Properties of the test matrices.
$A$ $M$ $N$ nnz $\sigma_{\max}$ $\sigma_{\min}$
Alemdar 6245 6245 42581 69.5 3.30e-3
bcsstk33 8738 8738 591904 75.8 2.18e-95
cat_ears_3_4 5226 13271 39592 5.49 3.06e-1
epb2 25228 25228 175027 3.34 1.28e-3
relat8 345688 12347 1334038 18.8 0
dw8192 8192 8192 41746 110 2.88e-5
garon2 13535 13535 373235 44.9 7.32e-7
tomographic1 73159 59498 647495 6.98 0
3elt 4720 4720 27444 6.03 2.12e-3
plddb 3069 5049 10839 1.43e+2 1.16e-2
bas1lp 5411 9825 587775 1.49e+3 3.62e-2
G66 9000 9000 36000 3.58 2.14e-4
stat96v4 3173 63076 491336 1.96e+1 2.07e-3
atmosmodl 1489752 1489752 10319760 6.21e+5 5.42e+2
cage15 5154859 5154859 99199551 1.02 8.56e-2
Table 1 lists the test matrices from the SuiteSparse Matrix Collection [2] together with some of their basic properties, where nnz denotes the total number of nonzero entries in $A$, the largest and smallest singular values $\sigma_{\max}$ and $\sigma_{\min}$ of $A$ are, only for experimental purpose, computed by the Matlab functions svds and svd for the very large matrices “relat8”, “tomographic1”, “atmosmodl”, “cage15” and all the other moderately large matrices, respectively.
In all the tables, denote by $I_{\mathrm{out}}$ and MVs the total number of correction equations and matrix-vector multiplications with $A$ and $A^{T}$ that the underlying algorithm was run for the same outer convergence tolerance, and by $T_{\mathrm{cpu}}$ the total CPU time in seconds. Notice that MVs dominates the overall efficiency of each algorithm. Since our codes were written in Matlab and the advanced PRIMME_SVDS was optimally programmed in C language, the latter should normally be much more efficient than ours in terms of CPU time if the same MVs are used.
6.1 A comparison of IPJDSVD and JDSVD
We took the default parameters as described in Section 5.3. For (9), (56) and (92), we took zero vector as the initial guesses and used the Matlab function minres to solve them until the relative residual norms met (93). To illustrate the wide applicability of our algorithms, we took $\tau$ at will, so that they correspond to both extreme and interior singular triplets.
[…]
Table 2: Results on computing the $10$ singular triplets with given targets.
$A$ $\tau$ Algorithm: IPJDSVD Algorithm: JDSVD
$I_{\mathrm{out}}$ MVs $T_{\rm cpu}$ $I_{\mathrm{out}}$ MVs $T_{\rm cpu}$
Alemdar 67.2 53 15998 1.63 59 24928 2.22
cat_ears_3_4 4.9 72 15262 2.32 86 22974 2.70
epb2 3.3 51 10968 3.50 60 19396 4.68
relat8 16 64 17036 77.2 70 25276 89.6
bcsstk33 56 75 15176 4.47 84 23680 6.12
dw8192 109.7 68 9952 1.59 227 26404 3.06
garon2 23 81 99810 22.2 91 161248 34.2
tomographic1 5.6 55 23494 28.7 61 36832 36.3
We computed $10$ singular triplets of eight test matrices in Table 1 with given targets $\tau$, where the desired singular values of the first four matrices are not very interior, and those of the last four matrices are interior and clustered ones. Both algorithms succeeded in computing all the desired singular triplets of all the test matrices and Table 2 lists the results. Regarding $I_{\operatorname{out}}$, we observe that IPJDSVD converged faster for all the test matrices and particularly much faster for “cat_ears_3_4” and “dw8192” and “garon2” than JDSVD, confirming the higher effectiveness of the new thick-restart. Regarding the overall efficiency, Fig. 1 depicts the average MVs per outer iteration of IPJDSVD and JDSVD to compute each desired singular triplet of “Alemdar” and “bcsstk33”. We can see that IPJDSVD, on average, saved $121$ and $80$ MVs at each outer iteration from JDSVD for the former and latter matrices, respectively, huge improvements. For all the test matrices, we can see from Table 2 that IPJDSVD reduced more than $32.74\%$ of MVs and more than $13.79\%$ of the CPU time by JDSVD.
Obviously, for these matrices, IPJDSVD is substantially superior to JDSVD.
6.2 A comparison of IPJDSVD_HYBRID and PRIMME_SVDS
We compare IPJDSVD_HYBRID with PRIMME_SVDS. For PRIMME_SVDS, we took the method in the first stage as the default ‘DYNAMIC’, dynamically switching between JDQMR and the GD+k method to minimize the time, and in the second stage we took the recommended ‘JDQMR_ETol’. For both algorithms, we took the maximum and minimum subspace dimensions $k_{\max}=30$ and $k_{\min}=3$ and the initial $u_{0}$ and $v_{0}$ as the same $m$- and $n$-dimensional random vectors generated by the Matlab function randn. All the other parameters in the two algorithms were set by default.
Table 3: Results on computing the $10$ largest singular triplets.
$A$ IPJDSVD_HYBRID PRIMME_SVDS IRLBD
$I_{\mathrm{out}}$ MVs $T_{\rm cpu}$ MVs $T_{\rm cpu}$ MVs $T_{\rm cpu}$
Alemdar 77 1804 2.05e-1 1858 1.27e-1 472 1.05e-1
3elt 103 2196 2.64e-1 2362 1.41e-1 774 1.58e-1
plddb 81 7680 3.46e-1 3806 1.12e-1 122820 20.1
bas1lp 46 522 3.56e-1 484 1.26e-1 86 6.44e-2
G66 83 2040 2.73e-1 2354 1.68e-1 726 1.72e-1
stat96v5 82 1376 2.82e-1 1306 2.10e-1 250 1.56e-1
atmosmodl 137 11334 6.05e+2 12926 2.02e+2 100184 3.97e+3
cage15 136 5364 1.02e+3 5528 4.99e+2 7168 1.12e+3
Table 4: Results on computing the $10$ smallest singular triplets.
$A$ IPJDSVD_HYBRID PRIMME_SVDS PRIMME_SVDS_J
$I_{\mathrm{out}}$ MVs $T_{\rm cpu}$ MVs $T_{\rm cpu}$ MVs $T_{\rm cpu}$
Alemdar 81 191462 10.6 312848 27.4 247150 69.0
3elt 190 117838 6.39 157086 12.2 146348 84.2
plddb 85 26484 6.59e-1 36328 1.77 37478 2.19
bas1lp 72 337718 69.2 416200 130 537622 189
G66 324 237528 29.5 81000005 9.19e+3 321304 77.9
stat96v5 171 16852 1.78 24972 5.78 24008 6.31
atmosmodl 62 625448 8.41e+3 382610 9.00e+3 313958 1.64e+4
cage15 90 4034 727 5316 722 2882 3.42e+3
We computed the $10$ largest and smallest singular triplets of the eight matrices in Table 1 with the outer stopping tolerance $\textit{tol}=10^{-12}$ by setting the target $\tau=\|A\|_{e}$ and 00 for IPJDSVD_HYBRID and the Sigma as ‘L’ and ‘S’ for PRIMME_SVDS, respectively. Table 4 reports the results. For the largest singular triplets, we also report the results obtained by the implicitly restarted LBD algorithm (IRLBD) [14] with the same parameters. For the smallest singular triplets, we also report the results computed by PRIMME_SVDS with the block Jacobi preconditioners (PRIMME_SVDS_J) with block size $10$ and $100$ for the first six and last two matrices, respectively.
When computing the largest singular triplets, IPJDSVD_HYBRID and PRIMME
_SVDS consumed comparable MVs and even runtime for all the matrices but “plddb”, and IRLBD outperformed them for five of the eight matrices by using significantly fewer MVs and less CPU time. For “plddb”, “atmosmodl” and “cage15”, the former two algorithms were substantially superior to IRLBD. We notice that IRLBD consumed much more MVs than PRIMME_SVDS for “plddb” and “atmosmodl”, and IPJDSVD_HYBRID did similarly for “plddb”. This is because some of the desired largest singular values are multiple, causing that IRLBD and IPJDSVD_HYBRID perform poorly while PRIMME_SVDS with block size bigger than or equal to the maximum multiplicity is able to compute multiple singular values.
As a whole, for computing the simple largest singular triplets, IPJDSVD_HYBRID is at least competitive with PRIMME_SVDS; IRLBD is generally preferred if the largest singular values are not clustered, otherwise the former algorithms may be advantageous.
For the computation of smallest singular triplets, except “atmosmodl” for which IPJDSVD_HYBRID used nearly as twice MVs as PRIMME_SVDS, for the other seven matrices, IPJDSVD_HYBRID performed considerably better than the latter regarding total MVs and CPU time; on average, IPJDSVD_HYBRID saves $25\%$ of MVs for each matrix, and it even used much less runtime than PRIMME_SVDS. Particularly, for “G66”, IPJDSVD_HYBRID converged fast and computed all the desired singular triplets correctly, while PRIMME_SVDS failed to converge when the maximum $N^{2}$ MVs were performed and only output five converged singular triplets.
When block Jacobi preconditioners were exploited in PRIMME_SVDS for the involved correction equations, we see they only accelerated the convergence of the inner iterations for “G66” and “cage15” dramatically and did not help for the other matrices in terms of CPU time and almost all the MVs. Overall, IPJDSVD_HYBRID outmatched PRIMME_SVDS_J considerably in terms of both MVs and CPU time, except for the last two matrices where IPJDSVD_HYBRID used more MVs. Therefore, IPJDSVD_HYBRID outperforms PRIMME_SVDS and PRIMME_SVDS_J greatly for computing the simple smallest singular triplets.
7 Conclusions
We have analyzed the convergence of MINRES for the correction equations in JDSVD, and proposed an inner preconditioning procedure for the correction equation and derived the preconditioned correction equation, whose solution has been proven to expand the subspaces as effectively as that to the standard correction equation. The resulting IPJDSVD method is essentially a new JDSVD method with a new correction equation solved at each outer iteration. We have established convergence results on the preconditioned MINRES, and shown that it can converge much faster than MINRES does for solving the standard correction equation.
We have presented a practical strategy that dynamically selects approximate singular triplets from current available ones and forms the preconditioned correction equation. Finally, we have developed a new thick-restart IPJDSVD algorithm with deflation and purgation to compute several singular triplets. Numerical experiments have illustrated that the new thick-restart IPJDSVD algorithm outperforms the standard thick-restart JDSVD algorithm considerably in terms of the outer iterations, MVs and CPU time; in the meantime, IPJDSVD_HYBRID is at least competitive with PRIMME_SVD in [27] for computing the largest singular triplets and outperforms the latter substantially for computing the smallest singular triplets in terms of MVs, provided that the desired singular values are simple.
Declarations
The two authors declare that they have no conflict of interest, and they read and approved the final manuscript.